Elements of Object-Oriented Program

Design

Robert “Corky” Cartwright

Copyright 1999-2002

Please send corrections and comments to cork@rice.edu August 9, 2002

Chapter 1

From Scheme to Java

1.1 Introduction

At first glance, Java and Scheme appear to have little in common. Java is written in notation similar to the widely-used C programming language, while Scheme is written in parenthesized prefix notation. In addition, Java expresses computations primarily in terms of operations attached to objects while Scheme expresses computations primarily in terms of functions applied to values— mechanizing the familiar world of algebra. In short, Java is a data-centered language while Scheme is a function-centered language.

Nevertheless, Java and Scheme are surprisingly similar beneath the surface. In this short monograph, we will discover how easy it is for a Scheme programmer to learn to write good Java code. The only obstacles are learning Java notation and learning how to organize programs in object-oriented (data-centered) form.

1.2 What is an Object?

Before discussing the specifics of Java’s object system, let’s define what an object is. Within a computer program, an object consists of

• a finite collection of variables called fields representing the properties of a specific physical or conceptual object, and

• a finite collection of designated operations called methods for observing and changing the fields of that object.

No code other than the designated operations can access or modify object fields. The fields and methods of an object are often called the members of the object. Each member of an object has a unique identifying name.

To make the notion of object more concrete, let us consider a simple example. Assume that we want to maintain a directory containing the office address and phone number for each person in the Rice Computer Science Department. In Java, each entry in such a directory has a natural representation as an object with three fields containing the person’s name, address, and phone number represented as character strings. We defer discussion about how to represent the directory itself until Section 1.6.

Each entry object must include operations to retrieve the name, address, and phone number fields, respectively.

Let’s summarize the form of a directory entry as a table:

         Fields:
             String name;
             String address;
             String phone;
         Methods:
             String getName();
             String getAddress();
             String getPhone();

This tabular description is not legal Java syntax. We will introduce the actual syntactic details in the next section of this chapter.

The three methods getName, getAddress, and getPhone do not take any explicit arguments because they are invoked by sending a “method call” to an object, called the receiver, which serves as an implicit argument for the method. In Java, the code defining the method can refer to this implicit argument using the keyword this, which is reserved for this purpose.

The syntax for invoking the method m (with no arguments) on the object o is

     m.o()

Consider the following example: assume that a Java program can access an Entry object e and needs to get the value of the name field of this object. The method invocation

     e.getName()

returns the desired result.

Finger Exercise 1.2.1: In the DrJava programming environment, load the program Entry.java, compile it, and type the following statements in the Interactions pane:

     Entry e = new Entry("Corky","DH 3104","x 6042");
     e.getName()

The first line defines a variable e as an Entry object with name field "Corky", address field "DH 3104", and phone field "x 6042". The second line is an expression that computes the name field of e. What value does the Java evaluator return for the name field of e?

Before we explain the Java code in the program Entry.java defining the methods getName, getAddress, and getPhone, we must discuss the syntax and primitive operations of Java.

1.3 Java Mechanics

1.3.1 Notation and Syntax

In Scheme, programs are constructed from expressions. In contrast, Java programs are constructed from statements. Statements are program phrases that do not have values, similar to Scheme expressions like (define ...) that return the dummy value (void). Nevertheless, many Java statements contain embedded expressions, so let us look briefly at Java expressions.

1.3.2 Java Expressions

In Java, arithmetic, boolean, and String expressions are written in conventional mathematical infix notation, adapted to the standard computer character set (called ASCII). For example, the Scheme expression

     (and (< (+ (* x x) (* y y)) 25) (> x 0))

is written in Java as

     (x*x + y*y > 25) && (x > 0)

The syntax of Java expressions is patterned after the C programming language. Like C, Java uses the symbol && for the “and” operation on boolean values (true and false) and the symbol == for the equality operation on numbers. (Warning: the symbols & and = are used in C and Java for other purposes.)

The following table lists the major infix operators provided by Java:

+ addition and String concatenation
- subtraction
* multiplication
/ division
% mod (remainder from integer division)
< less than
<= less than or equal
> greater than
>= greater than or equal
== equal
!= not equal
&& and

      || or

The Java arithmetic operators all perform the indicated operations using computer arithmetic instead of genuine arithmetic. Computer arithmetic does not exactly conform to the standard mathematical conventions. Calculations involving real numbers (Java types float and double) are approximate; the computer rounds the true result to the nearest real number expressible using the number of digits provided in the standard machine representation (scientific notation with a fixed number of digits for the fraction and exponent). Integer calculations are done exactly provided that the answer is an integer and that it can be represented using 31 binary digits plus a sign.^{1} Note that integer division always produces integer answers (unless you try to divide by zero which is an error). For example, the expression

     5/3

produces the result

     1

which is the quotient of 5 divided by 3. Integer division truncates the true rational result, dropping the digits to the right of the decimal point. Similarly, The expression

     5%3

produces the result

     2

which is the remainder of 5 divided by 3. In Java program text, spaces between symbols are ignored; the expression

1As we will explain shortly, Java supports several different sizes of integer representation; 31 binary digits plus sign is the default for integer constants.

     5 / 3

is equivalent to the expression

     5/3

Finger Exercise 1.3.2.1: In the DrJava programming environment, try evaluating the following expressions:

     5/3
     5 % 3
     5./3.
     5 / 0
     5./0.
     5 < 6
     5. < 6.
     3 + .1 * .1 - 3.

Java expressions directly in the Interactions pane. Did you get the answers that you expected?

All of the binary infix operators in Java are either arithmetic, relational, or boolean except for + when it is used in conjunction with strings. If either argument to + is of String type, then Java converts the other argument to a String. Object values are coerced to type String using their toString() methods. As we explain in Section 1.3.6, every object has a toString() method. The concatenation operator converts primitive values to strings using built-in conversion routines that we will discuss later.

Note that the order in which arguments appear and the use of parentheses in mixed integer and string expressions constructed from int and String values affects the conversion process. For example, the expression

     9 + 5 + 1 + "S"

evaluates to the String "15S" while the expression

     9 + (5 + (1 + "S"))

evaluates to the String "951S". The association rules for Java expressions are explained in Section 1.3.3.

Java also supports the unary prefix operators - (arithmetic negation) and ! (boolean “not”) used in conventional mathematical notation. Parentheses are used to indicate how expressions should be decomposed into subexpressions.

Finger Exercise 1.3.2.2: If the DrJava Interactions pane, try evaluating the following expressions:

     -5 + 3
     -(5 + 3)
     ! (5 < 6)

The only pure expression form in Java that deviates from conventional mathematical notation is the conditional expression notation

test ? consequent : alternative

borrowed from C. This expression returns the value of consequent if test is true and the value of alternative if test is false. It corresponds to the Scheme expression

     (cond [test consequent] [else alternative])

Note that when test is true, alternative is not evaluated. Similarly, when test is false, consequent is not evaluated. Hence, the expression

     (2 < 0) ?  2/(1 - 1) :  0

does not divide 2 by 0. The test expression must be a boolean value, true or

false.

Finger Exercise 1.3.2.3: In the DrJava Interactions pane, try evaluating the following expressions:

     (2 < 0) ? 2/(1 - 1) : 0
     (0 < 1) ? "foo" : "bar"
     17 ? true : false

The last example produces a syntax error because 17 is not a boolean value.

1.3.3 Precedence of Operations

Since Java uses conventional infix notation for expressions it relies on the notion of precedence to determine how expressions like

     12 * 5 + 10

should be interpreted. The Java operations given in the preceding subsection are divided into the following precedence groups:

prefix operators - !
multiplicative  * / %
additive  + -
relational  < > >= <=
equality  == !=
logical and  &&
logical or  ||
conditional  ? ... :

from highest to lowest. A higher precedence operator has greater “binding power”. For example, the expression

     72.  - 32.  * 1.8

is equivalent to

     72.  - (32.  * 1.8)

because * has higher precedence than infix -.

Finger Exercise 1.3.3.1: In the DrJava Interactions pane, try evaluating the following expressions:

     72. - 32. * 1.8
     (72. - 32.) * 1.8

All of infix operators listed above are left-associative: when infix operators of equal precedence are chained together, the leftmost operator has precedence. For example,

     72.  - 30.  - 12.

is equivalent to

     (72.  - 30.)  - 12.

Parentheses can be used to override the built-in precedence and associativity of operators. Hence,

     (72.  - 32.)  * 1.8

equals 40*1.8. Similarly,

     72.  - (30.  - 12.)

equals

     72.  - 18.

It is a good idea to use parentheses if you have any doubts about the precedence relationship between consecutive operators. The judicious use of parentheses makes complex expressions easier to read.

Finger Exercise 1.3.3.2: In the DrJava Interactions pane, try evaluating the following expressions:

     72. - 30. - 12.
     72. - (30. - 12.)

1.3.4 Java Statements

Since the Java expression language is not very rich, Java programs express computations as sequences of statements rather than as compound expressions. The most common form of Java statement is an assignment statement

type var = expr ;

where type is a Java type name, var is a Java variable name, and expr is an expression of type compatible with the type of var. The assignment statement

     int x = 5;

asserts that “the variable x has value 5”.

Finger Exercise 1.3.4.1: In the DrJava Interactions pane, try evaluating the following statements and expressions:

     int x = 5;
     x*x
     double d = .000001;
     double dd = d*d;
     dd
     dd*dd
     1. + dd
     1. + dd*dd

Did you get the answers that you expected?

Java variable names and type names must be identifiers. An identifier is any sequence of “alphanumeric characters” (letters, digits, and ) beginning with a letter or —except for the following keywords, which are reserved and may not be used as variable names or type names:

     abstract    default     if          private     throw
     boolean     do          implements  protected   throws
     break       double      import      public      transient
     byte        else        instanceof  return      try
     case        extends     int         short       void
     catch       final       interface   static      volatile
     char        finally     long        super       while
     class       float       native      switch
     const       for         new         synchronized
     continue    goto        package     this

Java is case-sensitive; the variable X is distinct from the variable x. There are three kinds of variables in Java: fields, method parameters, and local variables. Fields and method parameters are discussed in detail in the next subsection. We will defer a discussion of local variables until Section 1.12.1.

Java includes all of the basic statement forms found in the C programming language expressed in essentially the same syntax. In the remainder of this monograph, we will introduce these statement forms as they are needed. Although Java accepts most C syntax, many common C constructions are considered bad style in Java.

1.3.5 The Structure of Java Programs

Every Java program consists of a collection of classes—nothing else. A class is a template for creating a particular form of object. A Java class definition corresponds to a Scheme struct definition generalized to include all of procedures that process objects of the defined class. In Java, all program code is attached to some class.

Each object created by a class template contains the same members, each of which is either a field or a method. A field is a “container” that holds a value. A method is an operation on the fields of the object and any values that are passed as arguments to the method. The objects created by a particular class template are called the instances or objects of that class. Each instance contains the members specified in the class template.

Each member of a class has a name consisting of a Java identifier (defined above in Section 1.3.4). For now, we will require all such names to be unique within a class. We will slightly relax this restriction when we discuss overloading in Section 1.12.4.

Java objects are a generalization of Scheme structs. In Java, the collection of operations (methods) that operate an object is determined by the class definition for the object. In Scheme, the set of operations for manipulating a struct is fixed: a constructor, a recognizer, field-accessors, and field-modifiers.

The Java program in Figure 1 defines a class Entry suitable for representing entries in the department directory application described in Section 1.2:

     class Entry {
       /* fields */
       String name;
       String address;
       String phone;
       /* constructor */

       Entry(String n, String a, String p) {
         this.name = n;
         this.address = a;
         this.phone = p;
       }
       /* accessors */
       String getName() { return this.name; }
       String getAddress() { return this.address; }
       String getPhone() { return this.phone; }
     }

Figure 1: The Entry class

This class is the Java analog of the Scheme structure definition

     (define-struct Entry (name address phone))

where the name, address and phone fields are always bound to strings.

Let’s examine the syntax of the Entry class definition in detail. It consists of seven members:

• three fields name and address and phone which must contain String values;

• a constructor that specifies how these fields are initialized when an Entry object is constructed; and

• three methods getName, getAddress, and getPhone that define accessors for extracting the corresponding fields from an Entry.

An instance (object) of the class Entry is created by an expression of the form

     new Entry("SomeName","SomeAddress","SomePhone")

The expression is equivalent to the Scheme expression

     (make-Entry "SomeName" "SomeAddress" "SomePhone")

which creates a Scheme Entry object. The three accessor methods getName, getAddress, and getPhone are equivalent to the Scheme accessors Entry-name, Entry-address and Entry-phone that are automatically generated by the Scheme

define-struct definition.

Recall the syntax for method invocation in Java is quite different from the syntax for functional application in Scheme. Given an instance e of class Entry, the expression

     e.getName()

invokes the getName method of object e. The equivalent Scheme expression is written

     (Entry-name e)

The three methods defined in class Entry are extremely simple, yet they illustrate the most important characteristic of object-oriented programming: operations are attached to the data objects that they process. The methods getName, getAddress, and getPhone take no arguments yet they have access to the fields of the Entry object to which they are attached. The Entry object e in the method call

     e.getName()

is called the receiver because it “receives” the method call. (In the Java Language Specification, the term target is used instead of receiver.)

In Java, a method cannot return a value without executing a statement of the form

     return e;

where e is some expression. Since Java is statement-oriented language, the body of a method is a sequence of statements rather than an expression. When execution reaches a return statement, the specified expression is evaluated and the resulting value is returned as the value of the method call. Hence, the body of each of the three methods in the Entry class is a return statment.

In the code for the Entry class, the constructor and accessor methods all refer to fields of this, the hidden parameter bound to the object that is the receiver of the method invocation. For example, the expression

     this.name

returns the value of the name field of the object this. In constructor invocations, this is bound to the newly allocated object.

One attractive feature of Java is that the method syntax mandates a contract (declaration of input types and output type) as part of the method header. For example, the method header

     String getName()

indicates that the getName method takes no arguments (other than the “receiver” as an implicit argument this and returns a String value. In subsequent examples, we will define methods that take explicit arguments; the type of each such argument must be declared in the method header.

Finger Exercise 1.3.5.1: In the Definitions pane of DrJava, load the Java program Entry.java and compile the program. In the DrJava Interactions pane, try evaluating the following program text:

     Entry zung =  new Entry("Zung", "DH 3098", "x 3835");
     Entry cork =  new Entry("Cork", "DH 3104", "x 6042");
     zung.getName()
     cork.getName()
     cork.getPhone()
     zung == cork
     zung != cork

Did you get the results that you expected? If not, ask a teaching assistant to explain any of the results that you do not understand.

We will explain the syntax of Java class definitions in more detail in Section 1.5.

1.3.6 Printing Object Values

Given the definition of the Entry class given above, we could evaluate the expression

     new Entry("Corky","DH 3104","x 6042")

in the DrJava Interactions pane but the result would not be very informative. When the Java evaluator needs to convert an object to a printable String representation, it uses a method called toString() which is defined in class Object. Since every class is a subclass of Object, every class includes the

toString() method.

Every class inherits a definition of the toString() method from its superclass. The ultimate superclass Object contains a simple definition for toString() that is not very informative: it returns the String consisting of the class name for this followed by an @ followed by the address in memory where the object this is located.

Each class below Object in the superclass hierarchy either relies on the toString() method inherited from its superclass or introduces a new definition for toString() that overrides the definition in its superclass.

Finger Exercise 1.3.6.1 Load your file Entry.java into DrJava and compile it. Then evaluate

     new Entry("Corky","DH 3104","x 6042")
     new Entry("Corky","DH 3104","x 6042")

Did you get the results that you expected? Note that each new operation creates a distinct object.

The Entry class is an immediate subclass of the Object class which defines the toString() method. This definition of toString() simply generates the String consisting of the name of the class concatenated with an sign and an identifying serial number for the object.

To produce a better printed form for the instances of a class, we must define a toString() method specifically for that class. In the case of the Entry class, we could define toString() as follows:

     public String toString() {
       return "Entry[" + this.name + ", " + this.address + ", "
              + phone + "]";
     }

Java requires the public attribute in the header of the toString() method because it overrides an inherited method that is public. In Java, every class member has an associated visibility attribute. When an inherited method is overridden, its visibility cannot be reduced. We will discuss this issue in more detail in Section 1.7.7.

Finger Exercise 1.3.6.2 Load your saved program Entry.java into the Definitions pane of DrJava. Add the definition immediately above to the Entry class and compile the program. In the Interactions pane, evaluate the same print statement as in the last finger exercise. Is this output more helpful?

Java Design Rule: Redefine (override) the toString() method for every class that is used to represent data. (Recall that classes can also be used to group static fields and methods.)

1.3.7 Defining Constants

In addition to (instance) members (shown in the Entry class above), a Java class can include static members that are attached to the class rather than instances of the class. The static members of a class are not included in the template used to create class instances. There is only one copy of a static field for an entire class—regardless of how many instances of the class are created (possibly none). The most important use of static fields is to hold program constants. We will discuss static methods later in this monograph.

In Java, each program constant is attached to an appropriate class as a static final field of that class. The final attribute indicates that the field is a constant. For example, the built-in Java class Integer includes the constant

static final fields int MIN VALUE and int MAX VALUE specifying the minimum

and maximum values of int type.

The following Java class defines two constants: INCHES PER METER and METERS PER INCH:

class Conversions {
  static final double INCHES_PER_METER = 39.37;
  static final double METERS_PER_INCH = .0254;
}

Outside of the Conversions class, these two contants are denoted

     Conversions.INCHES PER METER

and

     Conversions.METERS PER INCH

respectively. Inside the Conversions class, the Conversions. prefix is unnecessary.

Finger Exercise 1.3.7.1: In the Definitions pane of DrJava, enter the Conversions class given above. Define a class Person with two fields name and

height and three methods getName, getHeightInMeters, and getHeightInInches.

Write a TestPerson class corresponding to Person that creates several different instances of Person and applies all three methods to them. Save your program in a file called Person.java.

1.3.8 Capitalization and Commenting Conventions

By convention, Java programs are written entirely in lower case characters with three exceptions.

• The first letter of class names are capitalized to distinguish class names from member names.

• The names of constant static final fields are written entirely capital letters. For example, the static final fields in the preceding subsection are all capitalized according to this convention.

• The first letter in each word of a multi-word identifier after the first is capitalized. For example, the built-in Java class Object includes a method called toString() that we will discuss later. The capital S signifies the beginning of a word within the multi-word name toString().

These conventions are not enforced by Java compilers, but it is considered bad style to violate them. A related convention is to never use the special character $ in a name; this character is reserved for the use of the Java compiler. Unfortunately, most Java compilers do not enforce this convention.

Java relies on commenting conventions similar to those in C++. A comment that is confined to a single line begins with the character sequence // and ends at the end of the line. Longer comments must be enclosed between the opening “bracket” /* and “closing” bracket */. Examples of both form of comments appear in Section 1.6. Note that a bracketed comment can appear in the middle of a line of code.

Finger Exercise 1.3.8.1: add both forms of comment to the Conversions and Person classes from the preceding exercise.

1.4 Java Data Types

Java programs manipulate two fundamentally different kinds of values: primitive values and object values.

1.4.1 Primitive Types

All primitive values belong to one of eight primitive types: int, float, boolean, char, byte, short, long, and double. Four of these types designate different sizes of bounded integers:

• byte contains the integers ranging from -128 to 127;

• short contains the integers ranging from -32768 to 32767;

• int contains the integers ranging from -2147483648 to 2147483647; and

• long contains the integers ranging from

          -9223372036854775808

to

          9223372036854775807.

In practice, only three of these primitive types are widely used: int, boolean,

and double.

A program should never contain explicit references to any of the specific integer constants given above. They can always be replaced by references to the static final fields MIN VALUE and MAX VALUE in the corresponding “wrapper” class: Byte, Short, Integer or Long. Java has a built-in wrapper class for each primitive type. For example, in the wrapper class Integer corre-

sponding to the primitive type int, MIN VALUE = -2147483648 and MAX VALUE

= 2147483647. For more information on the primitive wrapper classes, read the on-line documentation from Sun Microsystems for the built-in class Number and its subclasses.

The boolean type has two values true and false. The char type supports the Unicode character set which includes all conventional ASCII characters plus almost any foreign character imaginable. The char type is rarely used in Java programs because a flexible String object type is built-in to the language. The remaining two types float and double are used for approximate computations involving real numbers; they denote standard IEEE 32-bit and 64-bit formats, respectively.

Numeric Constants. Java interprets unadorned integer constants as values of type int. Long integer constants are indicated by attaching the suffix L to the number. For example, the constant 9223372036854775807L can be used in Java program text, while the same constant without the L is an error because it is too big to be an int. The double type is the default type for any floating point constant. On modern machines, there is little reason to use the less precise float.

Conversions Between Types [Optional]. Java will automatically convert any numeric type to a more “general” numeric type demanded by context. The following list gives the primitive numeric types in increasing order of generality:

     byte → short → int → long → float → double

Note that the notion of generality here is imperfect in some situations. The conversion of a long to a float, for example, will often lose precision. In fact, even the conversion of a really large long value to a double can lose precision.

Java provides explicit conversion operators called casts to convert a numeric type to a less general type. A cast is simply a type name enclosed in parentheses used a prefix operator. For example, the following expression casts the int constant 127 to the type byte

     (byte) 127

When converting from one bounded integer type to another, Java silently truncates leading digits if the output type is shorter than the input type. Watch out!

Finger Exercise 1.4.1.1: In DrJava, convert the maximum long value to double. (Do not type the 20+ digits for this constant! Recall the contant fields in the wrapper class Long.) What do you get? Convert the maximum long value to type float. What do you get? Why? Try converting the maximum long value minus 1 to double and back again. Do you get the same result?

1.4.2 Object Types

Object values are created by instantiating classes, which may either be builtin or program-defined. Classes are organized in a strict hierarchy with the special built-in class Object at the top. Every class C except Object has a unique parent in the hierarchy called the superclass of C. A descendant in the class hierarchy is called a subclass. Each subclass of a class C includes all of the members (fields and methods) of C and possibly additional members. For this reason, we say that each immediate subclass of (child of) C extends C. Note that subclass relation is transitive and reflexive. If class A is a subclass of class B and class B is a subclass of class C then A is a subclass of C (transitivity). In addition, every class A is a subclass of itself (reflexivity).

We will reserve the term extends to describe the immediate subclass (child) relation: A extends B iff A is an immediate subclass of B. Hence, the extends relation is neither transitive nor reflexive. Since Object is the top class of the hierarchy, all classes are subclasses of Object.

For example, the built-in classes Integer and Float extend the built-in class Number which extends Object. Hence, the superclass of Integer is Number, the superclass of Float is Number, and the superclass of Number is Object.

Object values are actually references to objects. For this reason, two different fields can be bound to exactly the same object. In Java, objects are never implicitly copied. When a field or method parameter v is bound to an object o, the value of v is a reference to the object o, not a copy of o! Scheme follows exactly the same conventions with regard to copying program data.

Every Java class C has an associated type C consisting of all instances of class C and all of its subclasses. Hence, the type Object contains all object values. The built-in class String has the class Object as its superclass. Since the class String has no subclasses, the only values of type String are instances of the class String. In contrast, the built-in class Number is a child of class Object and has several subclasses including Integer and Float. Hence, all instances of the classes Integer and Float are values of type Number.

In Java, every field and method has a declared type given as part of its definition. For a method, the declared type includes the type of the result and the types of the parameters. If a method does not return a value, then its declared return type is the degenerate type called void. There are no values of type void. A method with void return type does not return a value.

Java determines the type of every program expression using a simple set of rules and confirms that

• the value assigned to a field is consistent with the field’s declared type;

• the values passed as arguments to a method are consistent with the

corresponding parameter types;

• the value returned by a method is consistent with the declared return type of the method; and

• the member name in a field reference or method invocation is compatible with the static type of the receiver.

We will discuss these “type-checking” rules in more detail in Section 1.12.2.

1.5 Java Class Definitions

In Scheme, new forms of data are defined by the define-struct construct. A define-struct expression defines a template for creating instances of the specified structure. In Java, new forms of data are defined by the class construct. A class C serves as a template for creating instances of the class and as a basis for defining new classes that extend (enlarge or modify) C. The precise differences between the Scheme define-struct and the Java class constructs are explained below.

1.5.1 Defining Classes to Represent Compound Data

In Section 1.2, we introduced a class definition for department directory entries and briefly explained its various parts. We will now review the syntactic structure and semantic behavior of classes in more detail.

Recall that a directory entry is a object with the following fields and methods:

     Fields:
         String name;
         String address;
         String phone;
     Methods:
         String getName();
         String getAddress();
         String getPhone();

In Java, we can express the data definition by the class

     class Entry {
       /* fields */
       String name;

       String address;
       String phone;
       /* accessors */
       String getName() { return this.name; }
       String getAddress() { return this.address; }
       String getPhone() { return this.phone; }
     }

But this class definition is not complete because it does not define a suitable constructor for creating and initializing objects of this form.

1.5.2 Constructors

If we use define-struct to define a new form of Scheme data named Entry with three fields, Scheme generates a procedure make-Entry of three arguments that constructs a new instance of the structure containing the field values specified by the arguments. Java provides a similar mechanism called new for creating new instances of a class, but the class must provide a special method called a constructor that specifies how the fields are initialized. A constructor method has the same name as the class and does not contain the return type in the heading. The constructor for our sample class Entry has the following form:

       Entry(String n, String a, String p) {
         this.name = n;
         this.address = a;
         this.phone = p;
       }

Like a method with void return type, a constructor does not contain a return statement.

If the preceding constructor is included in the definition of class Entry, then the expression

       new Entry("Corky", "DH 3104", "x 6042")

constructs an instance of class Entry with the name "Corky", the address "DH 3104", and the phone "x 6042". The value of the new expression is the created object.

In general, a class may have several different constructors, but each constructor must have a distinct list of argument types so that the Java compiler can determine which one is meant. If a class does not include a constructor, Java generates a default constructor of no arguments that initializes each field of the object to the “null” value of its declared type. The “null” value for any class type is the value null, which refers to nothing. The “null” value of each primitive type is the “zero” value for that type: 0 for the six numeric primitive types, false for boolean, and the null character (which has code 0) for char. Any attempt to invoke a method on the null reference is a run-time error, which terminates program execution.^{2}

Let us summarize syntax for constructors. A constructor definition for a class C begins with name of the class followed by a (possibly empty) list of parameters separated by commas and enclosed in parentheses. The remainder of the constructor definition is a (possibly empty) sequence of Java statements enclosed in braces. When a new instance of class C is created, the body of the corresponding constructor is executed to initialize the fields of the created object. New objects are created by new expressions that have the form

     new C(E1, ..., En)

E1, ..., En are expressions with types matching the parameter declarations for some constructor of class C. When a new expression is evaluated, Java creates a new instance of class C and initializes the fields of C by binding the parameters of the matching constructor to the values of the argument expressions and executing the statements in the body of the constructor.

Finger Exercise 1.5.2.1: Add a second constructor to your saved Entry class. Your second constructor should take one argument, the value of String name, and initialize the other fields to the String "unknown". Test your code.

1.5.3 Defining Instance Methods

The three instance methods getName, getAddress and getPhone in class Entry all simply return the value of a field from the object that received the method call.

Let us define another method for class Entry that does more computation. Assume that we want to define an operation match for the class Entry that takes a string argument keyName and determines whether it matches the name field of the specified Entry object. We could include such a method definition within the definition of class Entry as shown in Figure 2:

     class Entry {
       String name;
       String address;

2Invoking a method on the null reference actually throws an exception which in principle can be caught using Java’s exception handling mechanism discussed in Section. Since this exception is the result of a coding error, catching such an exception in regular program code (as opposed to a test harness) is a bad idea.

       String phone;
       /* constructor */
       Entry(String n, String a, String p) {
         this.name = n;
         this.address = a;
         this.phone = p;
       }
       /* accessors */
       String getName() { return this.name; }
       String getAddress() { return this.address; }
       String getPhone() { return this.phone; }
       /* other methods */
       /** determines if this matches keyName */
       boolean match(String keyName) {
         return this.name.equals(keyName);
       }
     }

Figure 2: The expanded Entry class

The match method is implemented using the equals method on the String field name. Recall that the String class is built-in to Java. The equals method from the String class takes an argument and returns true if (and only if) it is a String with exactly the same contents as the receiver String object. Hence,

     (new Entry("Corky","DH 3104","x 6042")) . match("Corky")

returns

     true,

while

     (new Entry("Corky","DH 3104","x 6042")) . match("Matthias")

returns

     false.

Warning The Java infix operator == can be used to compare objects for identity. In many contexts, the results of such a comparison are problematic. On objects, the == operator returns true if (and only if) the both arguments are exactly the same object. Hence, if x is a variable of some object type T, the expression

     x == x

returns

     true

For distinct object arguments, the == operator returns false. Hence,

        new Entry("Corky","DH 3104","x 6042")
     == new Entry("Corky","DH 3104","x 6042")

returns

     false

because each occurrence of new creates a distinct object. For most Java object types including String, the == operator is not a reliable mechanism for testing equality! For example, Java does not guarantee that it creates only one copy of a String constant.

Finger Exercise 1.5.3.1

1. Add the match method to the Entry class in the file Entry.java. Test your code.

2. Modify your match method to use the == operator instead of the equals method. Find some test cases where it fails! Hint: A String constant that appears in the DrJava interactions pane will be distinct from any String constant defined in a class in the Definitions pane.

Java Design Rule: There are only two valid uses of the == operator:

• to compare values of primitive type; and

• to test object identity (not equality!).

The second use is relatively uncommon.

1.6 The Union and Composite Patterns

In our department directory example, an object of type Entry only has one form, namely an instance of class Entry. If we were designing the data for a more comprehensive directory such as a city phone directory, we would need more than one form of entry. At a minimum, we would need entry formats suitable for business listings, government listings, and residential listings. For such a phone directory, we might define an entry as follows.

A CityEntry is either:

   • a BusinessEntry(name,addr,phone,city,state),
   • a GovernmentEntry(name,addr,phone,city,state,gov), or
   • a ResidentialEntry(name,addr,phone),

where name is a string specifying the name for the listing, addr is a string specifying the street address for the listing, phone is a string specifying the phone number (with area code) for the listing, city and state are strings specifying the city and state for the listing, and gov is a string specifying the government entity for that the listing, e.g. the ”City of Houston”.

The BusinessEntry and GovernmentEntry forms include city and state information because businesses and government agencies that serve clients in cities outside their local calling area often elect to have their phone numbers included in the directories of other cities (in addition to the cities where they are located). In addition, government listings include a string specifying the government entity to which they belong. For example, a listing for the Federal Bureau of Investigation would specify the ”U.S. Government” as the gov field.

In Scheme, we would represent such an entry data type by defining three different structures:

     ;; a CityEntry is either:
     ;;   a BusinessEntry
     ;;     (make-BusinessEntry name addr phone city state),
     ;;   or a GovernmentEntry
     ;;     (make-GovernmentEntry name addr phone city state gov),
     ;;   or a ResidentialEntry
     ;;     (make-ResidentialEntry name addr phone).
     (define-struct BusinessEntry (name addr phone city state))
     (define-struct GovernmentEntry (name addr phone city state gov))
     (define-struct ResidentialEntry (name addr phone))

Note that the type CityEntry consisting of the union of the types BusinessEntry,

GovernmentEntry, and ResidentialEntry is not defined in program text be-

cause all union types in Scheme are implicit.

In Java, we can define the CityEntry type by introducing a “dummy” CityEntry class that we extend by “concrete” classes^{3} for each different form of Entry data. This technique, which is widely used in object-oriented programming, is called the union pattern. In this pattern, an abstract class serves as the root of a hierarchy of subclasses called variants, which are the component types of the union. In this example, there are three variant classes:

3Any class that is not declared as abstract is “concrete”.

CHAPTER 1. FROM SCHEME TO JAVA 24

BusinessEntry, GovernmentEntry, ResidentialEntry. The following Java code

defines the city-entry type:

/**  a CityEntry is either:
*   (i)   a BusinessEntry
*         new BusinessEntry(name,addr,phone,city,state),
*   (ii)  a GovernmentEntry
*         new GovernmentEntry(name,addr,phone,city,state,gov), or
*   (iii) a ResidentialEntry
*         new ResidentialEntry(name,addr,phone).
*/
abstract class CityEntry {
}
class BusinessEntry extends CityEntry {
/* fields */
String name;
String address;
String phone;
String city;
String state;
/* constructor */
BusinessEntry(String n, String a, String p, String c, String s) {
  this.name = n;
  this.address = a;
  this.phone = p;
  this.city = c;
  this.state = s;
}
/* accessors */
String getName() { return this.name; }
String getAddress() { return this.address; }
String getPhone() { return this.phone; }
String getCity() { return this.city; }
String getState() { return this.state; }
}
class GovernmentEntry extends CityEntry {
/* fields */
String name;

String address;
String phone;
String city;
String state;
String government;
/* constructor */
GovernmentEntry(String n, String a, String p, String c, String s, String g) {
  this.name = n;
  this.address = a;
  this.phone = p;
  this.city = c;
  this.state = s;
  this.government = g;
}
/* accessors */
String getName() { return this.name; }
String getAddress() { return this.address; }
String getPhone() { return this.phone; }
String getCity() { return this.city; }
String getState() { return this.state; }
String getGovernment() { return this.government; }
}
class ResidentialEntry extends CityEntry {
String name;
String address;
String phone;
/* constructor */
ResidentialEntry(String n, String a, String p) {
  this.name = n;
  this.address = a;
  this.phone = p;
}
/* accessors */
String getName() { return this.name; }
String getAddress() { return this.address; }
String getPhone() { return this.phone; }
}

Figure 3: The City Directory union class hierarchy

Note that each concrete class includes exactly the same fields as the corresponding Scheme structure definition. The pivotal differences between the Java code and the corresponding Scheme code are (i) the presence of the abstract class CityEntry in the Java Code identifying the union type which is left implicit in Scheme and (ii) the explicit definitions of constructors and accessors (selectors) in the Java concrete classes which are automatically generated by the Scheme structure definitions.

The Java code in the CityEntry example above involves several concepts that we have not discussed before.

• The attribute abstract attached to the class CityEntry indicates that CityEntry is a “dummy” class that cannot be instantiated. The class CityEntry is defined solely to group the concrete classes BusinessEntry,

     GovernmentEntry, and ResidentialEntry as a type that includes all three

kinds of data.

   • The concrete classes BusinessEntry, GovernmentEntry, and ResidentialEntry

are the only subclasses of the class CityEntry. Hence, the only values of type CityEntry are the instances of the classes BusinessEntry,

     GovernmentEntry, and ResidentialEntry.

• Each concrete class includes a constructor definition that specifies how new instances of the class are initialized. Each concrete class also includes getter (accessor) methods (like getPhone) to get the values of the various fields of the class.

The following expression creates a BusinessEntry for Rice University

     new BusinessEntry("Rice University", "6100 Main Street", "713-348-8101",
                       "Houston", "Texas")

This syntax is wordy but straightforward. Don’t forget to include the keyword new at the front on each constructor invocation!

Finger Exercise 1.6.1 Enter the preceding class definitions into the Definitions pane of DrJava. Compile this program and evaluate the following expressions in the Interactions pane:

     BusinessEntry e1 =
       new BusinessEntry("Rice University", "6100 Main St.", "713-527-8101",
                         "Houston", "TX");
     ResidentialEntry e2 =
       new ResidentialEntry("Robert Cartwright", "3310 Underwood St.",

                            "713-660-0967");
     e1.getName()
     e2.getName()

Did you get the results that you expected?

1.6.1 Member Hoisting

The preceding Java program can be improved by eliminating duplicated code in the subclasses extending CityEntry. The concrete classes forming a union are called variants. Note that the fields name, address, and phone appear in all three variants of the abstract class CityEntry. So do the definitions of the corresponding accessors getName, getAddress, and getPhone. These repeated member definitions can be hoisted into the abstract class CityEntry yielding the following Java code:

     abstract class CityEntry {
       String name;
       String address;
       String phone;
       /* accessors */
       String getName() { return this.name; }
       String getAddress() { return this.address; }
       String getPhone() { return this.phone; }
     }
     class BusinessEntry extends CityEntry {
       String city;
       String state;
       /* constructor */
       BusinessEntry(String n, String a, String p, String c, String s) {
         this.name = n;
         this.address = a;
         this.phone = p;
         this.city = c;
         this.state = s;
       }
       /* accessors */
       String getCity() { return this.city; }

String getState() { return this.state; }
}
class GovernmentEntry extends CityEntry {
/* fields */
String city;
String state;
String government;
/* constructor */
GovernmentEntry(String n, String a, String p, String c, String s, String g) {
  this.name = n;
  this.address = a;
  this.phone = p;
  this.city = c;
  this.state = s;
  this.government = g;
}
/* accessors */
String getCity() { return this.city; }
String getState() { return this.state; }
String getGovernment() { return this.government; }
}
class ResidentialEntry extends CityEntry {
/* constructor */
ResidentialEntry(String n, String a, String p) {
  this.name = n;
  this.address = a;
  this.phone = p;
}
}

Finger Exercise 1.6.1.1 In the preceding code, the abstract class CityEntry

has three concrete subclasses: ResidentialEntry, BusinessEntry, and GovernmentEntry.

By applying some very simple program transformations, you can eliminate more duplicate code in the CityEntry class and subclasses by inserting a new abstract class NonResidentialEntry between CityEntry and the concrete classes BusinessEntry and GovernmentEntry hoisting the common members of these concrete classes. After this addition, the class CityEntry still has only three concrete subclasses but only one of them is an immediate subclass. The other immediate subclass is NonResidentialEntry. Test your code using DrJava.

[Optional] Finger Exercise 1.6.1.2 Note that the constructor for each concrete subclass of CityEntry replicates the code for initializing the fields address and phone defined in the abstract class CityEntry. Similarly, the constructor for each concrete subclass of NonResidentialEntry replicates code for initializing the fields city and state. This code replication can be eliminated by defining constructors for the abstract classes CityEntry and NonResidentialEntry and using the special method name super to invoke these constructors at the beginning of each concrete class constructor.

In the body of a constructor for a class C, the reserved word super can be used as a method name to invoke a constructor in the superclass of C. In such an invocation, the method name super must be followed by an appropriate argument list enclosed in parentheses just like any other method call. (For more information on super calls, consult a Java reference book such as Thinking in Java by Eckel, The Java Programming Language by Arnold and Gosling, or Java in a Nutshell by Flanagan.)

Eliminate constructor code replication in the CityEntry class hierarchy. Test your code using DrJava.

Member hoisting is a special form of the general concept of code factoring. Code factoring is any transformation that eliminates repeated code. In functional languages like Scheme, the standard code factoring is typically accomplished by introducing a new λ-abstraction with a repeated code pattern as its body. Each instance of the repeated pattern is replaced by an appropriate call on the new abstraction. In some cases, the arguments to the pattern are procedures. This form of code factoring can be implemented in Java in several different ways. If all of the code repetitions appear within a class hierarchy for which the programmer has control of the source, then a method can be introduced in the most restrictive subclass that includes all of the occurrences of the repeated pattern. Each occurrence can be replaced by an invocation of the introduced method. In some cases, the arguments to the method are command objects (discussed in Section 1.11) representing procedures.

Java Design Rule: never repeat code in the variant classes in a union type. Hoist any repeated code into methods defined in an abstract superclass. If necessary, create a new abstract superclass to contain the factored code.^{4} while variants B and C share code for method n. In this case, either the code for method m or the code for method n can be hoisted but not both. More complex factoring schemes are possible (using, for example, the command pattern discussed in Section 1.11), but they are typically not worth the complication.

4In pathological cases, some repeated code may not be subject to factoring because of conflicts among possible factorings. For example, variants A and B may share code for method m

1.6.2 The Composite Pattern

Let’s return to our department directory example and show how to use the union pattern to represent department directory data.

   A DeptDirectory is either:

• an Empty directory, or

• a composite directory Cons(e,d) where e is the first Entry of the directory and d is the rest of the directory.

In Scheme, we could implement this definition using the following collection of structures:

     ;; a DeptDirectory is either:
     ;;   the empty directory (make-Empty), or
     ;;   the non-empty directory (make-Cons Entry DeptDirectory)
     (define-struct Empty ())
     (define-struct Cons (first rest))
     ;; an Entry is (make-Entry String String String)
     (define-struct Entry (name address phone))

Note that the preceding Scheme code leaves most of the corresponding data definition unstated! It never mentions the new type DeptDirectory. It does not express the fact that instances of Empty and Cons are elements of the type DeptDirectory, nor does it state that the first and rest fields of a Cons must be of type Entry and DeptDirectory, respectively. Similarly, it fails to state that the fields name, address and phone are all strings. In Scheme, we must compose comments to communicate this information. Unfortunately, since these comments are not code they are not checked for consistency.

In Java, each new type of data is represented by a class. Since the DeptDirectory type has two variants, we must use the union pattern to represent this type. The following collection of class definitions relies on the union pattern to define the DeptDirectory type. Since the DeptDirectory type is implemented by an abstract class, we will prepend the name DeptDirectory with the letter A to indicate that the class is abstract.

     /**  an DeptDirectory is either:
      *   (i)  the empty directory new Empty(), or

      *   (ii) the non-empty directory new Cons(Entry,DeptDirectory)
      */
     abstract class DeptDirectory {}
     class Empty extends DeptDirectory {}
     class Cons extends DeptDirectory {
       Entry first;
       DeptDirectory rest;
       /* constructor */
       Cons(Entry f, DeptDirectory r) {
         this.first = f;
         this.rest = r;
       }
       /* accessors */
       Entry getFirst() { return this.first; }
       DeptDirectory getRest() { return this.rest; }
     }

The Java code is wordier, but it captures all of the information in the data definition. Our data definition comment is redundant. The class Empty contains no fields, just like the corresponding Scheme struct. The class Cons contains two fields first and rest akin to the two fields in the corresponding Scheme struct. Similarly, the Entry class contains three fields name, address, and phone just like the Scheme struct Entry given above. The abstract class DeptDirectory is extended by only two classes: Empty and Cons. Hence, DeptDirectory is the union of Empty and Cons.

The Java code in the DeptDirectory example relies on one new feature that we have not seen before, namely the notion of a default constructor. The class Empty is concrete but does not include a constructor to initialize its fields because there are no fields to initialize! Java automatically generates a default zero-ary constructor for any class definition that does not include a constructor. As a result, the expression

     new Empty()

generates a new instance of the class Empty.

When Unions are Composite The use of the union pattern in the DeptDirectory example has an extra feature not present in the preceding CityEntry example. One of the variants of the union class DeptDirectory includes a field of type DeptDirectory which makes the structure of the union class self-referential. Since self-referential structures are ubiquitous in Scheme (and other functional languages), this feature is not at all surprising to programmers familiar with functional programming. In the OOP (“object-oriented programming”) community, which has strong historical ties to imperative programming, this feature is viewed as distinctive because it implies that methods that process the union class are naturally recursive. For this reason, the OOP community assigns a distinct pattern name, namely composite pattern, to the the special case of the union pattern where self-reference is present in the data definition. We will use this terminology in the remainder of the monograph.

The following expression creates a DeptDirectory containing the address and phone information for Corky and Matthias:

     new Cons(new Entry("Corky","DH3104","x 6042"),
       new Cons(new Entry("Matthias","DH3106","x 5732"), new Empty()))

This syntax is wordy but straightforward. Don’t forget to include the keyword new at the front on each constructor invocation!

1.6.3 Defining Instance Methods for a Composite Class

In Section 1.5.1, we showed how to define simple (instance) methods for the individual class Entry. But we did not show how to express operations that process all of the different forms of data defined by a composite hierarchy. Since each different form of data in a composite hierarchy is represented by a distinct concrete class, we can write a separate method definition for each kind of data.

Consider the following example. Assume that we want to define a method

     String firstAddress(String name)

on DeptDirectory that returns the address for the first person in the directory if the directory is non-empty and the null reference null if it is empty. We can write separate definitions for the method firstAddress in the concrete classes Empty and Cons as follows:

     class Empty {
       ...
       String firstAddress() { return null; }
     }
     class Cons {
       ...
       String firstAddress() { return this.first.getAddress(); }
     }

Now assume that x is a variable of type DeptDirectory. If we try to invoke the method firstAddress on x, Java will reject the code as erroneous because the class DeptDirectory does not contain a method named firstAddress. How can we enlarge the definition of firstAddress so that it applies to the class

DeptDirectory?

The answer is that we declare the method firstAddress in the class DeptDirectory as an abstract method:

     abstract class DeptDirectory {
       ...
       /* firstAddress() returns the first address in a DeptDirectory;
            it returns null if the DeptDirectory is empty */
       abstract String firstAddress();
     }

An abstract method is a method without a body. Abstract methods can only appear in abstract classes. Any class containing an abstract method must be declared abstract because it cannot be instantiated. Every concrete class extending an abstract class must provide concrete definitions for the abstract methods it inherits. This rule guarantees that abstract methods are never attached to objects.

Let us illustrate the process of defining a method over a composite class hierarchy in more detail by defining a method

     String findAddress(String name)

on DeptDirectory that finds the address for the person identified by name, assuming that name is in the directory.

First, we must insert the following member somewhere in the class DeptDirectory

     /* findAddress(s) returns the address of the person with name s;
          it returns null if no matching entry is found */
     abstract String findAddress(String name);

The abstract modifier in the definition indicates that the definition only describes the input and output types for the method, not how it is implemented. Each concrete class extending DeptDirectory must provide a definition for findAddress that includes a code body describing how it is implemented.

The ordering of members within a class typically does not affect program behavior. Nevertheless, it is good programming practice to list class members in a consistent order. We recommend placing static final fields at the beginning of the class followed by dynamic members and finally static method (which have been discussed yet). Within each category, we recommend the following order: fields, constructors, methods. (Of course, there is no such thing as a static constructor.) According to this convention, the abstract method findAddress should be the last member in the DeptDirectory class.

Second, we must provide a concrete definition of the findAddress method in each concrete subclass of DeptDirectory, namely Empty and Cons. Note that the composite pattern guarantees that a program includes code specifically to process each data variant. Moreover, in any variant containing a field f of parent type, the method typically invokes itself recursively on f. This approach to defining methods is the direct analog of natural recursion template for defining functions in Scheme. It is so common and so important that OOP community has labeled it as a separate pattern, called the interpreter pattern, enriching the composite pattern.

Let us return to defining the findAddress method. By definition, there is no Entry in an Empty directory matching the name passed as an argument to findAddress. Hence, findAddress must return a value signaling failure. In Java, the most convenient choice for such a value is null, the reference to no object. All object values in Java are actually references, so the same object can simultaneously appear as the value of many different variables. Scheme follows exactly the same convention regarding structures. Java also provides the special value null, which is the reference to no object. There closest analog to null in Scheme is the special value (void). Java null should only be used to represent a special failure value. It should never be used to represent one of the alternatives in a data definition, e.g., the empty DeptDirectory. The reason for this prohibition is simple: null is not an object. Any attempt to invoke a method on null will generate a run-time error aborting program execution.

The following code is an appropriate definition of the findAddress method in the Empty class.

     String findAddress(String name) { return null; }

The definition of findAddress for Cons objects is the only interesting chunk of code in this entire example. Since a Cons object always contains an Entry

first and a DeptDirectory rest, the findAddress method must check to see

if the passed name matches the name field of first and, depending on the outcome, either return the value of the address or recur on rest.

The object-oriented method has exactly the same recursive structure as the corresponding function definition.

The method can simply be coded as follows:

     String findAddress(String name) {
       if (this.name.equals(this.first.getName()))
         return this.first.getAddress();
       else return this.rest.findAddress(name);
     }

Every class contains the instance method equals which takes a single argument of type Object. The Object class at the top of the class hierarchy defines this method. For a String object name, the equals method returns true if and only if the argument object contains exactly the same sequence of characters

as name.

In the code above, the expression

     this.name.equals(this.first.getName())

invokes the equals method of object in field name on the argument

     this.first.getName().

The expression

     this.first.getName()

invokes the getName method of the Entry object in the field first to get its name field. Similarly, the expression

     this.first.getAddress()

invokes thSe getAddress method of the Entry object in field first to get its address field; the expression

     this.rest.findAddress(name)

invokes the findAddress method of the DeptDirectory object in the rest field on the object name.

Notice that a Java instance method with n arguments corresponds to a Scheme function of n+1 arguments because the object containing the method is the implicit argument this. Since the members of this are so frequently accessed in methods attached to the object this, Java allows the field prefix

     this.

to be omitted! Hence, the definition of findAddress could have been written

     String findAddress(String name) {
       if (name.equals(first.getName()))
         return first.getAddress();
       else
         return rest.findAddress(name);
     }

Since explicit references to this clutter code making it more difficult to read, we will generally omit them. (In some situations, a method must refer to the entire receiver object this rather than one of its fields. In such a situation, the use of the keyword this is essential.)

Finger Exercise 1.6.3.1 Load the file DeptDirectory.java into the Dr-Java Definitions pane. Write a simple test class for this program. Compile the program and try some test lookups in the Interactions pane. Write a method findPhone analogous to findOffice and test it. Remember to use the same design steps that you learned for Scheme. Save your program in a file called

DeptDirectory.java.

1.6.4 Conditional Statements

In the definition of the findAddress method, we used an conditional statement of the form:

     if (test) consequent else alternative

where test is an expression of boolean type and consequent and alternative are statements. Conditional statements are used to classify program values into disjoint sets or regions using logical tests that we call claims. In simple example given above, we distinguished between two claims:

     name.equals(first.name)

and

     !(name.equals(first.name))

1.6.5 Blocks

In Java, braces are to aggregate sequences of statements into individual statements. A sequence of statements

     {
       $s_1$;
       $s_2$;
       ...
       $s_n$;
     }

enclosed in braces is called a block. A block is a form of Java statement. The other forms of statements that we have seen so far are variable definitions, assignment statements, conditional statements, and method calls.

Suppose that we wanted to print a message every time the findAddress method failed to match a name in a DeptDirectory. We need to add a statement to the else clause of our conditional statement in the body of findAddress in class Cons. We can accomplish this task by surrounding the return statement in the else clause with braces and inserting our print statement before the return statement as shown below:

     String findAddress(String name) {
       if (name.equals(first.getName()))
         return first.getAddress();
       else {
         System.out.println(first.getName() + " does not match");
         return rest.findAddress(name);
       }
     }

Why not insert the print statement after the return statement instead of before?

1.6.6 Singleton Pattern

One of the most important uses of static final fields is storing the canonical instance of a class that only needs a single instance. For example, the Empty subclass of DeptDirectory only needs one instance because the class has no (dynamic) fields.

     class Empty extends DeptDirectory{
       ...
       static final Empty ONLY = new Empty();
     }

Instead of allocating new instances of Empty, code can simply refer to the canonical instance Empty.ONLY. The final attribute stipulates that the variable ONLY cannot be modified. This code pattern is called the singleton pattern because it constructs a single instance of the class.

The implementation of the singleton pattern shown above suffers from an annoying defect: the class definition for Empty does not prevent code in another class from creating additional instances of the Empty class. We can solve this problem by making the constructor for Empty private.

     class Empty extends DeptDirectory{
       ...
       private Empty() {}
       static final Empty ONLY = new Empty();
     }

Then code outside of class Empty cannot perform the operation

     new Empty();

A private member of a class C is only visible inside class C. We will discuss visibility modifiers in more detail in Section 1.7.7.

Finger exercise 1.6.6.1: Modify your DeptDirectory program to use the singleton pattern to represent the Empty DeptDirectory.

1.7 Basic Program Design

In the preceding sections of this monograph, we studied a Java subset suitable for explicating the basic principles of Java program design. As you recall from your Scheme background, the process of program design can be broken down into six steps:

• Data Analysis and Design

• Contract and Header

• Examples

• Template

• Body

• Test

which we collectively call the design recipe. Let us examine each of these six steps in the context of writing object-oriented programs in Java.

1.7.1 The Design Recipe

Data Analysis and Design

Our first task in writing a program is to understand and define the data that the program will process. We must compile an inventory of the various forms of data that can appear as program inputs and outputs and determine how to represent their “relevant” properties as Java data objects. In scientific problem solving, this process is often called “data modeling” or simply “modeling”. For each distinct form of data, we must define a Java class with fields that represent the relevant properties of each object of that form. We use the composite pattern to specify which different forms of data belong to the same more general category of data. In the preceding section, for example, we grouped the Empty department directory and non-empty Cons department directories together using the composite pattern to form the more general

category DeptDirectory.

Java class definitions are more general and flexible than Scheme struct definitions because they enable the programmer to determine (i) exactly which primitive operations, including constructors, the new form of data will support, (ii) how objects will be printed as strings, (iii) and the types of object fields and methods. In Scheme, the set of operations generated by a struct definition is rigidly determined and Scheme does not include any provision for specifying the types of struct fields and operations.

The extra generality provided by Java comes at a price. A Java programmer must write far more text to define a class than a Scheme programmer does to define a comparable struct. Of course, some of the classes in a Java data definition, e.g., the abstract class at the top of a composite class hierarchy, have no analog in Scheme, forcing the Scheme programmer to write explanatory comments instead.

It is a good idea to define a collection of examples for each concrete class in a data definition. These examples can be defined as static fields of the class.

Contract and Header

Since Scheme does not accept type declarations for function parameters, a wellwritten Scheme program must include a header for each function containing this information. In contrast, Java requires type declarations for all method parameters. Hence, every syntactically correct Java program includes headers for all methods. Half of this programming step is mandated by the Java language!

On the other hand, Java does not mandate the inclusion of a contract stating (i) what if any additional preconditions (beyond the the types of the arguments) must hold for the method to be called, and (ii) what relationship exists between the inputs and output of a method. The latter is often called a postcondition or output condition for the method. Well-written Java programs include preconditions and postconditions for all methods other than the trivial methods forming the data definition, e.g. stock constructors and accessors.

For methods defined over composite types (abstract classes at the top of composite class hierarchies), the contract information should be attached to the abstract method definition in the top class.

Examples

Each class should include a collection of sample inputs and corresponding outputs for each method. If the body of a method contains more than one control path (e.g. a conditional statement), the collection of examples should include at least one example for each control path. The examples for each method should be included in either a test method for the class or a separate test class. We will explain how to set up tests in separate test classes later in this monograph.

Template

In Java, much of the template selection process is mandated by the objectoriented programming model. When a data type T consisting of several different forms of data is represented by a composite class hierarchy, each method m defined on T must be defined in the abstract class T corresponding to T. With a few rare exceptions, the method definition for m in T must be abstract, forcing each concrete subclass of T to provide a definition for m. This decomposition corresponds to the cond template used to process (non-recursive) mixed data in Scheme. Moreover, the relevant data for each definition of m is simply the set of object fields in the class containing m!

The only features in the template for a Java method definition that are not dictated by the object-oriented programming model are the recursive method calls corresponding to circular references in the data definition. For any method m in a class C containing a object field f of type T where T is a supertype of C (e.g. C is a concrete class in a composite class hierarchy with T at the top), the body of m will usually invoke m recursively on each such field f. These recursive calls appear in the template. For a concrete example, see Section 1.7.2.

Body

The coding part of the design process consists of filling in the body of each method m in each concrete class C using the available object fields and the results of recursive calls from the template. In some cases, writing this code requires ingenuity. But in the vast majority of cases, the coding process is very easy given decomposition provided by the template.

Test

For each major class C or composite hierarchy of classes in a Java program the programmer should write a test class TestC including a test method that evaluates each publicly visible operation of C on appropriate sample data values (both instances of C and the method arguments). This process is described in much more detail in Section 1.8.

1.7.2 An Extended Example: Lists

Let us study how the design recipe described above applies to the process of writing some simple programs involving lists of integers. Lists are ubiquitous in programming. The DeptDirectory type introduced in Section 1.6 is simply a specialized form of list. For reasons that will become clear later in the monograph, we will use the name IntList rather than List for our list type. As you should recall from Scheme background, a IntList is either:

• an Empty list, or

• a composite list Cons(e,l) where e is the first Object of the IntList, and l is the rest of the IntList.

Compare this definition with the definition of DeptDirectory given earlier.

We can abbreviate the preceding definition using the following mathematical notation:

     IntList := Empty() + Cons(int, IntList)

which states that the set IntList is the union of the sets Empty and Cons. The set Empty contains a single object Empty() while the set Cons contains all objects Cons(o,l) where o is any int and l is any element of type IntList.

Assume that we are given the task of writing a program to perform some standard operations on lists of integers such as computing the sum of the numbers in a list, computing the product of the numbers in a list, and sorting the members of a list into ascending order.

Since we need to perform a variety of different operations on IntList, we will include a full set of getter operations (also called selectors or accessors), namely getFirst and getRest methods corresponding to first and rest in Scheme. The following collection of Java classes provides a minimalist definition for the IntList type:

     /** Composite Data Definition:
      *  IntList := new Empty() + new Cons(int, IntList)
      */
     abstract class IntList {
       /** Returns first element of non-empty list.
        *  Throws an IllegalArgumentException on Empty.
        */
       abstract int getFirst();
       /** Returns the ``rest'' a non-empty list (all elements but first).

 *  Throws an IllegalArgumentException on Empty.
 */
abstract IntList getRest();
/** Returns "" for Empty
 *          " e1 e2 ... en" for IntList with elts e1,...,en
 */
abstract String toStringHelp();
}
class Empty extends IntList {
/** The only instance of class Empty */
static final Empty ONLY = new Empty(); // singleton pattern
/* constructor */
private Empty() {}
/* Cons accessors */
int getFirst() { throw new IllegalArgumentException(
 "first requires a non Empty IntList");
}
IntList getRest() { throw new IllegalArgumentException(
 "rest requires a non Empty IntList");
}
/* other methods */
public String toString() { return "()"; }
String toStringHelp() { return ""; }
}
class Cons extends IntList {
/* private fields */
private int first;
private IntList rest;
/* constructor */
Cons(int f, IntList r) {
  first = f;
  rest = r;
}

       /* accessors */
       int getFirst() { return first; }
       IntList getRest() { return rest;  }
       /* other methods
       public String toString() {  // no leading space before first
         return "(" + first + rest.toStringHelp() + ")";
       }
       String toStringHelp() {     // leading space before each elt
         return " " + first + rest.toStringHelp();
       }
     }

These three classes form a conventional composite class hierarchy with IntList at the top. These class definitions rely on Java exceptions to handle erroneous uses of the getFirst and getRest methods. We will discuss exceptions in detail in Section 1.12.3. For now, all you need to know is that (i) all exceptions are elements of the the built-in type Exception and (ii) the throw statements in the bodies of getFirst and getRest in class Empty abort the computation and print the specified strings as part of the error message. Their meaning in this context produces exactly the same results as invoking the Scheme construct

error.

The IntList class includes the method toStringHelp() as a “help” method for printing lists with exactly the same spacing as in Scheme. The String representation for every list is enclosing in parentheses, but leading blanks appear every element after the first. The toStringHelp() method is responsible for printing the elements of a list tail that excludes the first element. Hence, toStringHelp() generates a space before each element that it processes. In contrast, toString() generates the enclosing parentheses and the first element if present; it delegates formatting the remaining elements of the list to

toStringHelp().

Let us compare the IntList defined above with the built-in lists of Scheme. Scheme provides the primitive operations empty, cons, getFirst, and getRest

akin to the Java operations new Empty(), new Cons(...), first, and rest.

Scheme also provides recognizers for empty and compound lists called empty? and cons? which have no visible analog in the preceding Java program. We could try to define analogous operations in our Java data definition by including abstract boolean operations isEmpty and isCons in the abstract class IntList and define them appropriately in the concrete subclasses Empty and Cons. But this approach does not achieve our objective or defining operations equivalent to the Scheme recognizers, because the isEmpty and isCons methods are applicable only to objects of type IntList.

1.7.3 Type Predicates and Type Casts in Java

Most programming languages do not have an analog of Scheme predicates like empty? because they do not have a universal type that contains all other types. Java is almost identical to Scheme in this regard. All object types are subtypes of the universal type Object. If we ignore the eight primitive types (which all have corresponding wrapper types in the Object type hierarchy), then the data models of Scheme and Java are essentially identical.

To test membership in any object type, Java provides a collection of postfix operators of the form

     instanceof T

where T is any defined object type. Hence, given the preceding program defining type IntList, Java interprets the program expressions below as follows:

     new Empty() instanceof Empty ⇒ true
     new Cons(0, new Empty()) instanceof Empty ⇒ false
     "A" instanceof Empty ⇒ false

The instanceof operator has the same precedence as the relational operators. (Although the second “argument” to instanceof must be a type name, the Java parser initially recognizes this argument as an expression.)

Finger exercise 1.7.3.1: Load the sample program IntList.java into the DrJava Definitions pane. Add definitions for isEmpty and isCons. In the Interactions pane try evaluating the following sequence of interactive computations:

     IntList empty = Empty.ONLY;
     empty
     IntList oneElt = new Cons(1, empty);
     oneElt
     empty.isEmpty()
     empty.isCons()
     oneElt.getFirst()
     oneElt.isEmpty()
     oneElt.isCons()
     IntList twoElts = new Cons(0, oneElt);
     twoElts.getFirst()
     twoElts.getRest()
     twoElt.getRest().isCons()
     empty.getFirst()

     empty.getRest()
     "A".isEmpty()
     "A".isCons()

Perform the equivalent sequence of membership tests as in the previous exercise using instanceof operators instead of the operations isEmpty and isCons.

To accomodate static type checking, Java includes a second form of type predicate not present in Scheme called a cast. You may recall that Java includes operations for casting one primitive type to another. These primitive type casts convert values of one type to “corresponding” values of another type. The casting operations for object types have a completely different meaning; casting a value v to an object type T peforms an instanceof check on v! If the check returns false, then Java throws a ClassCastException indicating that the cast failed. If this exception is not caught (see Section 1.14.3), Java aborts execution and prints an error message indicating which cast failed. In contrast, primitive type casts never fail!

If object type casts can only cause a program to abort execution, what good are they? Since the cast prevents execution from continuing if the instanceof test fails, the compiler knows that the result of object casting expression

     (T) e

has type T. Consequently, the static type checker in the compiler assigns the static type T to this casting expression. By inserting object casting operations in a program, you can tell the static type checker that a particular expression has a narrower (more precise) type that the type that would otherwise be assigned by the static type checking rules.

Finger exercise 1.7.3.2: Load the program IntList.java into the Definitions pane of DrJava. In the Interactions pane try evaluating the following sequence of interactive computations:

     IntList empty = new Empty();
     IntList oneElt = new Cons("B", empty);
     oneElt
     oneElt.first
     ((Cons) oneElt).first
     oneElt.rest
     ((Cons) oneElt).rest

Perform the equivalent sequence of membership tests as in the previous exercise using instanceof operators instead of the operations isEmpty and isCons.

1.7.4 A Sample IntList Program

We are now ready to define a simple program to sum the integers in a lists. We will add a definition of the method

int sum();

to each class in our composite class hierarchy for IntList. Let’s begin by writing the contract and header for sum in the abstract class IntList:

/** Composite Data Definition:
*  IntList := new Empty() + new Cons(int, IntList)
*/
abstract class IntList {
...
/** Compute the sum of the elements in this. */
abstract int sum();
}

Next we need to generate examples showing the expected behavior of the method:

class TestIntList {
/** Sample IntList values */
static final oneElt  = new Cons(1, Empty.ONLY);
static final twoElts = new Cons(5, oneElt);
static final threeElts = new Cons(-10, twoElts);
/** Test the sum() method, assuming that ans is
 *  the correct answer for this.sum() */
void testSum(int ans) {
  int result = sum();
  System.out.println("computed sum of " + this + " = " + sum());
  if (result != ans)
    System.out.println("FAILURE: correct answer is " + ans);
}
/** Test driver method */
public void test() {

         Empty.ONLY.testSum(0);
         Cons.oneElt.testSum(1);
         Cons.twoElts.testSum(6);
         Cons.threeElts.testSum(-4);
       }
     }

As the fourth step, we select a template for writing the sum method:

     class Empty {
       ...
       int sum() { ... }
     }
     class Cons extends IntList {
       int first;
       IntList rest;
       ...
       int sum() { ... first  ... rest ... rest.sum() ... ; }
     }

Finally, we complete the coding process by filling in the bodies of the methods in the template:

     class Empty {
       ...
       int sum() { return 0; }
     }
     class Cons extends IntList {
       int first;
       IntList rest;
       ...
       int sum() { return first + rest.sum(); }
     }

To finish the design recipe, we test our code using the examples in the main

method of TestIntList.

Finger Exercise 1.7.4.1: Load your saved file IntList.java into Dr-Java Definitions pane and add the definition of the sum method as described above. In the Interactions pane, run the tests specified in the test() method of IntList. Using the same design recipe, add a definition of a method prod to compute the product of a list of numbers and test it. Note that the Data Analysis and Design step has already been done in IntList code. Save your revised program in the file IntList.java.

1.7.5 Inheritance and the Composite Pattern

Up to this point, we have used methods in Java essentially like functions in Scheme with the exception of overriding the definition of toString() in several classes. What makes object-oriented programming truly powerful is the ability to add new forms of data to our program without modifying any old code. For example, if we later decide to insert links to other directories as additional form of DeptDirectory data, we can simply define a new subclass Link of DeptDirectory with a subDir field referring to the embedded directory (which can be searched using the findAddress method. The new class must define findAddress for the new form of data that it introduces, but none of the existing classes requires any change whatsoever.

In defining a program extension, the added data does not have to be a new subclass of an abstract class like DeptDirectory. The new data can be a subclass of an existing concrete class. For example, we could extend our directory program by defining a class EntryWithTitle extending Entry with a String field title specifying the person’s title (graduate student, instructor, professor, chair, etc.) and a getter method getTitle() to retrieve it. No revision of the Entry class would be required. Unfortunately, to extract this information using the programming techniques discussed so far, we would have to add a new method findTitle to the composite class hierarchy DeptDirectory— modifying existing code. We will introduce a design pattern, called the visitor pattern near the end of this chapter that addresses this problem.

When a class C extends another class D, every instance (non-static) member m of D is automatically an instance member of C with one exception: if C redefines the method m^{5} We say that the instance members of class D are inherited by the subclass C. The linguistic convention of automatically incorporating the instance member definitions of a class in each of its subclasses is called inheritance. If an inherited method m is redefined in class C, we say that the new definition overrides the inherited definition. We have already seen some simple examples of inheritance and overriding involving the toString() method.

We have already made extensive use of a limited form of inheritance in the composite pattern. All of the variant classes of in a composite hierarchy provide definitions for the abstract methods inherited from the abstract superclass. Recall the DeptDirectory and IntList programs. When an abstract method from an abstract class is overridden in a concrete subclass (e.g. findAddress from DeptDirectory in the classes Empty and Cons), the “missing” definition inherited from the abstract class is overridden. This special form of overriding is sometimes called method extension because the inherited meaning of the method (nothing) is extended rather than modified.

5The redefined method must have exactly the same name and input and output types as the method m that would have been inherited.

When a class C overrides a method m in its superclass D, code in the body of C can still invoke the overridden method m from D using the special notation

     super.m( ...  )

The feature can be used to add pre-processing and post-processing code to an inherited method m. The overriding definition of m can check or transform its inputs, invoke super.m on the transformed inputs, and subsequently check or transform the result produced by the super.m call.

It is important to remember that all unqualified member references in inherited code implicitly refer to the implicit method argument this which is bound to an instance of the inheriting class! Hence, the meaning of a method call appearing in inherited code changes if the specified method has been overridden! In practice, this semantic convention gives the behavior that programmers expect for instances of subclasses, but it can occasionally produce surprises—particulary if the method overriding is the result of an accidental rather than an intentional name match.

Finger exercise 1.7.5.1: Load the superCall sample program into the DrJava Definitions pane. The body of method length in class Vector includes calls on the instance method getLeft and getRight. The class TranlatedVector which extends Vector contains a super call on length in its overriding definition of length. To what object is this bound in the inherited length method when it is invoked by the super call?

The meaning of field names in inherited code is even more subtle than the meaning of method names, because fields are never overridden. If the extending class C defines a field with the same name as an inherited field, the inherited field is merely “shadowed” exactly as it would be by a local variable with the same name. The inherited field still exists in each instance of C and can be accessed by code in the body of C (assuming that it not private, see Section 1.7.7) by casting the C object to its superclass type D before extracting the field.

If you follow the programming guidelines recommended in this monograph you can generally avoid the pathologies discussed in the preceding paragraph. According to our stylistic rules, field references should never span class boundaries. Hence, shadowed members are only accessed through getter methods inherited from the superclass. The only field that can be accessed directly is the one defined in the current class.

1.7.6 Overriding equals

A particularly important and subtle example of inheritance in Java involves the method

     public boolean equals(Object o);

which is defined in the class Object, the superclass of all Java classes. Any Java class that is defined without a designated superclass is an immediate subclass of the Object class. In the class Object, equals is defined to mean Object identity. Two object references are identical if and only if they refer to exactly the same object (produced by a particular new operation).

For some classes, identity is the appropriate definition for equality, but for many others it is not. In the built-in class String, the equals method is overridden to compare the sequences of characters in strings, so that copies of the same string are considered equal. The redefinition of equals only affects the class String and any subclasses that it might have.

The overriding of the equals method is particularly delicate because the Java libraries all assume that equals defines an equivalence relation on the universal type Object, excluding the special value null, which is treated as a special case. In particular, for all non-null x and y, x.equals(y) iff y.equals(x). If the argument to equals is null, then the Java API specification stipulates that equals must return false.^{6} If the class containing the overriding definition of equals can be extended (subclassed) then the coding of equals is quite subtle.

In particular, the overriding definition must confirm that the argument o belongs to exactly the same class as this. Assume that we are overriding the definition of equals in the composite class hierarchy IntList given in Section 1.7.2 above. The following code for the definition of equals in the Cons does not work in general!

       public boolean equals(Object o) {
         return (o != null) && (o instanceof Cons) &&
           (first == ((Cons)o).first) && rest.equals(((Cons)o).rest);
       }

This code can fail if Cons has a subclass ExtCons because equals can report that an instance of ExtCons is equal to an instance of Cons. Even worse, if equals is overridden in ExtCons using the same instanceof pattern,

6This part of the equals specification is poorly designed because it unnecessarily complicates the behavior of equals without providing any useful benefit. A better specification would have stated that the behavior of equals on null was unspecified because null is outside the intended domain of the method.

       public boolean equals(Object o) {
         return (o != null) && (o instanceof ExtCons) &&
           (first == ((ExtCons)o).first()) && rest.equals(((ExtCons)o).rest());
       }

a.equals(b) does not imply b.equals(a) For example, if a is an instance of Cons and b is an an instance of ExtCons with exactly the same first and rest fields as a, a.equals(b) will be true while b.equals(a) will be false (because

a instanceof ExtCons is false.

The problem with the instanceof pattern for writing equals is that instanceof does not test for an exact class match. We can compare the classes of objects by using the method getClass() which is inherited by all classes from Object. This method returns an instance of the built-in class Class representing the class of the receiver object. In addition, we can get the Class object for any specific class C, simply by typing C.class. Every class has exactly one Class object representing it. Hence, the equals method for Cons above can be rewritten:

       public boolean equals(Object o) {
         return (o != null) && (this.getClass() == o.getClass()) &&
           (first == ((Cons)o).first) && rest.equals(((Cons)o).rest);
       }

Finger Exercise 1.7.6.1. Load the sample program intList into the DrJava Definitions pane. Override the definition of equals for both Empty and Cons to match the definition of the equal? function in Scheme on lists of integers. The Scheme equal? function compares two lists to determine if they contain the same sequence of elements. Try evaluating a substantial set of test cases in the Interaction pane of DrJava.

1.7.7 Helper Methods, Packages, and Visibility

The coding of non-trivial methods often involves the use of auxiliary methods called “help” methods. The specified operation may be easily derivable from another operation that has a simpler definition. For example, in the preceding definition of the composite class hierarchy IntList, we introduced a helper method toStringHelp to help support the printing of lists in the same format that Scheme uses. The toStringHelp prints the rest of a non-empty list with a leading blank before each element but no trailing blanks or closing parenthesis.

Since helper methods are defined strictly for the use of code within the class, we would like to prevent the definition from “leaking” outside the class. Java provides a mechanism for preventing such leaks. Class members can be assigned one of four visibility levels private, default, protected or public. private members are visible only within the class in which they are defined. Default members are visible only within the package in which they are defined. protected members are visible in the package in which they are defined and in subclasses of the defining class. public members are visible everywhere.

In section 1.3.5, we stated that the only way to access the fields of an object is through “getter” methods provided by the class definition. If we always declare the instance (non-static) fields of a class as private, then this statement is completely accurate. We strongly recommend following this convention; it supports to the object-oriented principle of separating the implementation of a class from its interface.

We have avoided mentioning the Java package system until now because it is not helpful in writing programs of modest size. Large Java programs typically are partitioned into packages analogous to the file directories in a tree-structure file system. Each package, except for the “default” package discussed below, has a unique name consisting of one or more Java identifiers separated by periods. Hence java, java.lang, and java.awt.event are are all valid package names.

Every Java class belongs to some package. If a source file does not mention a package name, then it is considered part of a special “default” package with no name. In this monograph, we will use the default package for all of the code that we write. On the other hand, all of the Java core library code that we will use resides in named packages. The Java libraries are partitioned into packages

like java.util, java.awt, java.awt.event and javax.swing. Packages are not

nestable. There is no connection between java.awt and java.awt.event other than a common name prefix.

The private attribute is well-suited to hiding helper methods that aren’t required in subclasses. The protected attribute is useful when helper methods are referenced in subclasses residing in other packages. In our example above, toStringHelp is accessed by all of the subclasses of IntList. Hence, the appropriate protection mechanism for our toStringHelp is either default or protected. Since all our program classes reside in the same package, it doesn’t matter. However, if we wanted to define subclasses of IntList in another package, we would need to declare the toStringHelp method as protected to make it visible within these subclasses.

When an inherited method is overridden, it cannot be made less visible. Hence, an overridden public method must be declared as public. On the other hand, an overridden protected method may be declared as public.

Finger Exercise 1.7.7.1 Load the sample IntList program into the Dr-Java Definitions pane. Convert the testSum() method to a private method. Confirm that the test() method for the IntList class still executes the test suite for the sum method. In the Interactions pane, try evaluating the following sequence of statements:

     IntList l = new Cons(17, new Cons(13, Empty.ONLY));
     l.testSum(30);

Finger Exercise 1.7.7.2 Load the sample IntList program into the Dr-Java Definitions pane. Add a method

     IntList insert(int i);

to the IntList class that inserts int i in this assuming that this is already sorted into ascending order. Test your code.

1.8 Unit Testing in Java

Program testing can be conducted at different levels of program granularity: expressions, statements, individual methods, public methods, and the user interface. Testing at both fine and coarse levels of granularity is essential for achieving program reliability. Testing at a fine-grain of granularity is called unit testing. The goal of unit testing is to check the behavior of each identifiable unit of program behavior in isolation from all other units. The smallest unit of program behavior that is accessible in compiled Java code is the individual method (excluding private methods which cannot be accessed directly).

The practice of unit testing can be a very effective tool for building robust programs of any size. A strong suite of tests, developed in tandem with the code they test, can aid our efforts to develop robust code in many respects. Errors can be caught more quickly. Code can be refactored more easily. Other programmers can understand the functionality of other programmer’s code more quickly. But writing unit tests to maximum advantage requires that we employ several important principles.

1.8.1 Tests Should Check Themselves

First and foremost, the tests we write should be capable of checking themselves. Consider the following example of a testing method:

     public void testCityEntryToString() {
       System.out.println(new CityEntry(...).toString());
     }

This test checks the result of calling toString() on a CityEntry by printing the result to the console. The programmer must check this output and verify that it matches the intended result. This is a waste of programmer time and effort; it would be much more efficient to write test methods that checked themselves. For example, we could rewrite the above test as follows:

     public void testCityEntryToString() {
       assertEquals("CityEntry[a,b,c]",
                    new CityEntry("a","b","c").toString());
     }

where we call an assertEquals method that compares the value of its arguments to ensure that they are equal. If they aren’t equal, assertEquals throws an exception. This way, we know whenever we run this test that it passed unless we were notified with an error message.

1.8.2 Tests Should be an Integral Part of the Coding Process

When we write tests, we should keep in mind that the tests we write are a very powerful form of documentation. When another programmer reads our tests, he can quickly learn how to invoke our code, and how it will behave What’s more, he can verify that our code meets this documentation by running the tests (compare this with English comments). For this reason, we should be careful to develop small and relatively simple tests that are easily readable by others.

Additionally, these tests should capture the most interesting places where the code may break. Determining where these places are is best done when the code is actually written. For this reason, it’s a good practice to alternate writing tests and writing code in small steps, checking at each step that all tests still pass. Not only will this help you to write penetrating tests while the code is fresh in your mind, but it will also help you to discover bugs much more quickly and easily, as you will know when a test fails that only a very small block of code could be the culprit.

1.8.3 Write The Tests First

When interleaving the writing of tests with code, it’s best to write the tests for each section of code before the corresponding code is written. By writing the tests before we write the code, you can help to solidify your understanding of what you want that code to do. You also know when we are done writing the code: when it passes all of the tests you’ve written. And if the tests don’t capture all of our intended functionality, it’s time to write more tests and repeat!

Another advantage of ”test-first” programming is that the code you write will be designed from the very start to be easily testable, as you will never even write code until you’ve determined how it will be tested. Avoid the temptation to declare, for a particularly hard problem, ”this code can’t possibly be tested”. Instead, ask yourself, ”how could I design this code in such a way that I can test it?” In virtually all cases, you will find that if you approach the problem from this perspective, you will be able to devise a design that you can test quite naturally.

1.8.4 Unit Testing and JUnit

JUnit has provided a powerful facility for running unit tests in Java. In fact, the code for DrJava itself is tested using an extensive suite of JUnit tests. Of the 53,000 lines of code that comprise the latest build of DrJava, 16,000 lines consist of unit tests. As part of our build process, we ensure that these tests must be run before any change to the code is committed.

To help users test their own programs, support for running JUnit tests is included in DrJava. To use JUnit on a test file, the following programming conventions must be used:

At the top of the file, include:

     import.junit.framework.TestCase;

This line is an import statement that tells Java to include the JUnit class TestCase in the namespace of the file. The main class of the file must:

• be public,

• extend TestCase, and

• have a constructor of the form:

          public classname(String name) { super(name); }

Note: If you forget to do this, you will get an error message that begins:

”TestCase() is not public...”

These conventions are necessary is because class TestCase has a single public constructor, and this constructor takes a String as its parameter. As with all Java classes, we have to call a constructor of the superclass within the constructor of the extending class. By default, the zero-argument will be called implicitly, but there is no zero-argument constructor to call on class TestCase.

The test methods of this class (methods to be run automatically when the Test command is invoked) must:

• be public and not static,

• return void,

• take no arguments, and

• have a name beginning with test

Test methods in this class can call any of the following methods:

• void assertTrue(String, boolean) which issues an error report with

the given String if the boolean argument is false.

• void assertEquals(String, int, int) which issues an error report with

the given String if the two integers are not equal. The first int is the expected value, and the second int is the actual (tested) value.

• void assertEquals(String, Object, Object) which is just like the assertEquals

method on ints, except that it takes Objects as its last two parameters.

• void fail(String) which immediately causes the test to fail, issuing an error report with the given String. Note that test methods are permitted to throw any type of exception, as long as it is declared in the throws clause of the method contract. If such an exception is thrown, the test fails immediately.

If there is any setup work to be done before running the tests (such as initializing instance variables), do it in the body of a method with the following contract: protected void setUp() This method is automatically run before each test in the class. There is a corresponding method, tearDown() that is called after each method. tearDown() is useful for releasing shared resources, such as open files or database locks.

Alternatively, a test suite can be constructed to invoke a series of test methods under programmer control with a method of the form:

 public static Test suite() {
   TestSuite suite = new TestSuite();
   suite.addTest(new classname("methodname");
   ...
   return suite;

}

It is then also necessary to import TestSuite and Test from junit.framework. There is also a version of the addTest() method that takes a Test, so test suites can be composed.

To start JUnit on a TestCase in DrJava, simply click the Test button (or choose ”Test Using JUnit” from the Tools menu) when a valid test file is open, and DrJava will run the tests and display any test failures in the ”Test Output” tab at the bottom of the window, much the same as compiler errors are displayed.

1.8.5 Conclusion

As we’ve seen, unit tests can be a great tool for writing robust code. But as programs become increasingly complex, the tests over that code become more complex as well. JUnit provides many sophisticated ways to test code, beyond just what we’ve presented in this short introduction. For more examples and information, see the JUnit Cookbook, available at http://www.junit.org.

1.9 Interfaces

We have not yet discussed one of most important object-oriented features of Java, namely the notion of an interface. In essence, an interface is a special lightweight form of abstract class. We use the term lightweight to describe an abstract class with no fields and no concrete methods; the only members are abstract methods. The key difference between an interface and the corresponding abstract class is that a class or interface can have unlimited number of immediate superinterfaces but it can only have one superclass.

An interface definition has almost the same syntax as a class definition:

     interface name {

... member declarations ...

}

The only significant difference is the keyword interface instead of class.

In this monograph, we will follow the convention that interface names begin with “I” followed by a capitalized name. Hence, “IList” is the name of an interface, while “List” is the name of a class. There is no generally accepted convention for distinguishing interface names from class names in the Java programming culture. In the Java libraries, for example, there is no way to tell a class from an interface based on its name.

A class can implement an arbitrary number of interfaces. A class definition can optionally include an implements clause immediately following the extends clause (assuming one is present) in the class header. Sample programs using interfaces appear in 1.11.

1.9.1 Multiple Inheritance

Some object-oriented programming languages like C++ permit a class to have multiple superclasses. This form of inheritance, which is called multiple inheritance, is very controversial. While clever uses of multiple inheritance abound, the semantics of code and field inheritance with multiple superclasses is problematic because of name clashes between members and, worse, the possibility of inheriting the same class in more than one way. Recent programming language research suggests that there are better approaches to inheritance that combine the flexibility of multiple inheritance with the simplicity of single inheritance, but they are currently the subject ongoing research and experimentation.

Java supports multiple interface inheritance, which the most attractive proven alternative to multiple inheritance. In multiple interface inheritance, a class can extend multiple lightweight abstract classes, but only one class that includes method code and fields. Since no method code or fields appears in the lightweight abstract classes, there is no problem with name clashes (since all interface methods are publicly visible) or inheriting the same interface in multiple ways. We will illustrate the utility of Java interfaces in the next section. At this point, all we can say is that the abstract classes DeptDirectory and IntList could be declared as interfaces instead if we were willing to make all of their methods public and delete the test method from IntList.

Since the abstract class DeptDirectory above does not contain any members, it could be trivially rewritten as follows:

     interface IDeptDirectory {}

We will follow the convention that interface names begin with the capital letter I followed by a name capitalized like a conventional Java class. This convention is not a commonly accepted standard, but it clearly distinguishes interfaces from classes.

This change forces a small change in the definition of any immediate subclass of IDeptDirectory: the word implements must be used in the header instead of extends:

     class Empty implements IDeptDirectory  ...
     class Cons implements IDeptDirectory  ...

In short, a class implements an interface but extends a class.

We generally recommend using abstract classes as the root classes in composite hierarchies because they frequently need to introduce methods that should be hidden (not public) and methods that are not abstract. On the other hand, some programs involve multiple composite class hierarchies that share concrete subclasses. In this case, you must define the root classes as interfaces or define two copies of the “shared” classes.

1.9.2 Implicit Polymorphism

A cardinal rule in Scheme programming is “Never repeat code”. If a program repeats essentially the same code in two or more places, then the programmer failed to identify a common abstraction that should only be written once. The repeated code sections can be replaced by calls on a procedure that defines the repeated operation. Any minor differences between the repeating code sections can be accommodated by passing appropriate arguments that “fill in” the differing expressions.

The same rule applies to Java, but the notational details are more cumbersome because (i) methods cannot be directly passed as arguments in Java and and (ii) Java is statically typed. Passing methods as arguments is such an important programming technique that object-oriented programmers have developed a design pattern, called the command pattern, that enables Java programs to indirectly pass methods as parameters by embedding them in “dummy” objects called commands. This pattern is discussed in the next subsection. The complicating effects of Java’s static type discipline are illustrated by the following example.

Consider a program that manipulates lists of several different types of elements. One approach is to define a separate composite class hierarchy for each kind of list. But this approach requires replicating essentially the same code in the definition of each class. To avoid this code replication, we can define a single composite class hierarchy for lists of type Object. Since all Java object types are subtypes of Object, such a list type can be used in place of any specific list type. However, when we extract an element from such a list, we will generally have to cast it to the specific type required by the context in which it is used. These casting operations clutter the code and reduce the precision of static type checking. Nevertheless, the advantages conferred by avoiding code replication usually outweigh these disadvantages.

Finger Exercise 1.9.2.1 Load the IntList.java program into the DrJava Definitions pane. Convert it to a definition of a class ObjectList where the list elements have type Object instead of type int. Test this program and save it in a file ObjectList.java for future use. Finger Exercise 1.9.2.2 Load the program in your saved file ObjectList.java into the Definitions pane. Define a method

     ObjectList sort()

that sorts a list of Integer into non-descending order. Your code will need to use casting operations confirming that the elements in the receiver of a sort invocation have type Integer. Test your code. What happens if you try to sort a list containing elements that are not of type Integer? Finger Exercise 1.9.2.3 Load your program ObjectList.java program into the DrJava Definitions pane. Modify the sort method to work for any element type that supports the Comparable interface from the Java API.

1.9.3 Interface Types

An interface identifies a program type independent of any specific class. This mechanism enables Java to express computations in more abstract and flexible terms. Consider the following example. Java includes a built-in interface named Comparable with the following definition:

     interface Comparable {
       int compareTo(Object o);
     }

All of the methods in an interface are automatically abstract and public. Let us define a class CompList similar to IntList where list elements have the type Comparable. An object has type Comparable iff it is an instance of a class that implements the Comparable interface. Interface implementation is inherited: if a class C implements an interface I then all subclasses of C implement I also.

In this context, we can define the class CompList as follows:

     abstract class CompList {
       abstract Comparable getFirst();
       abstract CompList getRest();
       abstract String toStringHelp();
       public static void main(String[] args) {
         Empty empty = new Empty();
         System.out.println("twoElts is: " + Cons.twoElts);
         System.out.println("fourElts is: " + Cons.fourElts);
       }
     }

class Empty extends CompList {
Comparable getFirst() { throw
   new IllegalArgumentException("getFirst() requires a non-Empty CompList");
}
CompList getRest() { throw
  new IllegalArgumentException("getRest() requires a non-Empty CompList");
}
public String toString() { return "()"; }
String toStringHelp() { return ""; }
public static void main(String[] args) {
  CompList e = new Empty();
  System.out.println("() = " + e);
}
}
class Cons extends CompList {
Comparable first;
CompList rest;
Cons(Comparable f, CompList r) { first = f; rest = r; }
Comparable getFirst() { return first; }
CompList getRest() { return rest; }
public String toString() {
  return "(" + first + rest.toStringHelp() + ")";
}
String toStringHelp() {
  return " " + first + rest.toStringHelp();
}
static Cons twoElts =
  new Cons(new Integer(1), new Cons(new Integer(2), Empty.ONLY));
static Cons fourElts =
  new Cons(new Integer(-1), new Cons(new Integer(-2), l1));
}

Now assume that we want to modify the CompList class so that it implements the Comparable interface. The compareTo method in the Comparable interface has the following contract. Any class C that implements the Comparable interface must have an associated binary relation that is totally ordered: for every pair of objects a and b in C, either (i) a is less than b, (ii) a equals b, or (iii) a is greater than b. For any instance o of the same class as this, compareTo(o) returns (i) a negative number if this is less than o, (ii) zero if this equals o, and (iii) a positive number if this is greater than o. If o be-

longs to a different class than this, compareTo(o) throws a ClassCastException

indicating an erroneous use of the compareTo method.

In the CompList class, we can impose a lexicographic total ordering on lists. This ordering is a generalization of the familiar alphabetic ordering on strings. In such an ordering, a precedes b iff either

• a is empty and b is not,

• the first element of a precedes the first element of b, or

• the first element of a equals the first element of b and the rest of a precedes the rest of b.

Finger Exercise 1.9.3.1 Load your saved file objectList.java int the DrJava Definitions pane. Convert it to a definition of the class CompList given above. Modify this CompList class to implement the Comparable interface as described above. Include test examples in your code and run them to confirm that your program works in these cases.

1.10 Using Classes and Interfaces to Enforce Invariants

Some data objects have an associated invariant (boolean condition) which must be maintained for the object to be well-formed. For example, the elements in a sorted list must appear in ascending order. In many cases, we can use an interface to ensure that such an invariant always holds.

Consider the example that we already cited: a sorted list. Can we define an OrdList class hierarchy with subclasses Empty and OrdCons similar to the IntList class hierarchy in Section 1.7.2 that guarantees that all instances are sorted? The answer is yes, but we have to change the visible interface (members) of the class. In particular, we cannot allow clients of the OrdList type to perform new operations on the OrdCons class. To add an element to an OrdList, clients must use a method

     OrdList insert(int f)

that inserts f in proper position in this.

The OrdCons class includes a binary constructor just like IntList except for the fact that it is private, implying that no code outside of class OrdCons can use it. This visibility restriction raises a minor problem: how can we write the insert method for the Empty subclass? The binary OrdCons constructor is not accessible! The answer is to define a second constructor for the OrdCons class that takes a single int argument and initializes the rest field to Empty.ONLY.

Finger Exercises 1.10

1. In this exercise, we will build an implementation of ordered lists as a subtype of unordered lists. Load the program NewIntList.java into the Definitions pane of DrJava. This program is similar to IntList.java but formulates the IntList type as an interface rather than an abstract class. Define OrdList as an interface extending IntList interface with the method:

          OrdList insert(int i);

and narrow the output type of the empty to OrdList since the empty list is ordered! Define the class OrdCons as a subclass of Cons. The OrdCons and Empty classes implement the interface OrdList. The member fields of Cons can be private if you rely on a super call to initialize these fields in your OrdCons constructors.

2. Define a sort method for the class IntList that sorts a list converting all Cons nodes to OrdCons nodes. Test your code. Save your program as

     OrdList.java.

3. The Empty and Cons classes contain repeated code. Hoist the common code in these two classes into a separate abstract class IntListClass.

1.11 The Command Pattern

In a finger exercise in Section 1.6.3, we extended the DeptDirectory program by writing a method findPhone(String name) to look up a person’s phone number. We implemented findPhone(String name) in exactly the same way as findAddress(String name), replicating method code. A better strategy would be to implement a method

     Entry findEntry(String name)

that returns the Entry matching a given name, and then to define both findPhone

and findAddress in terms of findEntry.

In this section, we will explore a far more general technique for eliminating code replication called the command pattern. To accommodate returning different Entry fields, we will define a method

     String findField(Operation f, String name)

that takes an Operation object as an extra argument specifying which field to return. This approach mimics the familiar “code factoring” process in Scheme: repeated code patterns are abstracted into functions that take parameters that “customize” the code appropriately. In many cases, these parameters are functions.

Code factoring involving functions as parameters cannot be directly implemented in Java because methods are not values that can be passed as arguments. Some object oriented languages such as SmallTalk and Self classify methods as data values, permitting code factoring to be implemented directly. Fortunately, it is not difficult to get around this restriction by explicitly representing methods as objects. All we have to do is introduce an appropriate abstract class Operation containing a single abstract method execute( ... ) and define a separate concrete subclass of Operation for each method that we want to pass as an argument. Each concrete subclass defines the abstract method execute appropriately. In the general case, the Operation subclasses may contain fields that correspond to the free variables appearing in procedural arguments in Scheme. These free variables must be bound when the Operation is constructed, exactly as they are in a language supporting procedures as data values.

In the object-oriented design literature, this technique is called the command pattern in homage to the dominant role that imperative operations have played in object-oriented computation. Here we are using this pattern in a purely functional fashion.

To illustrate the command pattern, let us continue our DeptDirectory example. If we independently write findPhone and findAddress, they differ only in the field name used in the return expression.

     class Empty extends DeptDirectory {
       ...
       String findAddress(String name) {
         return null;
       }
       String findPhone(String name) {

         return null;
       }
     }
     class Cons extends DeptDirectory {
       ...
       String findAddress(String name) {
         if (name.equals(first.name))
           return first.getAddress();
         else return rest.findAddress(name);
       }
       String findPhone(String name) {
         if (name.equals(first.name))
           return first.getPhone();
         else return rest.findPhone(name);
       }
     }

We can “abstract out” this difference by writing a single findField method embodying the common code in the methods findPhone and findAddress. To accommodate differing choices for the returned Entry field, the method takes an Operation that performs the appropriate field extraction on the Entry. The following code includes a new mechanism for defining concrete subclasses, called anonymous classes, that we have not discussed before. We will explain anonymous classes in detail below. In this example, anonymous classes are used to generate instances of new subclasses of the interface IOperation; the static fields address and phone are bound to objects of type Operation that define the execute method as the method extracting the address and phone fields, respectively, of an Entry.

     interface IOperation {
       String execute(Entry e);  // implicity public and abstract
     }
     abstract class DeptDirectory {
       ...
       abstract String findField(IOperation c, String name);
       String findAddress(String name) {
         return findField(opddress, name);
       }

String findPhone(String name) {
  return findField(opPhone, name);
static IOperation opAddress = new IOperation() {
  // ANONYMOUS class
  public String execute(Entry e) { return e.getAddress(); }
}
static IOperation opPhone = new IOperation() {
  // ANONYMOUS class
  public String execute(Entry e) { return e.getPhone(); }
}
}
class Empty extends DeptDirectory {
...
String findField(IOperation c, String name) {
  return null;
}
}
class Cons extends DeptDirectory {
...
String findField(IOperation c, String name) {
  if (name.equals(first.name)) return c.execute(first);
  else return rest.findField(c,name);
}
}

Each brace construction

{  // ANON CLASS
  public String execute(Entry e) { return e. ...; }
}

following a new IOperation() expression above defines a unique instance of a new anonymous (unnamed) class implementing the interface IOperation. In Java, anonymous classes are simply an abbreviation mechanism. The IOperation class could have been written without anonymous classes as follows:

interface IOperation {
String execute(Entry e);
}
}

     class AddressOperation implements IOperation {
       public String execute(Entry e) {
         return e.getOffice();
       }
     }
     class PhoneOperation implements IOperation {
       public String execute(Entry e) {
         return e.getPhone();
       }
     }
     abstract class DeptDirectory {
       ...
       static IOperation opAddress = new AddressOperation();
       static IOperation opPhone = new PhoneOperation();
     }

at the cost of introducing the new class names AddressOperation and PhoneOperation.

In general, a single instance of a new class extending class (implementing interface) C can be created using the notation:

     new C(...) { ... members ...}

where C(...) specifies what superclass initialization should be performed on the instance. If C is an interface, then the argument list in C(...) must be empty. No constructors can appear in the list of members because the class is nameless and cannot be instantiated again. Any required initialization of fields inside the instance can be specified directly in the code defining the class.

If we ignore the ugly notation, an anonymous class extending the abstract class IOperation has a direct analog in Scheme that you may have recognized, namely a lambda-expression. In any situation in Scheme where it is appropriate to use a lambda-expression, you can use an anonymous class in Java! The failure to make such an identification is the single most glaring failure of most expositions on Java. Of course, they are typically written for readers with a background in C or C++ rather than Scheme or other language that support procedures as general data objects.

If an anonymous class appears inside a dynamic method, it can contain references to the fields of the enclosing class instance—akin to the free variables that can appear in Scheme lambda-expressions. The only complication is the treatment of the variable this. Since an anonymous class defines an instance of a new class, the variable this inside an anonymous class refers to the new class instance. It does not refer to the enclosing class instance. To refer to the “entire” enclosing class instance, Java uses the notation C.this where C is the name of the enclosing class.

Finger Exercises 1.11:

1. Add a map method to the IntList class that takes an IOperation and applies it to each element of this to produce a new IntList with exactly the same number of elements as this.

2. Assume that a vector

< a_{0}, a_{1}, a_{2}, ..., a_{n} >

is represented by the list

(a_{0} a_{1} a_{2} ... a_{n}).

where the coefficients a_{i} are objects of type Double. Add a method

    double norm()

to IntList computes the norm by this.v by squaring the elements, adding the squares together and taking the square-root of the sum. You can compute the vector of squares using map and the define a method double sum() to compute the sum.

3. Assume that a polynomial

a_{0}x + a_{1}x + a_{2}x^{2} + ... + a_{n}x^{n}

is represented by the list

(a_{0} a_{1} a_{2} ... a_{n})0

where the coefficient’s a_{i} are objects of type Double. Write a method

    double eval(IntList p, Double x)

to evaluate the polynomial at coordinate x. Use Horner’ rule asserting that

a_{0} + a_{1}x + a_{2}x^{2} + ... + a_{n}x^{n} = a_{0} + x · (a_{1} + x · (a_{2} + ...x · a_{n}))

Remember to use the structural design pattern for processing lists (and by association polynomials as we have represented them).

1.11.1 Static Members of Classes

In addition to (instance) members, a Java class can include static members that are attached to the class rather than instances of the class. We have already seen how static final fields provide a simple way to define constants.

The static members of a class are not included in the template used to create class instances. There is only one copy of a static field for an entire class—regardless of how many instances of the class are created (possibly none). Similarly, the code in a static method cannot refer to this or to the fields of this because there is no class instance to serve as the receiver for such an access./footnoteOf course, a static method can invoke an instance method (or extract an instance field) of class if it explicitly specifies a receiver for the invocation.

Static methods are useful because we occasionally need to write methods where the primary argument is either a primitive value or an object from a class that we cannot modify. For example, the library method Integer.toString(int i) converts an int to the corresponding String. Since an int is not an object, there is no int class to hold such a method.^{7} Consequently, the Java library

provides a static method toString(int i) in the class Integer.

Similarly, an operation

public String squeezeWhiteSpace(String s);

that returns a String identical to s with all spaces and tabs removed should be expressed as a static method because the String class cannot be modified or extended.

Finally, all operations on arrays must be expressed in static (procedural) form because array types do not have conventional class definitions; they are built-in to the Java virtual machine. We will discuss arrays in Chapter 2 when we address imperative programming in Java.

Finger Exercise 1.11.1.1: In DrJava, try typing the followin expressions in the Interactions pane:

     Integer.toString(7)
     Boolean.toString(true)
     Math.abs(-10.)
     Integer.toString(7) + Boolean.toString(true)
     7 + Boolean.toString(true)

Why do the last two expressions return the same result? Java automatically inserts the appropriate toString conversions on the argument of the + operator.

7Of course, any static method can be converted to an instance method in some class, but the conversion is gratuitous since the static method code ignores this.

This conversion process, however, only works when Java can determine that you are using + to denote string concatenation rather than addition. Try evaluating

     7 + true
     7 + 7
     true + true
     Integer.toString(7) + 7
     Boolean.toString(true) + true

If one of the arguments to + is a string, then Java interprets + as the string concatenation operator. One way to force this conversion is to include the empty string "" in a concatenation sum. Try evaluating

     "" + 7 + true

Static members are really familiar notions from Scheme and C in disguise; static fields behave exactly like ordinary global variables in Scheme or C and static methods behave like ordinary Scheme or C procedures. Java forces these global variables and procedures to be attached to classes in the guise of static fields and static methods.

Since static fields and methods (other than static final constants) are outside the realm of object-oriented computation, we will rarely use them. Finger Exercise 1.11.1.2: Define a class MathUtil that includes a method

     public static int max(int[] x)

that computes the maximum element of the array x. If the array is empty, what should you return?

Optional Finger Exercise 1.11.1.3 Add a method

     public static int map(IFunction f, int[] x)

to the MathUtil class where IFunction is the interface

     interface IFunction {
       int apply(int y);
     }

The map method takes an IFunction f and an int[] x and applies f to each element of the array x, returning a new array of the same length as s as the result. map does not modify the array x.

1.11.2 Complete Java Programs

Every complete Java program must contain a root class where execution can begin. A root class must contain a main method defined with the header

     public static void main(String[] args)

To execute a Java program, a user must identify the name of the root class for the program. In most Java computing environments, the command required to execute the Java program with root class C is simply

     java C

Note that a class may contain a main method even if it is not intended to be the root class of an actual program. Of course, evaluating the DrJava expression

     Motto.main(null);

is equivalent to executing this conssoled. For the sake of programming convenience, the DrJava Interactions pane accepts the command line syntax as well as the explicit invocation of the main method shown above.

When execution begins in the main method of a root class, no instances of the root class exist. In fact, most Java programs never instantiate the root class because it serves as a framework for organizing program execution rather than a definition of a new form of data. Java classes really serve two distinct purposes: defining data objects and organizing static methods and variables.

The args parameter of the main method in the root class is used to pass command line options to the Java program. We will not use this feature in this monograph.

A Complete Java Program

The following simple program

     class Motto {
       public static void main(String[] args) {
         System.out.println("Java rules!");
       }
     }

prints the String output

Java rules!

and stops when the Java program with root-class Motto is executed. Before any Java program can be executed, it must be compiled into a form suitable for machine execution called a class file Motto.class. The name of the class file generated for a class C is simply C.class.

In DrJava, the Compile button applies the compiler to each open Java files and write out the resulting class files to the file system—assuming that the files did not contain any syntax errors. Each class file is stored in the same direcoty as the corresponding source (.java) class.

Finger Exercise 1.11.2.1: Using DrJava, define the Motto class given above and execute it as program.

1.12 Loose Ends

We have nearly finished covering the core “functional” subset of Java. Only a few loose ends remain. They include:

• local variables;

• casts and static type checking;

• exceptions; and

• name and method overloading

1.12.1 Local variables

Method parameters in Java play exactly the same role as function parameters in Scheme. They are local variables bound to the corresponding argument values. They are destroyed when the method returns.^{8} Java also provides a mechanism, akin to Scheme’s local construct, for introducing local variables to perform a subcomputation without invoking a helper method. In Java, local variable definitions can be inserted as statements in any statement sequence (such as a function body or compound statement). Each such variable is accessible only in program text between its definition and the end of the statement sequence.

For example, it is convenient in writing test methods to introduce local variables to hold test data values. The following testDeptDirectory method for the class DeptDirectory relies on this technique.

8In Scheme, this statement is false in general; local variables continue to exist as long as a “closure” referring to them exists.

     class TestDeptDirectory {
       public static void testDeptDirectory(){
         Entry cork = new Entry("Corky","DH 3104","x 6042");
         Entry matthias = new Entry("Matthias","DH 3106","x 5732");
         Entry ian = new Entry("Ian","DH 3102","x 3843");
         DeptDirectory dd = new Cons(ian, new Cons(cork,
           new Cons(matthias, Empty.ONLY)));
         System.out.println("ian " + dd.findAddress("ian") + " " +
           dd.findPhone("ian"));
         System.out.println("cork " + dd.findAddress("cork") + " " +
           dd.findPhone("cork"));
         System.out.println("matthias " + dd.findAddress("matthias")
           + " " + dd.findPhone("matthias"));
       }
     }

Java imposes an important restriction on the use of local variables in anonymous classes. Any local variable mentioned in an anonymous class definition must be declared final. In practice, this is not a significant restriction. In fact, most attempted uses of non-final local variables in anonymous classes correspond to errors in program logic. In any case where you need to mention a non-final local variable in an anonymous class, you can simply introduce an extra final variable with the required value. Ironically, this transformation often converts a logically incorrect program in to a correct one! We will revisit this issue in later in the monograph in connection with programming graphical user interfaces.

1.12.2 Casts and Static Type Checking

In Scheme every primitive operation dynamically checks that its arguments have the appropriate form (type) as it executes. If an operation is applied to data of the wrong form, Scheme aborts execution and prints an error message, much as Java does when an exception is thrown and not caught.

Java also performs some argument checking during program execution (run-time), but most argument checking is done statically by the compiler before a Java program executes. A Java compiler enforces a syntactic discipline on program text called static typing. The compiler uses a collection of type-checking rules to determine whether a program is well-typed or not. Programs that are not well-typed are rejected with an explanation of which rules were broken.

The type-checking rules embody simple “common sense” inferences and consistency checks. The rules assign a type to every program expression and subsequently check that these type assignments are consistent. A Java compiler assigns types to program expression as follows:

1. every program constant has a type specified by the language definition;

2. every program variable (field, parameter, or local variable) has the type declared in the program;

3. each method invocation has the declared return type of the method;

4. each application of an arithmetic operator (e.g., +, *) has the return type stipulated by a table in the language definition;^{9}

5. each application of a relational operator and instanceof test has type

   boolean;

6. each conditional expression

test ? consequent : alternative

has the more general type of the consequent type and alternative;^{10} and 7. the type of any cast expression

T e

is T. Given these type assignments, a Java compiler checks their consistency by

enforcing the following rules:

1. the type of the receiver of a field selection or method invocation includes the specified field or method in its signature (a list of the member headers for the class or interface);

2. the type assigned to each argument expression in a method invocation is a subtype of the declared type of the corresponding parameter;

3. the type of the right-hand side of an assignment is a subtype of the type of the left-hand-side; and

4. the type of each return expression in a method body is a subtype of the declared return type of the method.

9The return types of arithmetic operators generally depend on the types of their arguments. For some operator applications, the table immediately reports a type error.

10If neither arm of the conditional expression has a more general type, the program is not well-typed.

Note that Java type checking rules do not capture the logical consequences of instanceof tests. As a result, Java program text often must include apparently redundant casting operations in code following an instanceof test.

This phenomenon is illustrated in the following simple example. Consider the following method which could be added to the IntList class above.

     static Object first(IntList l) {
       if (l instanceof Cons) return ((Cons) l).first;
       else throw
         new ClassCastException("first requires a non-Empty IntList");
     }

In the method, all occurrences of the parameter l have the same type, namely IntList as declared in the method header. The

     l instanceof Cons

test has no effect on type-checking. As a result, the occurrence of l preceding the field extraction operation .first must be explicitly converted to type Cons using the casting operation (Cons) written as a prefix in front of l. Since the field .first is not defined in the abstract class IntList, the definition of the method first does not type check if the casting operation (Cons) is omitted.

Applying a casting operation ( T ) to a Java expression e of some static object type U has consequences for both program execution and compilation. First, it inserts a run-time check to confirm that the value of e belongs to the type T as claimed by the casting operation. Second, it converts the static type of the expression e from U to T.

1.12.3 Exceptions as Errors

Some operations on data are inherently partial. For example, there is no mathematically sensible definition for the result of integer division by zero. Similarly, there is no sensible definition for the first element of an empty sequence. Java provides an elegant mechanism for coping with this problem. Program operations can “throw” a “run-time error” condition called an unchecked exception that aborts program execution and prints a diagnostic error message and a list, called a traceback, of the stack of pending method calls. For example, attempting to divide any int by 0 generates an ArithmeticException. Exceptions are ordinary Java data objects descended from the built-in class called Exception. Unchecked exceptions are descendants

of the class RuntimeException extending Exception.

In Java, any method can “throw” an exception simply by executing the statement

throw e;

where e is any expression that evaluates to an object of type Exception. The classes Exception and RuntimeException both include a zero-ary constructor and a unary constructor. The latter takes an error message String as an argument. The string argument is printed as a diagnostic message if it is provided.

The following Java code implements “functional” lists of objects:

abstract class List {
abstract Object first();
abstract List rest();
abstract String toStringHelp();
// List -> String without enclosing parentheses and leading blanks
}
class Empty extends List {
Object first() {
  throw new
    ClassCastException("first requires a non-Empty List");
}
List rest() {
  throw new
    ClassCastException("rest requires a non-Empty List");
}
public String toStringHelp() { return "()"; }
String toStringHelp() { return ""; }
}
class Cons extends List {
Object first;
List rest;
Cons(Object f, List r) {
  first = f;
  rest = r;
  }
Object first() { return first; }
List rest() { return rest; }

       public String toString() { return "(" + first + rest.toStringHelp() + ")"; }
       String toStringHelp() { return " " + first + rest.toStringHelp(); }
     }

The class ClassCastException is a built-in Java class extending RuntimeException that other built-in Java classes use to signal improper method applications where this or some other argument belongs to the wrong class.

Exception objects that do not belong to the type RuntimeException are called checked exceptions. We will discuss how they are used later in this monograph.

1.12.4 Name and Method Overloading

In Java, the same name can simultaneously be used for a local variable or method parameter, different fields of the same object, and different methods of the same object. Java uses context and type information to determine the meaning of a name. For example, many Java programmers would write the constructor for the class Cons above as follows:

       Cons(Object first, List rest) {
         this.first = first;
         this.rest = rest;
         }

In this example, the names first and rest each have two different meanings. Java resolves potentially ambiguous uses of names by using the innermost (most local) definition of the name. Hence, the prefix this. is required to distinguish the first and rest fields of this from the constructor parameters of the same name.

While the use of the same name for different kinds of program entities is widely accepted as good programming practice, the use of one name for several different fields or methods is more controversial. A Java subclass B can introduce a field with exactly the same name n as a field in its superclass A. The inherited field is not overridden (what would overriding mean in this context?); it is merely “shadowed”. When the name n appears in a method of B, its meaning depends on the type of the receiver. If the receiver is this, then the new field n introduced in B is meant. But if the receiver has type A rather than B, then the old field n introduced in A is meant. Warning: duplicate field names can be the source of insidious program errors that are very difficult to diagnose. For this reason, we strongly recommend against using them.

Duplicate method names are less controversial but can still be dangerous. In a class, Java permits the same method name to be used for different methods as long as their argument lists do not identical the same length and same types. The practice of defining more than one method in a class with same name is called method overloading. Java resolves overloaded method names using the types of the argument expressions. When the Java compiler encounters a method invocation involving an overloaded method, it determines the types of the method arguments and uses this information to select the “best” (most specific) match from among the alternatives. If no best method exists, the program is ill-formed and will be rejected by the Java compiler.

We urge restraint in using the same name for different methods involving the same number of arguments. Since static type information is used to resolve which method is meant, program errors may be difficult to find because the programmer may not infer the correct static type when reading the code.

1.13 The Visitor Pattern

The composite and interpreter patterns enforce an elegant discipline for writing object-oriented functional programs. But this elegance comes at a price: every time we need to define a new operation we must modify every class in the composite class hierarchy to add a new method. As a result, program modification involving new methods is painful. Moreover, we cannot use these patterns to define new operations if we do not control the source code for the composite class. Without access to the source code, the only way we can define new methods is to write them in procedural style as static methods in another class, losing all the benefits of object-oriented organization.

In this section, we introduce a design pattern, called the visitor pattern, that completely and elegantly solves the problem of adding new operations to composite class hierarchies without modifying the text of either the composite class or its variants.

Before delving into the technical definition of the visitor pattern, we present a motivating example: an interpreter for arithmetic expressions.

1.13.1 Interpreting Arithmetic Expressions

An ArithExpr is either:

   • Const(c)

• Sum(left right),

• Prod(left,right), or

   • Neg( left)

where c is an int and left and right are ArithExprs.

As before, to represent this data type in Java, we employ the composite pattern. We define an abstract class to represent the union of these types, and four concrete subclasses, one for each of the different variants.

abstract class ArithExpr {
}
class Const extends ArithExpr {
/* fields */
private int value;
/* constructor */
Const(int v) { value = v; }
/* getters */
int getValue() { return value; }
/* toString */
public String toString() { return Integer.toString(value); }
}
}
class Sum extends ArithExpr {
/* fields */
ArithExpr left, right;
/* constructor */
Sum(ArithExpr l, ArithExpr r) { left = l; right = r; }
/* getters */
ArithExpr getLeft() { return left; }
ArithExpr getRight() { return right; }
/* toString */
public String toString() {
  // here we have recursive calls to toString,
  // as we would expect in an inductively-defined type
  return "(" + left + "+" + right + ")";
}
}

The two remaining classes, Prod and Neg, are defined similarly.

Next we need a way to evaluate our expressions. First, let’s try defining an eval method that returns a constant for any ArithExpr.

We add the abstract method

     abstract Const eval();

to the ArithExpr class. We could easily define eval to return an int, but we’ve chosen to return a Const instead because it buys us a little flexibility later.

In the Const class, we add a concrete version of the abstract eval method:

       Const eval() { return this;  }

But now we encounter a minor problem: to evaluate products and sums, we need to be able to multiply or add two instances of Const. Multiplication and addition are not defined for instances of Const, but they are defined for the int values embedded inside instances of Const. Thus, we can use the accessor getvalue() to retrieve the value of a Const. To the class Prod, we add the method

       Const eval() {
         return new Const((left.eval().getValue()) * (right.eval().getValue()));
       }

The eval methods for Sum and Neg are defined similarly.

Let us amplify this example by adding variables as a syntactic category to ArithExpr. Writing down our revised data type in a shorthand form, we have

     ArithExpr ::= Const(int)
                  | Sum(ArithExpr, ArithExpr)
                  | Prod(ArithExpr, ArithExpr)
                  | Neg(ArithExpr)
                  | Var(String)

where the variant Var represents variables. The String field in a Var is the variable’s name.

This notation is merely shorthand for the follow ing prose. An ArithExpr is either:

   • Const(c ),

• Sum(left,right ),

• Prod(left,right) ),

• Neg(left), or

   • Var(s)

where c is an int, left and right are ArithExprs, and s is a String.

In order to evaluate expressions including variables, we introduce environments, which store collections of bindings. A binding is simply a pair containing a variable name and a value. We will also have to modify eval so that its signature becomes

          Const eval(Environment env)

We can implement Environments using functional lists of string-value pairs. The details of the implementation are left as an exercise. (Hint: look back at the departmental directory example.)

The definition of eval will have to change accordingly for each of the existing concrete subclasses, but only the Var subclass will actually make use of the environment parameter. For example, Sum.eval will become

       Const eval(Environment env) {
         return new Const(left.eval(env).getValue() + right.eval(env).getValue());
       }

The parameter env is not used directly in the eval code for Sums, but it is passed to the recursive calls to eval in case there is a Var further down in the expression. It is only in class Var that we need to use the environment parameter to look up the value of a variable:

     class Var extends ArithExpr {
       /* fields */
       private String name;
       /* constructor */
       Var(String n) { name = n; }
       /* accessors */
       public String getName() { return name; }
       /* toString */
       public String toString() { return name; }
       Const eval(Environment env) { return env.lookup(name); }
     }

Here env.lookup(name) fetches the Const value associated with name in the environment env (if there is no entry for name, lookup should raise some kind of exception).

Having to pass the environment as a parameter to all of the eval methods, when it is directly used in only one of them, is clumsy. As we shall see, there is a different way to implement expression evaluation that avoids this problem.

Finger Exercise 1.13.1.1 Finish writing the program to evaluate arithmetic expression given in this section.

1.13.2 Openness in Data Design

Recall our definition for an arithmetic expression without variables:

ArithExpr := Const(int) | Sum(ArithExpr, ArithExpr) ....

Our implementation of this data definition using the composite pattern would be more robust and more flexible if we could define new operations on ArithExprs without modifying any existing code. Fortunately, there is a clever design pattern called the visitor pattern that lets us do this. The idea underlying the visitor pattern is to bundle the methods defining the new operation for each concrete subclass together in a new class called a visitor class. An instance of such a class is called a visitor.

First, we will define a new interface IVisitor that specifies what methods must be included in every visitor class for ArithExpr:

     interface IVisitor {
       int forConst(Const c);
       int forSum(Sum s);
       ...
     }

Notice that each method takes an instance of the class that it processes. This argument, called the host, is needed to give it access to all the information that would be available through this if the method were defined inside that class, e.g., the values of the object’s fields returned by accessors.

Now we will create a new concrete class EvalVisitor to hold all the methods for evaluation of an ArithExpr:

     class EvalVisitor implements IVisitor {
       int forConst(Const c) {
         return c.getValue();
       }
       int forSum(Sum s) {
         return s.left().accept(this) + s.right().accept(this);
       }
       ...
     }

We need to install a hook in each subclass of ArithExpr to execute the corresponding visitor method. The hook is a new method, accept, which takes a visitor as an argument and calls the appropriate method in that visitor.

     abstract class ArithExpr {}
       abstract int accept(IVisitor v);
     }
     class Const {
       ...
       int accept(IVisitor v) {
         return v.forConst(this);
       }
     }
     class Sum {
       ...
       int accept(IVisitor v) {
         return v.forSum(this);
     }
     ...

To evaluate an arithmetic expression, we simply call

     a.accept(new EvalVisitor())

If we wish to add more operations to arithmetic expressions, we can define new visitor classes to hold the methods, but there is no need to modify the existing subclasses of ArithExpr.

Notice that, since a visitor has no fields, all instances of a particular visitor class are identical. So it is wasteful to create new instances of the visitor every time we wish to pass it to an accept method. We can eliminate this waste by using the singleton design pattern which places a static field in the visitor class bound to an instance of that class.

     class EvalVisitor {
       static ONLY = new EvalVisitor();
       ...
     }

Then, instead of

     accept(new EvalVisitor()),

we may simply write

     accept(EvalVisitor.ONLY).

Another elegant way to define visitors is to define each visitor as an anonymous class. Since an anonymous class definition defines only one instance of the new class, it produces results similar to the singleton pattern. The principal difference is that the new class has no name; the unique instance must be bound to a local variable or field declared in the enclosing program text.

Recall that an anonymous class has the following syntax:

new className(arg_{1}, ..., arg_{m}) { member_{1}; ...;member_{n} }

In most cases, the class className is either an abstract class or an interface, but it can be any class. The argument list arg_{1}, ..., arg_{m} is used to call the constructor for the class className; if className is an interface, the argument list must be empty. The member list member_{1}; ...;member_{n} is a list of the member definitions for the new class separated by semicolons.

For example, to create an instance of a visitor that evaluates an arithmetic expression, we write:

     new IVisitor() {
       int forConst(Const c) {...}
       int forSum(Sum s) {...}
       ...
     }

Since we generally want to use a visitor more than once, we usually bind the anonymous class instance to a variable, so we can access it again! The statement:

     visitor ev = new IVisitor() {
                    int forConst(Const c) {...};
                    int forSum(Sum s) {...};
                    ...
                    };

binds the variable ev to our anonymous class instance.

Finger Exercise 1.13.2.1 Convert your solution to exercise 1.13.1.1 to a visitor-based implementation using the IVisitor interface given in the subsection.

1.13.3 Polymorphic Visitors

In our application of the visitor pattern above, the for methods of the visitor classes and the accept methods for the ArithExpr classes returned values of type int. This convention is acceptable as long as all the computations we ever want to perform over ArithExprs have integer results. But if not, we are forced to declare a new interface visitor type and new accept methods for each distinct result type. Since the whole point of the visitor pattern is to avoid having to modify a data type every time we wish to perform some new computation over it, we have a potential problem.

We can address this problem by redefining the for and accept methods so that they return a more general type. Before getting into the details, let’s step back and give visitors a more general definition.

A visitor is an object containing

• a concrete description of an operation on composite data, with a separate method for processing each alternate form of the data; and

• the argument values required by the operation.

The properties common to all visitors for a particular composite type are collected in corresponding the visitor interface. Any arguments that are processed by a particular visitor are typically stored in fields declared in the corresponding visitor subclass to avoid compromising the generality of the interface. A typical visitor interface has the form

     interface Visitor {
       // a "for" method for each concrete
       // subclass of the visited class
       Object forC1(...);
       .
       .
       .
       Object forCn(...);
     }

We use Object as the return type of the for methods so that we can accommodate concrete visitor classes that produce almost any type of result. This convention exploits the polymorphism inherent in class inheritance: every object belongs to all of the types associated with its superclass, super-superclass, etc. Since the class Object is perched at the root of the class hierarchy, all objects belong to the type Object, and support the operations defined in class Object. The only types that do not belong to Object are the primitive types like int. Fortunately, we can get around this by using the corresponding wrapper classes (e.g. Integer instead of int). In the event we want a visitor operation to have a void return type, we can either return the null reference or the only instance of a singleton class VoidType.^{11}

Let us return to our ArithExpr example. We can generalize our visitor class as follows:

     abstract class AE { // AE is short for ArithExpr
       abstract Object accept(IVisitor v);
     }
     class Const extends AE {
       /* fields */
       int value;
       /* constructor */
       Const(int v) { value = v; }
       /* getters */
       int getValue() { return value; }
       /* toString */
       public String toString() {
         return Integer.toString(value);
       }
       Object accept(IVisitor v) {
         return v.forConst(this);
       }
     }
     . . .

The code for the Sum and Prod ares nearly identical except for the fact that the accept methods in those classes respectively invoke the forSum and forProd, methods of the Visitor argument v. They all have Object as their return type.

Finger Exercise 1.13.3.1 Convert your solution to exercise 1.13.2.1 to use the polymorphic version of the IVisitor interface.

11The Java libraries include a Void class, but it cannot be instantiated, so we must define our own VoidType.

1.13.4 Polymorphic Visitors with Arguments

To make our visitor example more interesting and realistic, let us include variables in the type ArithExpr.

class Var extends ArithExpr {
String name;
Var(String n) {
  name = n;
}
public String toString() {
  return name;
}
Object accept(Visitor v) {
  return v.forVar(this);
}
}

Then the visitor interface for ArithExprs has the form:

interface IVisitor {
Object forConst(Const c);
Object forSum(Sum c);
Object forProd(Prod p);
Object forVar(Var v);
}

The concrete visitor class that implements expression evaluation is:

class EvalVisitor implements Visitor {
Env env;  // an environment for looking up variables
EvalVisitor(Env e) { env = e; }
Object forConst(Const c) { return c; }
Object forSum(Sum s) {
  return new Const( ((Const)s.left.accept(this)).value +
                    ((Const)s.right.accept(this)).value );
}
Object forProd(Prod p) {
  return new Const( ((Const)p.left.accept(this)).value) *
                    ((Const)p.right.accept(this)).value));
}
Object forVar(Var v) { return env.lookup(v.name); }
}

The environment env, which was an explicit parameter of the eval method in our method-based implementation for evaluation, is now a field of the visitor. As before, it is directly used only for evaluating instances of Var, but now we don’t need to explicitly pass the environment through method argument lists.

Since we are programming in a functional style, the forConst method need only return its argument as the result, rather than allocating a copy. The forSum and forProd methods are mostly straightforward, evaluating the subexpressions first and combining the results. The only subtlety is that since accept now returns an instance of Object rather than an int, we need to perform an explicit type cast to get the values for the left and right subexpressions. For example, to obtain the value for the left subexpression in a Sum, we have

          ((Const)s.left.accept(this)).value

The expression

     s.left.accept(this)

computes a Const whose value field is the value of the expression s.left. But the declared return type for accept is Object, and since an Object has no field named value, we cannot extract the value directly. Since we know that the Object is in fact a Const, we can insert an explicit type cast to Const, and then extract the value.

The forVar method looks up the value of the variable in the current environment. The environment is passed in when an EvalVisitor is created, and is presumably given bindings for the existing variables beforehand.

Finger Exercise 1.13.4.1 Extend your solution to exercise 1.13.3.1 to support variables. You will need to define an Environment class in addition to the ArithExpr classes.

1.14 Unusual Situations versus Runtime Errors

In this section, we study Java exceptions, a language construct to process unusual situations and run-time errors. In the process, we will identify the limitations of checked exceptions.

Many production programs have to detect erroneous input, report it to the user, and in some cases recover to handle more input. Since handling erroneous input deviates from the expected flow of program control, Java provides a mechanism called exception handling that is tailored for this purpose. To be more specific, let us consider an example based on the arithmetic expression evaluator from the previous section.

1.14.1 A Motivating Example

Recall the visitor class for evaluating arithmetic expressions with variables:

     class evalVisitor implements Visitor {
       Env e;
       Object forConst(Const c) {return c; }
       Object forSum(Sum s) {
         Const l = (Const)(s.left.accept(this));
         Const r = (Const)(s.right.accept(this));
         return new Const(l.value + r.value);
       }
       Object forVar(Var v) { return env.lookup(v.name); }
     }

The casting operations (Const) ... in the body of the forSum method are required by the Java type checker. The Java type system is too imprecise to determine that the recursive invocations of the visitor in forSum will never return any value other than a Const. The Java type system simply uses the declared return type for a method as the type of invocation of that method.

You might wonder why the designers of Java adopted such a simple, imprecise type system. Precise type systems have two crippling disadvantages. First, they perform a complex inference process that is difficult to understand. If a programmer makes a type error, it is difficult to determine what program revisions are required to satisfy the type checker. Second, precise type inference is expensive. The time complexity of very precise type checking (depending on the specific algorithm) may grow with the square or cube of program size or worse.

Now let us augment Arithmetic Expressions with a variable binding operator let. For example, we might want to evaluate the expression:

     let x = 17 in x + y

in an environment where y is bound to 10 to produce the value 17. This extension is reflected in the definition for data type ArithExpr by adding a Let form to the list of variants:

     ArithExpr ::= Const(int)
                 | Sum(ArithExpr, ArithExpr)
                 | ...
                 | Let(Var, ArithExpr, ArithExpr)

Similarly, in the object-oriented implementation of ArithExpr, we must add the variant class

class Let extends ArithExpr {
Var bindVar;
ArithExpr bindExpr;
ArithExpr body;
...
Object accept(Visitor v) { v.forLet(this); }
}

To define operations for our generalized Arithmetic Expressions using visitors, we need to define a new visitor interface with for methods for all of the variants:

interface IVisitor {
Object forConst(Const c);
Object forSum(Sum c);
Object forProd(Prod p);
Object forVar(Var v);
Object forLet(Let l);
}

Since the evaluation of a Let form involves extending the environment, let us write the code for manipulating environments:

class Env {
abstract Const lookup(String name);
}
class Empty extends Env {
Const lookup(String name) {
  return null;
}
}
class Cons extends Env {
String firstName;
Const firstVal;
Env rest;
Cons(String name, Const val, Env env) {
  firstName = name;
  firstVal = val;
  rest = env;
}
Const lookup(String name) {
  if (name.equals(firstName)) return firstVal;

         else return rest.lookup(name);
       }
     }

To evaluate generalized Arithmetic Expressions, we must define an appropriate concrete class implementing IVisitor:

     class EvalVisitor implements Visitor {
       Env env;
       Object forConst(Const c) { return c; }
       ...
       Object forLet(Let l) {
         Const bv = (Const) l.bindExpr.accept(this);
         return l.body.accept(
           new evalVisitor(new Cons(l.bindingVar.name, bv, env));
       }
     }

Notice the cast to Const in the definition of bv. Java requires this cast operation because the declared return type of accept is Object. But the value we bind to a variable to in a let expression must be a Const. What happens if we try to evaluate a Var that is not bound in the environment? To explain how Java will behave in a such a situation, we need to discuss Java exceptions in more depth.

Finger Exercise 1.13.5.1 Extend your solution to exercise 1.13.4.1 to support the let construct. You will need to define an Environment class in addition to the ArithExpr classes.

1.14.2 Using Java Exceptions

A Java exception is an object of type Exception, which is a built-in Java class. There are two basic forms of exceptions that can occur during Java program execution:

1. Unchecked exceptions, which extend the class RuntimeException, usually signal a program coding error.

2. Checked exceptions, which extend the class Exception but not the class RuntimeException, signal unusual but legal conditions that require deviation from the normal flow of control.

When an EvalVisitor encounters a Var not bound in Env, it has detected an error in the input expression. If the Arithmetic Expression evaluator is being used in a larger program that can prompt the user for corrected input, then such an input error should be handled as part of valid program execution. It does not indicate a coding error in the Arithmetic Expression evaluator. Hence, when an EvalVisitor encounters an unbound exception, it should throw a checked exception, which the larger program can intercept and interpret, printing an error message such as

     I'm sorry, that's not a valid expression

and prompt the user for corrected input with a message like

     Please enter a valid expression:

Java requires an explicit throws clause in the header for any method that can generate a checked exception, directly or indirectly by invoking another method. The EvalVisitor class defined above will return null or generate a

     NullPointerException

if it encounters an unbound variable. The lookup method will return null as the value of an unbound variable. Any subsequent attempt to use such a value as a Const (e.g., in computing a Sum) will generate a

     NullPointerException

. Since this exception is unchecked, it does not need to be declared in throw clauses.

If we rewrite lookup to throw a checked UnboundException instead of returning null, the change has a dramatic impact on the rest of the program. The revised code appears below:

     class UnboundException extends Exception {
       UnboundException(String name) {
         super("Variable " + name + " is unbound");
         String varName = name;
       }
     }
     class Env {
       abstract Const lookup(String name) throws UnboundException};
     }
     class Empty extends Env {
       Const lookup(String name) throws UnboundException} {
         throw UnboundException(name);

}
}
class Cons extends Env {
String firstName;
Const firstVal;
Env rest;
Cons(String name, Const val, Env env) {
  firstName = name;
  firstVal = val;
  rest = env;
}
Const lookup(String name) throws UnboundException} {
  if (name.equals(firstName)) return firstVal;
  else return rest.lookup(name);
}
}
class EvalVisitor implements IVisitor {
Env env;
Object forConst(Const c) { return c; }
Object forSum(Sum s) throws UnboundException { ... };
Object forProd(Prod p) throws UnboundException { ... };
Object forVar(Var v) throws UnboundException { ... };
Object forLet(Let l) throws UnboundException { ... };
}
}
interface IVisitor {
Object forConst(Const c) { return c; }
Object forSum(Sum s) throws UnboundException;
Object forProd(Prod p) throws UnboundException;
Object forVar(Var v) throws UnboundException;
Object forLet(Let l) throws UnboundException;
}

The preceding code cleanly handles input errors, but it pollutes the signatures of nearly all of the for methods in the class Visitor and its descendants. In this case, an unchecked exception is preferable. The code for this variation is identical the code above except for the extends clause in the definition of class UnboundException and the elimination of all throws clauses in for methods.

Checked exceptions and polymorphic programming do not mix well. Consequently, it should not be surprising that the Java libraries use unchecked exceptions far more than they use checked exceptions. Our advice is use checked exceptions to signal unusual conditions in code that does not involve polymorphism. If polymorphism is present, use unchecked exceptions instead.

1.14.3 Exception Handling

If the evaluation of a Java statement generates an exception, that exception can be caught (consumed) and processed by an appropriate handler associated with the statement. The handler typically restores the program to a recoverable state.

The Java construct for associating exception handlers with program statements is called a try-catch block. It has the following syntax

     try {

statement_{1};
...
statement_{m};
}
catch(ExceptionT ype_{1} e) { handler_{1} }
... catch(ExceptionT ype_{n} e) { handler_{n} }

This statement associates the handlers described by handler_{1}, ..., handler_{n} with the statements in the sequence

statement_{1}; ...statement_{m};

If any statement in this sequence generates an exception e, it is matched against the types ExceptionT ype_{1}, ..., ExceptionT ype_{n} in order. The handler code for the first type containing e is executed and the exception e is consumed. Then execution resumes at the statement immediately following the try-catch block. If no type in the catch clauses matches the generated exception e, program execution searches back up the chain of pending method calls until it finds the next “youngest” pending try-catch block with a matching catch clause. If this search exhausts the chain of pending method calls, the Java Virtual Machine prints an error message, a traceback of method calls from the point of exception generation, and aborts program execution.

Since the exception is consumed by the matching catch clause, the program text surrounding the try-catch block does not see the exception. More generally, if a try-catch block includes a catch clause for exception type E, the surrounding program text can never see an exception of type E emerge from the try block. As a result, if the exception type E is checked, the containing method need not declare a throws clause for type E unless other code, unshielded by an enclosing try-catch block, can generate exceptions of type E.

A try/catch block can optionally be followed by a finally clause

     finally {
     cleanUp

}

This clause is always executed when the try block terminates (including any required handler code). It typically contains code to perform any clean up that might be necessary after executing the statement sequence enclosed by the try/catch block, regardless of whether the statement sequence generates an exception. In other words, the finally clause is always executed (assuming the statement sequence does not loop infinitely). The most common usage of the finally clause is to release a explicitly allocated resource such as an opened file.

We illustrate the use of the try-catch-finally block by the following example. Suppose the main method of the ArithExpr class contains test code for expression evaluation. The main method can protect itself from the exceptions eval might throw by enclosing its invocation in a try-catch block as follows:

     ArithExpr a = ...;           // set the expression
     Env.e  = ...;                // set up the environment
     ...
     try {                        // try the operation that might fail
       ...
       result = a.eval(e);
       ...
     }
     catch (UnboundException u) { // handle any error that occurred
       System.out.println ("Unbound exception thrown: " + u.varName +
                           " is undefined.");
     }
     finally {                    // optional clean-up
       ...
     }
     ...                          // continue with processing

The operation that might throw the exception, eval, is placed inside a try block, and the handling code for the exception appears in a following catch block. There may be more than one catch block, each for a different type of exception. The type and a name for the caught exception is declared at the top of the catch block, and the code in the block can use the exception’s fields to perform some kind of error recovery. Here this recovery is simply printing a message that describes the problem.

The finally clause would be useful if the try block read each expression a from a separate file and eval could throw other exceptions besides UnboundException. In this case, the finally clause could close the file regardless of how execution of the try block terminated.

If the code in the try block raises no exceptions, or raises an exception that matches a catch clause, execution continues immediately after the try-catch block and optional finally clause. Of course, the finally clause is executed as part of the try-catch block.

Object-Oriented Data Structures

The traditional programming curriculum focuses far more attention on efficient algorithms than on clean program design. Nearly all existing books on data structures including those that use C++ and Java fail to apply good object-oriented design principles in presenting interesting data structures. In this chapter, we will show how to formulate some of the most common data structures in an object-oriented design framework. Enjoy!

2.1 Sequences

The first data structures that we will discuss are the common representations for sequences of elements. A sequence is an ordered collection s_{0}, s_{1}, s_{2}, ..., s_{n}, n >= 0 of data objects drawn from some base data type D. Sequence representations can be classified based on the set of operations that they support. In Java, it is convenient to describe these various sets of operations as Java

interfaces.

We have already used a common representation for sequences, namely functional lists, in many of our examples of functional programming in the preceding chapter. Before we explore general representations for sequences, we need to discuss a particular formulation of sequences called arrays, which are built-in to Java as they are in nearly all other programming languages.

2.1.1 Arrays

An array is an indexed sequence s_{0}, s_{1}, s_{2}, ..., s_{n}, n >= 0 of data values of fixed length. In contrast to more general representations for sequences an array object cannot grow or shrink. It has a specific length n >= 0 when it is created and retains this length during its entire lifetime. Java arrays are almost identical to Scheme vectors; the only difference is that every array in Java has a declared type T[] asserting that all elements of the array have type T. All of the primitive operations on arrays in Java preserve the declared types attached to arrays. Given an array a of type T[], the expresssion a[i] extracts the ith element of the array, which must be a value of type T. Array indexing begins with the integer value 0 as in Scheme, C, and C++. Hence, an array of length n has valid indices of 0, 1, ..., n − 1.

The assignment operator = is used to update array elements. Given an array a of type T[], the statement

a[i] = e;

updates the value of the ith element to the value of the expression e. The values of all other elements of a remain unchanged.

In Java, arrays of every type are built into the language. If you define a class or interface C, then the array type C[] is automatically supported by Java. Every array type is a subtype of type Object, but array types cannot be extended. In essence, Java arrays are identical Scheme vectors augmented by an enforced type constraint. If a Java variable has type T[] for some primitive or object type T then it can only be bound to arrays containing elements of type T. Every Java array has a int field length that contains the length of the array.

Since Java arrays are objects, a Java array value is actually a reference to an array. Hence, a variable of type T[] in Java can have the value null. A variable of type T[] can appear on the left-hand side of an assignment

     a = e

where e is any expression of type T.

Since array values are references to arrays, two variables can be bound to exactly the same array. If array variables a and b are bound to the same array object, then updating an array of a

     a[i] = e;

changes the value of b[i]. Scheme variables bound to vectors have precisely the same property.

Arrays are allocated in Java using the new statement just like other objects. The array form is

     new T[length]

where T is any type and length is an expression that evaluates to a nonnegative int. Arrays can have length 0. Each element of the array is set to the ‘zero” value of the declared type T. If T is an object type, the initial value of each element is null. Hence, the expression

     new String[1000]

allocates an array of 1000 String elements all bound to null.

Java has alternate form for constructing an array called an anonymous array. The expression

new T[] {v_{0}, ...v_{n−1}}

allocates an array of length n of type T containing the specified sequence of elements.

Java uses two mechanisms to enforce the declared type T[] of an array object. First, when an array T[] is allocated, all of the initial element values must be of type T. Second, when an array object a of type T][] is updated by an assignment

     a[i] = e

Java confirms that the the new value e for the ith element belongs to the type T. During program execution, the Java virtual machine confirms the the value e belongs to type T each time that the assignment is executed.

The array-assignment check must be performed at run-time because an array object of type S[] can be stored in a local variable or field of type T[] where S is a subtype of T. Hence, in the preceding array assignment, the declared (static) type T of e could match the declared type T[] of the array variable a, yet the assignment could fail at run-time because the value of a is an array of type S where S is a proper subtype of T.

The Java type system permits an array variable of type T[] to be bound to an array A of type S[] provided S ⊆ T. This property of the Java type system is called covariant subtyping. Note that covariant subtyping implies that the array type Object[] is a supertype for all object array types, which permits arrays to be treated polymorhically in many situations.

Recipes for Processing Arrays

Arrays do not have the same internal structure as lists: an array of length n does not contain an array of length n−1 as a component. Hence, the structural design recipes for processing lists do not directly apply to lists. But it is easy to mimic the structural decomposition of lists as follows. Given an array

A = a_{0}, ..., a_{n−1}

we define the slice

A<k,l>

where k ≥ 0 and l ≤ n as the sequence of elements

ak, ..., a_{l−1}

Since array slices are not arrays, we cannot pass them to helper functions as array values. Instead we must pass three values in the general case: the array A, and the two int bounds k and l. Fortunately, in most list processing recipes one of two bounds is always fixed (either k at 0, or l at n), so we only need to pass two values.

Assume that we want write a static method int sum that takes an argument a of type int[] and returns the sum of the elements of a. If we were summing the elements of a list instead of an array, we could use the natural recursion scheme on lists to reduce summing a compound list (a Cons) to summing its tail (the rest component of the list). (See Section 1.7.2.) We can use exactly the same scheme to sum an array provided that we use array slices instead of array values. To process slices, we must write a helper method sumHelp that takes arguments a and k of type int[] and int respectively and returns the sum of the elements in the array slice a<k,n> where n is the length of the array a. An empty slice corresponds to the case where k ≥ n. A compound slice corresponds to the case where k < n.

The following Java code implements the sum method

     class ArrayUtil {
       public static int sum(int[] a) {
         // returns a[0] + ... + a[a.length-1]
         return sumHelp(a,0);
       }
       public static int sumHelp(int[] a, int k) {
         // given 0 <= k < a.length
         // returns a[k] +  ... +  a[a.length-1]
         if (k >= a.length) then return 0;
         else return a[k] + sumHelp(a, k+1);
       }
     }

From the standpoint of computational efficiency, neither the natural recursion program or the equivalent program on array slices written above is optimal because neither one is tail-recursive. A method definition is tail-recursive if recursive calls only appear in tail-position, the last operation before returning from the method. In Scheme, the standard recipe for converting such a computation to tail recursive form involves writing a help function with an accumulating parameter and summing the elements in the opposite order (leftto-right instead of right-to-left). We can convert our array slice solution to tail-recursive form using essentially the same transformation.

The following Java code implements a tail-recursive solution using array slices:

     class ArrayUtil {
       public static int sum(int[] a) {
         return tailSumHelp(a,0,0);
       }
       public static int tailSumHelp(int[] a, int k, int accum) {
         // given 0 <= k < a.length
         // returns accum + a[k] +  ... +  a[a.length-1]
         if (k >= a.length) then return accum;
         else return tailSumHelp(a, k+1, accum+a[k]);
       }
     }

In languages that do not support the efficient translation of tail-recursive procedures to machine code, tail recursive (also called iterative) computations must be expressed in the more restrictive framework of for and while loops to produce efficient code. A tail-recursive procedure is a more general framework for expressing iterative computations than structured loops! In contrast to structured loops, tail-recursive procedures gracefully accommodate iterations with exit conditions; each procedure return clause that is not a tail-recursive call is an exit. To translate the an iterative program expressed using tail recursion to one expressed using a loop, the corresponding loop construction must have multiple exit jumps (implemented as break or go to).

Java has three familiar looping constructs: while loops, do ... while loops, C-style for loops. The first two constructs are completely standard. A while loop has syntax:

     while (test) do statement

where statement is usually a block. A block is simply a sequence of local variable declarations and statements enclosed in braces. The test expression must have boolean type. A do while loop has syntax:

     do statement while (test);

The only different between the two looping constructs is the obvious one. In a while loop the test is executed before the loop body is executed. In a do while loop the loop body is executed before the test expression is evaluated.

The Java for loop is borrowed from C. It has the form

CHAPTER 2. OBJECT-ORIENTED DATA STRUCTURES 102

for (init-expr; test; incr-expr) statement

which simply abbreviates the following code fragment containing a while loop:^{1}

init-expr;
while (test) { statement;
incr-expr; }

Let us return to our tail-recursive Java program that sums the elements of an array. Fortunately, we can translate this tail-recursive procedure directly to a simple while loop. All that we have to do is replace the recursive call a block of code that updates the procedure parameters to reflect values passed in the tail call^{2} and jumping back to the beginning of the procedure instead performing the tail call.

class ArrayUtil {
  public static int sum(int[] a) { return tailSumHelp(a,0,0); }
  public static int sumHelp(int[] a, int k, int accum) {
    // given 0 <= k < a.length, accum = accum'
    // returns accum' + a[k] +  ... +  a[a.length-1]
    while (true) {
      if (k >= a.length) return accum;
      // accum == accum' + a[0] + ... + a[k-1]
      accum = accum + a[k];
      k = k+1;
      // assignment to accum depends on k; k must be modified last
    }
}

This single exit loop can be rewritten as a conventional for loop and folded back in the sum method as follows:

class ArrayUtil {
  public static int sum(int[] a)
     returns a[0] + ... + a[a.length-1]
     int accum = 0;
     for (int k = 0; k < a.length; k++) {
      // accum == a[0] + ... + a[k-1]
      accum = accum + a[k];
    }
    return accum;
  }
}

1With one minor exception involving the use of continue in the loop body. Since our Java subset does not include continue, it is not an issue.

2Taking care to avoid interference from side effects!

The expression k++ is an abbreviation for

k = k+1;

The resulting program uses the most attractive idiom in imperative programming: the for loop. This form of processing forms the basis for the most commonly used imperative design pattern: the iterator pattern. We will discuss this pattern in detail in Section 2.1.7.

We now turn our attention to more general data representations for sequences that accommodate operations that change sequence length.

2.1.2 Lists

By convention, linked representations of sequences are called lists. This representation of sequences is so pervasive that the terms sequence and lst are often used interchangeably (conflating the abstraction and the implementation). A particularly important distinction between sequence interfaces is whether or not an interface includes operations that mutate (destructively modify) the object this. For example, if a sequence object x contains the Strings "Corky" and "Matthias", does any operation on the object x permit the contents of x to be modified, i.e., changing, adding, or subtracting elements? Operations on x that construct new sequences that incorporate the contents of x are not mutators because the object x is left unchanged.

Immutable data types are easier to define, to use, and to implement than mutable data types, but they they have two important limitations. First, they do not support some computations efficiently. Second, object mutation plays a critical role in the natural modeling of some computational problems. The functional model of computation that we studied in Chapter 1 is exclusively concerned with immutable data. The term “functional list” is synonymous with “immutable list”.

We will focus first on immutable sequences and their representations. Then we will investigate what adjustments must be made to support mutation.

2.1.3 Immutable Sequences

All of the sequence classes presented in this monograph—immutable and mutable— support the operations in the following Seq interface

     interface Seq {
       Seq empty();
       // returns Seq that is empty
       Seq cons(Object newElt);

       // returns the Seq with elts newElt, s[0], ..., s[n]
       Object first();     // returns the element s[0]
       Seq rest();         // returns an object representing s[1], ..., s[n]
       Object eltAt(int i);  // returns the element s[i]
       boolean isEmpty();    // returns n == 0
       public Object execute(SeqVisitor host);  // applies the visitor code host
     }
     interface SeqVisitor {
       Object forEmpty(Seq host);
       Object forCons(Seq host);
     }

The contracts for all of these operations stipulate that they do not modify the observable state of a sequence object.

Immutable sequence classes also support the two additional functional operations in the following interface:

     interface FinalSeq extends Seq {
       Seq updateFirst(Object val);  // returns val,this[1], ...,this[n]
       Seq updateRest(Seq r);
       // given r[0],...,r[m] returns this[0],r[0], ..., r[m]
     }

These two operations return new sequence values; they do not modify this.

There are two widely used representations for immutable sequences: linked and contiguous.

Linked Representation

In the linked representation, a sequenceis a (reference to) an object, which is either an empty node, representing the empty sequence, or a cons node with a field of type T containing the first element of the sequence and a field of type Seq containing a pointer to the first node in the rest of the sequence. This data representation, which is often called a linked list, directly corresponds to the standard inductive definition of sequences. We defined this sequence representation in Section 1.12.3 as the class List, but that definition did not support all of operations listed above. The following modification of the List composite class hierarchy from Section 1.12.3 defines a linked representation for lists of objects; it includes all of the FinalSeq operations:

     abstract class List implements FinalSeq {

/* function methods */
public Seq empty() { return EMPTY; }
public Seq cons(Object first)  { return new Cons(first, this); }
public abstract Object first();
public abstract Seq rest();
public abstract Object eltAt(int i);
abstract public boolean isEmpty();
/* mutators */
public abstract Seq updateFirst(Object f);
public abstract Seq updateRest(Seq r);
abstract String toStringHelp();
// List -> String without any parentheses and leading blanks
static final Empty EMPTY = new Empty();
private static class Empty extends List {
/* constructor */
  private Empty() {}
/* methods */
public Object first() {
  throw new IllegalArgumentException("first() applied to empty list");
}
public Seq rest() {
  throw new IllegalArgumentException("rest() applied to empty list");
}
public int isEmpty() { return true; }
public Seq updateFirst(Object o) {
  throw new IllegalArgumentException("updateFirst() applied to empty list");
}
public Seq updateRest(Seq s) {
  throw new IllegalArgumentException("updateFirst() applied to empty list");
}
public Object eltAt(int i) {
  throw new IllegalArgumentException("out-of-bounds index in List.eltAt");
}
public Object execute(SeqVisitor v) { return v.forEmpty(this); }
public String toString() { return "()"; }

public String toStringHelp() { return ""; }
}
class Cons extends List {
/* fields */
private final Object first;
private final List rest;
/* constructor */
Cons(Object f, List r) {
  first = f;
  rest = r;
}
/* methods */
public Object first() { return first; }
public Seq rest() { return rest; }
public int isEmpty() { return false; }
public Object eltAt(int i) {
  if (0 == i) return first;
  else return rest.eltAt(i-1);
}
public Object execute(SeqVisitor v) { return v.forCons(this); }
public Seq updateFirst(Object o) { return rest.cons(o); }
public Seq updateRest(Seq r) { return r.cons(first); }
public String toString() { return "(" + first + rest.toStringHelp() + ")"; }
String toStringHelp() { return " " + first + rest.toStringHelp(); }
}
}

The definition of the List class contains nested class definitions for the classes Empty and Cons. The static attribute identifies these classes as nested classes rather than inner classes. Nested classes are identical to conventional “toplevel” classes except for two minor differences.

• First, nest classes have qualified names (of the form containing-class-name.nested-class-name).^{3} The full names for the classes Empty and Cons are List.Empty and List.Cons, respectively. Within the body of the List class, the unqualified names Empty and Cons are synonymous with the qualified names.

3Since classes may be nested to any depth, multiply nested classes have multiple qualifiers, one for each level of nesting.

• Second, nested classes can be declared as private, making their names invisible outside the body of the containing class. Hence, the classe names List.Empty and List.Cons are not defined outside of the body of the List class. If we removed the private attribute for the Empty class above, then the classes

In contrast to instances of inner class, instances of nested classes do not have enclosing instances. Section 2.1.6 discusses nested and inner classes in more detail.

The for.. methods in the SeqVisitor interface all take a host argument of type Seq because the implementation is not constrained to use the composite pattern to represent immutable sequences. The following visitor class implements sequence concatenation:

     class Append implements SeqVisitor {
       // returns sequence host || that
       /* fields */
       Seq that;
       /* constructor */
       Append(Seq t) { that = t; }
       /* methods */
       public Object forEmpty(Seq host) { return that; }
       public Object forCons(Seq host) {
         return host.updateRest((Seq) host.rest().execute(this));
       }
       public static void main(String[] args) {
         Seq s1 = List.EMPTY.cons("B").cons("A");
         Seq s2 = List.EMPTY.cons("C");
         System.out.println("s1 = " + s1);
         System.out.println("s2 = " + s2);
         System.out.println("s1 * s2 = " + s1.execute(new Append(s2)));
       }
     }

The following figure shows a picture of linked list of integers.

In the figure, the nodes with two fields are Cons instances, and the crossedbox is an Empty instance. References (pointers) are represented by the heavy arrows. The reference fields in the cells are in fact memory addresses. In Java, these addresses are always interpreted as references to objects. Java only supports operations on references that are consistent with this abstraction, e.g. you cannot perform arithmetic on a reference. In lower-level languages like C and C++, references (pointers) can be manipulated as ordinary integers.

Finger Exercise 2.1.3.1 Write a SeqVisitor class to reverse a Seq. Test your code using DrJava. Hint: introduce a helper visitor with an extra parameter to accumulate the result.

Contiguous Representation

In the contiguous representation, a sequence is represented by a reference to an immutable array of fields of type T. An immutable array is an array that, once initialized, is never modified. Java doesn’t directly support immutable arrays, but a Java program can enforce immutability by defining a wrapper class for arrays (akin to the Integer class for wrapping ints in objects) with a single private field holding the embedded array and a collection of public methods that do not mutate this field. A lighter weight but less robust protocol for supporting immutable arrays is to use comments to indicate which arrays are immutable and to follow the discipline of never modifying arrays documented as immutable. In either case, a new array object generally must be created whenever an element of the represented sequence is changed, added, or removed. Creating an array object is a costly operation proportional in time and space to the number of elements in the new array.

In the linked representation of sequences, every operation in the collection listed above except eltAt can be performed in constant time. On the other hand, the eltAt operation takes time proportional to the length of the sequence in both the worst case and the typical case. The List implementation of sequences given in chapter 1 has this property.

The performance trade-offs embodied in the immutable array implementation are very different. In this implementation, the operations empty, first, length, eltAt, can be performed in constant time. (In the Java array implementation, length is stored in a separate “field” in the block of storage holding the array.) With the exception of the rest operation, the remaining operations all take time proportional to the length of the sequence. The running time of the rest operation depends on an interesting implementation detail. If immutable arrays are implemented as instances of a “wrapper” class, then the rest operation can be performed in constant time at the cost of making an extra field reference in the implementation of eltAt. A wrapper object can store an integer offset that is added to the index passed as an argument to eltAt. In this scheme, the rest operation constructs a new wrapper object containing a pointer to the same array object as this but an increased offset (by one). If immutable arrays are implemented directly by Java arrays, then rest operation must construct a completely new array one element shorter than the original.

Finger Exercise 2.1.3.2 Construct two implementations of an ImmutArray wrapper class that represents sequences as arrays. Do not include an offset field in the first implemenation. Include a offset field in the second implementation. Test the Append and Reverse visitors written in the context of the linked representation above and your contiguous implementations. Conduct some experiments to measure the performance impact of including the offset pointer. For each implementation, can you devise a test program that favors it?

In practice, array representations of immutable sequences are generally not used in computations that make extensive use of the cons and empty operations to construct new sequences. The repeated copying of arrays required to support these operations is very inefficient (proportional to the square of the length of the constructed array!)

Finger Exercise 2.1.3.3 Let n be the length of a sequence host represented either as a linked list or an array. Prove that the computation

     host.execute(new Append(host.empty()))

runs in time O(n) in the linked representation, O(n^{2}) in the contiguous representation (with or without an offset field).

The usual way to avoid this source of inefficiency is to include an operation in the immutable array class that constructs an array representation for a sequence given either a corresponding linked representation or mutable array representation. The Java Foundation classes include both the immutable string (sequence of char) class String and the mutable string class StringBuffer for this reason.

The array implementation of immutable sequences is a good choice when new sequences are generally built from scratch rather than constructed by applied operations to existing sequences. For this reason, many computations involving immutable sequences of characters (strings) rely on the array representation. The Java String class implements the array representation for immutable sequences of characters (the primitive type char in Java). Note that Java includes an operation for converting mutable strings (represented by the Java class StringBuffer) to immutable strings (represented by the Java class String). Strings can be incrementally constructed from characters far more efficiently using the StringBuffer class that they can be using the

String class.

2.1.4 Mutable Sequences

A sequence implementation is mutable if it includes operations that modify the value of this. A class representing mutable sequences implements a subset of the following operations:

     interface SeqObject extends Seq {
       void setFirst(T f);          // this = this.updateFirst(f)
       void setRest(Seq r);         // this = this.updateRest(r)
       void set(Seq v);             // this = v
       void setEltAt(int i, T val); // changes s[i] in this to val
       void insert(Object o);       // inserts o in front of s[0] in this
       void remove();               // removes s[0] from this
     }

As with immutable sequences, there are two basic implementation schemes for mutable sequences: linked and contiguous. Mutation complicates the implementation of linked representations, which we examine in detail below.

Singly-linked Mutable List Representation

The various linked mutable list representations that we will study are all derived from the standard linked representations for immutable sequences. A particularly simple approach to sequence mutation is to represent a mutable sequence as a variable x of immutable list type and implement mutation by assigning a new value to the variable, e.g.

       List x = empty();
       .
       .

       .
       x = cons(0,x);

But this approach fails to represent mutable lists as objects and to encapulate list mutation as an ordinary method. This representation cannot implement the insert method given above. In the list container representation, the insert operation modifies the program variable representing the mutable sequence, but variables are not objects! When we pass an immutable lists represented by an assignable variable as a method arguments, we can only pass the immutable list value to which the variable is bound.

Finger Exercise 2.1.4.1 Try to write an insert method for mutable sequences represented by variables bound to immutable sequences. What goes wrong?

2.1.5 List Containers

A better approach is to define a container class with a single field that holds an immutable sequence. Then we can update the mutable sequence by modifying the contents of the field in the container object. For example, suppose we have a class List that defines a list representation for immutable sequences. The following container class works for any implementation of the Seq interface:

     class ListBox implements SeqObject {
       private static Seq prototype = new ...;
       // any instance of the class implementing Seq
       /* fields */
       private Seq value;  // contents of ListBox: s[0],s[1],...,s[n-1]
       /* constructors */
       ListBox() { value = prototype.empty(); }
       ListBox(Seq v) { value = v; }
       /* visible accessor methods */
       public Seq empty() { return new ListBox(); }
       public Seq cons(Object newElt) { return new ListBox(value.cons(newElt)); }
       public Object first() { return value.first(); }
       public Seq rest() { return value.rest(); }
       public Object eltAt(int i) { return value.eltAt(i); }
       public boolean isEmpty() { return value.isEmpty(); }
       /* visible mutator methods */

public void setFirst(Object o) { value = value.updateFirst(o); }
public void setRest(Seq r) { value = value.updateRest(r); }
public void set(Seq v) { value = v; }  // set contents of box to v;
public void setEltAt(int i, final Object val) {  // changes s[i] to val
  return execute(new UpdateEltAt(i), val);
}
public void insert(Object o) { value = value.cons(o); }
// changes contents of this from s[0],...,s[n] to o,s[0],...,s[n]
public void remove() { value = value.rest; }  // removes s[0] from the sequence
public Object execute(SeqVisitor v) { return value.execute(v); }
// apply visitor v to value and return result; value is UNCHANGED
/* inner classes */
private class UpdateEltAt implements SeqVisitor {
  /* fields */
  int index;          // index of element to be updated
  Object eltValue;    // new value for updated element
  /* constructor */
  UpdateEltAt(int i, Object e) { index = i; eltValue = e; }
  /* visit methods */
  Object forEmpty(Seq host) { throw
    new IllegalArgumentException("out-of-bounds index in UpdateEltAt");
  }
  Object forCons(Seq host) {
    if (index == 0) return new Cons(val, host.rest());
    else return host.rest().execute(new UpdateEltAt(i-1));
}
}

The variable holding the Seq instance is now wrapped inside an instance of a class (ListBox above) implementing the SeqObject interface. A method that accepts a SeqObject as an argument can modify it.

In the preceding example, the use of the inner class UpdateEltAt warrants careful study. This inner class is a private member of the ListBox class and hence inaccessible outside of the class body. Since the inner class referenced only within the setEltAt method, we could have placed the definition of UpdateEltAt as a declaration inside the body of this method! In this case, the UpdateEltAt class would have been visible only within the method setEltAt But such an embedded inner class definition can be hard to read, so we elected to make a private member of the ListBox instead.

We have already seen a special case of inner classes, namely anonymous classes that appear inside dynamic methods. The only difference between an inner class and a conventional class is that every instance of an inner class has an enclosing instance of the class C that textually contains its definition. The free variables in the inner class refer to this class. In addition, the notation C.this refers to the “entire” enclosing instance. Inner classes impose exactly the same restrictions on references to local variables of the enclosing instance as anonymous classes do: any such local variable must be declared final. Inner classes are discussed in more detail in Section 2.1.6.

In ListBox class, the methods insert, remove, and set modify the receiver of the method invocation, so there is no need for them to return values. Consequently they have return type void. Mutator methods typically have the return type void because they embody commands that modify objects rather than functions that compute new values based on the value of this.

The “lists as containers” representation of mutable lists is a very simple example of the state pattern. In the state pattern, a mutable object contains a field of union type (denoted by an abstract class or interface) representing the state of the object. The object can easily change “shape” by updating the field to contain an instance of a different class in the union. In the ListBox class, an empty list object can mutate to a non-empty list object (or viceversa) by modifying the contents of the value field containing a Seq, which is a union type.

Since the SeqObject interface extends the Seq interface, it inherits the visitor interface from the immutable Seq interface. As a result, no visitor class implementing the SeqVisitor interface can mutate a SeqObject! In particular, to mutate fields of a ListBox object, we must use explicit assignment. Given a ListBox l and SeqVisitor v that returns a ListBox, the assignment

     l = (ListBox) l.execute(v);

updates l to the new value returned by the visitor operation.

The efficient operations on ListBox objects are precisely the efficient operations on the underlying functional List class, namely, adding and removing elements at the front of the list. Mutable lists in which elements can only added or removed at the front are called stacks or LIFO (”last in, first out”) lists. Representing mutable lists as containers holding immutable list values is well-suited to this form of usage. The operations first, insert, and pop precisely match the usual operations push, top, and pop on stacks.

The “container” representation for mutable lists is simple and easy-to-use but it is poorly suited to many applications because it fails to support certain list operation efficiently. This representation forces list nodes (Cons objects) to be recopied whenever the list is changed. To modify the list element with index i or insert and element in front of the list element with index i, a computation must construct a new List, copying the elements from the old List with indices less than i. We can avoid this recopying process and avoid the potentially confusing distinction between immutable list values and mutable list objects by using a mutable variant of functional lists developed by Dung Nguyen and Steve Wong.

2.1.6 Quasi-Functional Lists

From the perspective of the public interfaces, quasi-functional lists differ from lists as containers in two respects. First, quasi-functional lists require the list arguments for setRest and set to be mutable list objects rather than immutable list values. Second, quasi-functional lists support visitor operations that mutate list structure in addition to the “purely functional” visitor operations corresponding to the SeqVisitor interface. To capture these differences in the Java type system, we introduce two new interfaces: a new mutable sequence interface called MutSeq:and a mutable visitor interface called

MutSeqVisitor:
     interface MutSeq extends Seq {
       void setFirst(Object f);         // changes this.first = f
       void setRest(MutSeq r);          // changes this.rest = r
       void set(MutSeq m);              // changes this = m
       void setEltAt(int i, Object val);     // changes this[i] = val
       void insert(Object o);           // changes this.first,this.rest = o,this
       void remove();                   // changes this = this.rest
       Object execute(MutSeqVisitor m); // applies visitor operation m to this
     }
     interface MutSeqVisitor {
       Object forEmpty(MutSeq host);
       Object forCons(MutSeq host);
     }

The MutSeq interface stipulates that the arguments to the operations setRest and set must be list objects, (objects of type MutSeq) rather than the list values (objects of type Seq) given in the SeqObject interface. The MutSeq interface also introduces an execute operation to support visitor operations (objects of type MutSeqVisitor) that mutate list structure. The MutSeqVisitor interface differs from the SeqVisitor interface in one key respect: the host object must be a mutable list (object of type MutSeq) rather than a list value (object of type Seq) enabling a visitor to destructively modify the host. This MutSeqVisitor interface is not applicable to lists as containers because the component rest fields embedded in the immutable list value are not mutable! The pivotal difference between the QuasiList and ListBox classes is the type of the rest field. A MutSeqVisitor can destructively modify both the first and rest fields of the host by using MutSeq mutator methods setFirst and setRest.

class QuasiList implements MutSeq {
/* fields */
public static final Empty EMPTY = new Empty();
private List value;
/* constructor */
QuasiList() { value = new Empty(); }
private QuasiList(List v) { value = v; }
/* visible methods */
Seq empty() { return new QuasiList(); }
Seq cons(Object newElt) { return new QuasiList(value.cons(newElt)); }
Object first() { return value.first(); }
Seq rest() { return value.rest(); }
// rest returns a MutSeq (QuasiList) but the Seq interface mandates
//   the weaker type!
Object eltAt(int i) { return value.eltAt(i); }
boolean isEmpty() { return value.isEmpty(); }
void insert(Object o) { value = new Cons(o, new QuasiList(value)); }
public String toString() { return value.toString(); }
void setFirst(Seq v) { value = value.updateFirst(v); }
void setRest(MutSeq m) { value = value.updateRest(m); }
void set(MutSeq m) { value = m.value; }
void setEltAt(int i, final Object val) {
  /* inner class */
  class UpdateEltAt implements MutSeqVisitor {
    /* fields */
    final int index;       // index of element to be updated

  /* constructor */
  UpdateEltAt(int i) { index = i; }
  /* visit methods */
  Object forEmpty(MutSeq host) { throw
    new IllegalArgumentException("out-of-bounds index in UpdateEltAt");
  }
  Object forCons(MutSeq host) {
    if (index == 0) return host.setFirst(val);
    else return host.rest().execute(new UpdateEltAt(index-1));
  }
}
execute(new UpdateEltAt(i));
}
void remove() { value = value.rest(); }
Object execute(SeqVisitor v) { return value.execute(v); }
// apply visitor v to value and return result; value is UNCHANGED
Object execute(MutSeqVisitor v) { return value.execute(v,this); }
// apply visitor v to value and return result; value may be CHANGED
/* inner classes */
private interface List {
abstract String toStringHelp();
// List -> String without any parentheses and leading blanks
}
private class Empty extends List {
/* constructor */
private Empty() {}
/* methods */
Object first() {
throw new IllegalArgumentException("first() applied to empty list");
}
MutSeq rest() {
throw new IllegalArgumentException("rest() applied to empty list");
}
Object eltAt(int i) {
  throw new IllegalArgumentException("out-of-bounds index in List.eltAt");

}
Object execute(SeqVisitor v) { return v.forEmpty(this); }
Object execute(MutSeqVisitor v) { return v.forEmpty(QuasiList.this); }
public String toString() { return "()"; }
public String toStringHelp() { return ""; }
}
private class Cons extends List {
/* fields */
private final Object first;
private final MutSeq rest;
/* constructor */
Cons(Object f, List r) {
  first = f;
  rest = r;
}
/* functional methods */
Object first() { return first; }
Seq rest() { return rest; }
/* MutSeq is the correct output type but Java does not support it */
Object eltAt(int i) {
  if (0 == i) return first;
  else return rest.eltAt(i-1);
}
/* mutator methods */
void setFirst(Object o) { first = o; }
Object execute(SeqVisitor v) { v.forCons(this); }
Object execute(MutSeqVisitor v) { v.forCons(QuasiList.this); }
public String toString() {
  return "(" + first + rest.toStringHelp() + ")"; }
String toStringHelp() { return " " + first + rest.toStringHelp(); }
}
}

The QuasiList implementation given above uses the state pattern to represent each tail (rest component) of the list. Each tail is a quasilist object that can mutate between two forms: empty and non-empty. Since each tail is a mutable object supporting the state pattern, a MutSeqVisitor can modify the state of any tail in the process of traversing a list.

The QuasiList code uses inner classes to hide the classes implementing the state of a QuasiList and to eliminate passing the QuasiList host as an extra parameter to the execute(MutSeqVisitor v) methods in the Cons and Emtpy subclasses of List. If the List class is moved outside of QuasiList, the QuasiList object containing a given List object is not accessible to the List object.^{4}

Nested Classes vs. Inner Classes

A nested class is a class whose definition appears inside the definition of another class, as if it were a member of the other class. For example, if a program contains

class A {
  class B {
    // fields, methods of class B...
  }
  // fields, methods of class A...
}

then class B is a nested class of class A. Code outside of the methods of class A can refer to class B by calling it A.B, using the same dot notation as for field and method references. Within the methods of class A class B can be used without qualifying the name. B could be hidden from code outside of class A by declaring it private, just as with fields and methods.

A nested class like B is known as an inner class. An inner class has access to the fields of an instance of its enclosing class. For this reason, an instance of B can only be created in association with an instance of A, using the expression

     instanceA.new A.B(...)

outside of A’s methods, where instanceA is an instance of A, and

     new B(...)

inside of A’s methods. The new instance of B also knows about the enclosing instance of A and can refer to it using the expression

4Another alternative is add an owner field to the abstract class List that refers to the containing QuasiList object but this approach complicates the form of the constructors for Cons and Empty, which must take an additional argument to initialize the owner field.

     A.this

We can think of an instance of an inner class as having two this references, one for itself and one for its enclosing instance. An inner class may be nested within another inner class, so an inner class can even have multiple levels of this references. Nesting inner classes more deeply than one level is quite uncommon, however, and should usually be avoided.

A nested class can be declared static, in which case it has reduced access to its enclosing class. For example, if B were declared static above, it could no longer access the instance variables of A, and there would be no associated instance A.this. Static nested classes are known as nested top-level classes, because they are exactly like classes declared outside any other class, except for the way they are named. Instances of a static nested class are created using regular new, as in

     new A.B(...)

We’ll see uses for both static nested classes and inner classes when we present the full implementation of imperative lists.

Mutable Visitors

Quasi-functional lists are more flexible than lists as containers because the MutSeq interface includes support for visitor operations that mutate the structure of a list. The following visitor implements the operation of destructively inserting an element at the rear of the sequence:

     class InsertRear implements MutSeqVisitor {
       /* given the embedded Object elt and a host with elements s[0], ... s[n],
          host.execute(this) destructively updates host so that host =
          s[0],..., s[n],elt */
       /* field */
       private Object elt;
       /* constructor */
       InsertRear(Object e) { elt = e; }
       Object forEmpty(MutSeq host) {
         host.insert(elt);
         return null;  /* dummy return value; this operation has return type void!
       }

       Object forCons(MutSeq host) {
         ((MutSeq) host.rest()).execute(this);
         return null;  /* dummy return value; the return ``type'' is void!
       }
     }
     class MutAppend implements MutSeqVisitor {
       /* given the embedded MutSeq tail with elements t[0], ..., t[n] and a host
          with elements s[0], ... s[n], host.execute(this) destructively
          updates host so that host = s[0],..., s[n],tail[0],...tail[m] */
       /* field */
       private MutSeq tail;
       /* constructor */
       MutAppend(Object t) { tail = t; }
       Object forEmpty(MutSeq host) {
         host.set(tail);
         return host;  /* dummy return value; this operation has return type void!
       }
       Object forCons(MutSeq host) {
         return ((MutSeq) host.rest()).execute(this);
       }
     }

The primary disadvantage of quasi-functional lists is that sharing list tails between two list objects can produce unexpected results when list objects are mutated. Mutating a shared list tail changes all of the list objects that share that tail! In the programming literature, the sharing of mutable data objects is often called “aliasing”.

Finger Exercise 2.1.6 To be provided: an example involving aliasing.

2.1.7 Extended Mutable Lists

Both of the preceding representations of mutable sequences—lists as containers and quasi-functional lists—are inefficient at inserting elements at the rear of a sequence. In each of these representations, the code for the operation must scan the entire sequence to reach the end. The container representation is particularly inefficient in this regard because the entire sequence must be reconstructed starting with a singleton List containing the new element.

Mutable sequence implementations that efficiently support adding elements at rear of the list and removing them from the front are important because this access protocol, called a queue or a FIFO (first in, first out) list, frequently arises in practice. Procedural formulations of linked lists discussed in traditional textbooks on data structures provide constant-time access to the end of the list by maintaining a “pointer” to the last node of the list. This strategy is conceptually simple but prone to coding errors because the empty list state requires special treatment. Moreover, the traditional procedural approach to representing lists exposes the concrete data structures (nodes and links) used in the implementation. We can exploit the same strategy in an object-oriented representation of lists that hides the concrete data structures in private object fields—provided that we deal carefully with the potentially troublesome “boundary” cases in the definition of list operations that involve the empty list.

Formulating Traditional Linked Lists as Objects

The quasi-list representation of mutable sequences includes an extra level of object nesting in the representation of list tails beyond what is present in the conventional “singly-linked list” representations that are widely used in procedural programming. A major disadvantage of this data representation is the extra memory required to hold the extra object allocated for each node. The basic singly-linked list representation avoids this extra overhead; it relies on the exactly same data representation as the “lists as containers” representation given in Section 2.1.5 with one critical modification: the first and rest fields of Cons objects are mutable. The following Java code implements the MutSeq interface using conventional singly-linked lists rather than quasi-lists.

     class MutList implements MutSeq {
       /* fields */
       static final Empty EMPTY = new Empty();  // singleton empty list
       List value;
       /* constructors */
       MutList() { value = EMPTY; }
       private MutList(List v) { value = v; }
       /* visible methods */
       Seq empty() { return new MutList(); }
       Seq cons(Object newElt) { return new MutList(value.cons(newElt)); }

Object first() { return value.first(); }
// returns the element s[0]
Object rest() { return MutList(value.rest()); }
// returns a MutList containing elements s[1],...,s[n-1]
Object eltAt(int i) { return value.eltAt(i); }
// returns the element s[i]
boolean isEmpty() { return value.isEmpty(); }
//  yields the number of elements in the sequence
void insert(Object o) { value = value.cons(o); }
// returns new MutList with elts o, s[0], s[1], ..., s[n]
void setFirst(Object o) { value = value.setFirst(o); }
void setEltAt(int i, final Object val) {  // changes s[i] to val
  class UpdateEltAt implements MutSeqVisitor {
    /* fields */
    int index;       // index of element to be updated
    /* constructor */
    UpdateEltAt(int i) { index = i; }
    /* visit methods */
    Object forEmpty(MutSeq host) { throw
      new IllegalArgumentException("out-of-bounds index in UpdateEltAt");
    }
    Object forCons(MutSeq host) {
      if (index == 0) {
host.setFirst(val);
return null;
      }
      else host.rest().execute(new UpdateEltAt(i-1));
  }
  value = execute(new UpdateEltAt(i));
}
void remove() { value = value.rest; }
// removes s[0] from the sequence

Object execute(SeqVisitor v) { return value.execute(v); }
// apply visitor v to value and return result; value is UNCHANGED
Object execute(MutSeqVisitor v) {
// apply visitor v to this; value may CHANGE
if (value == EMPTY) then return v.forEmpty(this)
else return v.forCons(this);
}
private static abstract class List implements Seq {
abstract void setFirst(Object o};
abstract List cons(Object o};
abstract Object first();
abstract Seq rest();
abstract boolean isEmpty();
}
private class Empty extends List {
public void setFirst(Object o) {
  throw new IllegalArgumentException(setFirst() applied to empty list");
}
public List cons(Object o} { new Cons(o,MutList.this); }
public Object first() {
  throw new IllegalArgumentException("first() applied to empty list");
}
public Seq rest(); ???

  throw new IllegalArgumentException("rest() applied to empty list");
}
public int length() { return 0; }
}
private static class Cons extends List { ... }
}

Note that we have defined the List classes as inner classes to hide them from clients of MutList. This feature distinguishes our representation of basic linked lists from the traditional representation used in procedural languages. By embedding the Node class hierarchy inside the definition of the MutList class, we have completely hidden the fact that we are using a conventional linked list representation! To client code, MutList is semantically indistinguishable

from QuasiList!

What have we gained? First, the MutList class is a more efficient implementation of the MutSeq interface corresponding to quasi-lists because it CHAPTER 2. OBJECT-ORIENTED DATA STRUCTURES 124

allocates only one object for each list node instead of two. Second, we can easily expand the MutList class to include constant-time methods for adding and element to the end of a list and appending to lists. The extended class maintains a reference to the subsequnce containing last element The following ExtMutList class provides these new methods. Since the implementation relies on maintaining references to both the first and last nodes of the list (value and last, the changes to MutList required to create ExtMutList are non-trivial.

class ExtMutList implements ExtMutSeq {
/* fields */
static final Empty EMPTY = new Empty();  // singleton empty list
List value;
/* constructors */
MutList() { value = EMPTY; }
private MutList(List v) { value = v; }
/* visible methods */
Seq empty() { return new MutList(); }
Seq cons(Object newElt) { return new MutList(value.cons(newElt)); }
Object first() { return value.first(); }
// returns the element s[0]
Object rest() { return MutList(value.rest()); }
// returns a MutList containing elements s[1],...,s[n-1] where n=length(this)
Object eltAt(int i) { return value.eltAt(i); }
// returns the element s[i]
int length() { return value.length(); }
//  yields the number of elements in the sequence
void insert(Object o) { value = value.cons(o); }
// returns new MutList with elts o, s[0], s[1], ..., s[n]
void setFirst(Object o) { value = value.setFirst(o); }
void setEltAt(int i, final Object val) {  // changes s[i] to val
  class UpdateEltAt implements MutSeqVisitor {

    /* fields */
    int index;       // index of element to be updated
    /* constructor */
    UpdateEltAt(int i) { index = i; }
    /* visit methods */
    Object forEmpty(MutSeq host) { throw
      new IllegalArgumentException("out-of-bounds index in UpdateEltAt");
    }
    Object forCons(MutSeq host) {
      if (index == 0) {
host.setFirst(val);
return null;
      }
      else host.rest().execute(new UpdateEltAt(i-1));
  }
  value = execute(new UpdateEltAt(i));
}
void remove() { value = value.rest; }
// removes s[0] from the sequence
Object execute(SeqVisitor v) { return value.execute(v); }
// apply visitor v to value and return result; value is UNCHANGED
Object execute(MutSeqVisitor v) {
  // apply visitor v to this; value may CHANGE
  if (value == EMPTY) then return v.forEmpty(this)
  else return v.forCons(this);
}
private static abstract class List implements Seq {
  abstract void setFirst(Object o};
  abstract List cons(Object o};
  abstract Object first();
  abstract Seq rest();
  abstract int length();
}
private class Empty extends List {
  public void setFirst(Object o) {
    throw new IllegalArgumentException(setFirst() applied to empty list");
  }

  public List cons(Object o} { new Cons(o,MutList.this); }
  public Object first() {
    throw new IllegalArgumentException("first() applied to empty list");
  }
  public Seq rest();

    throw new IllegalArgumentException("rest() applied to empty list");
  }
  public int length() { return 0; }
}
private static class Cons extends List { ... }
}
interface ExtMutSeq extends MutSeq {
void insertRear(Object e);
void mutAppend(ExtMutSeq t);
}
class ExtMutList extends MutList implements ExtMutSeq {
/* fields */
Node value;
Node last;
/* relies on default constructor that calls super() */
/* visible methods */
Object first() { return value.first(); }
Seq rest() { return value.rest(); }
Object eltAt(int i) { return value.eltAt(i); }
// returns the element s[i]
int length() { return value.length(); }
//  yields the number of elements in the sequence
void insert(Object o) { value = value.cons(o); } /* last ?? */
// returns new ExtMutList with elts o, s[0], s[1], ..., s[n]
void setFirst(Object o) { value = value.setFirst(o); }

void setEltAt(int i, final Object val) {  // changes s[i] to val
class UpdateEltAt implements MutSeqVisitor {
  /* fields */
  int index;       // index of element to be updated
  /* constructor */
  UpdateEltAt(int i) { index = i; }
  /* visit methods */
  Object forEmpty(MutSeq host) { throw
    new IllegalArgumentException("out-of-bounds index in UpdateEltAt");
  }
  Object forCons(MutSeq host) {
    if (index == 0) {
host.setFirst(val);
return null;
    }
    else host.rest().execute(new UpdateEltAt(i-1));
}
value = execute(new UpdateEltAt(i));
}
void remove() { value = value.rest; }
// removes s[0] from the sequence
Object execute(SeqVisitor v) { return value.execute(v); }
// apply visitor v to value and return result; value is UNCHANGED
Object execute(MutSeqVisitor v) {
// apply visitor v to this; value may CHANGE
if (value == EMPTY) then return v.forEmpty(this)
else return v.forCons(this);
}
private static abstract class Node {
abstract void setFirst(Object o};
abstract void cons(Object o};
abstract Object first();
abstract Node next();
abstract Object eltAt(int i);
abstract int length();

       }
       private static class Empty extends Node { ... }
       private static class Cons extends Node { ... }
     }

Spaghetti References (akin to spaghetti code)

We strongly endorse the slogan ”pointers are the root of all evil.” In the 1970s, structured programming was developed, which placed a discipline on the control flow of programs. Up until that point, languages (and programmers) had been very liberal in their use of the notorious goto statement or unconditional branch. Undisciplined use of goto statements to programs whose control flow is almost impossible to trace by hand (spaghetti code), and which are hence almost impossible to debug. In his Turing Award Lecture in the early 1970s, C.A.R. Hoare asserted that multiple references to the same data in a program can introduce a similar form of complexity if that data can be mutated through those references. With multiple references to mutable data in a program, we can stumble into serious trouble when one part of our program modifies an object when another part is unaware of or unprepared for the modification.

Even in programs written and maintained by a single programmer, the unrestricted use of pointers is foolhardy and invites disaster. A programmer almost certainly does not remember the precise reasoning used involved in writing every section of a significant application. As a result, mistakes are inevitable. Just as structured programming placed constraints on the control flow in programs, we will place constraints on when and where mutable data can be changed. This is why we hide mutable nodes inside a list container.

The Iterator Pattern

Before we present a complete implementation of singly-linked imperative lists, we describe a new pattern which allows us to process lists similarly to arrays. The pattern is called the Iterator pattern, and consists of two interfaces, an abstract list, with methods for building and modifying lists

     interface ItSeq extends {
       void insertFront(Object o);
       void insertRear(Object o);
       boolean isEmpty();
       void remFront();          // remove the first element
     }

and an abstract iterator, with methods for traversing a list and examining its contents

     interface IIterator {
       void front();             // move to first element
       void next();
       boolean atEnd();          // test whether past the last element
       Object currentItem();     // contents of current element
     }

The iterator’s atEnd method returns true if and only if the iterator has moved past the last element of the list. When the iterator is in this state, the currentItem method will throw an exception.

With such a list and such an iterator we could easily implement a queue, since we can remove from the front and add at the back.

It would be nice if the list were more flexible however. For example, we may wish to sort a list. We can already do this in a functional style, by building a new list while using insertion sort, but since this is a mutable list we should ideally be able to sort the list without copying the nodes, changing it from an unsorted list into a sorted one. We can implement such a sort if we add two more methods to the Iterator I implementation:

       void insertBefore(Object o); // add new element before current
       void remove();               // remove current

A given list may have more that one iterator active on it at one time, so the remove and insertBefore methods must be used with some care.

The definition of an iterator class implementing the Iterator I involves two subtle issues: First, for an iterator to remove the current element, it must have a reference to the element immediately before the current element.

Second, we can treat the empty list like any other list if we include a dummy node which is always at the head of the list. This dummy node simplifes the implementation of element removal when that element is the last element of the list. When the list is empty, the last field refers to dummy node.

2.1.8 An Implementation

All of our implementation sketches of the last few sections now culminate in the following real list implementation. Notice that we make use of both static nested and inner classes in an effort to hide imperative details.

// (Singly-Linked) Mutable Lists
// The IList interface includes a method newIterator that creates
// a new iterator for a list.  The List Class implementing IList
// hides the Iterator class implementing IList.  As a result,
// invoking newIterator is the only way to create new iterators.
interface IList {
IList newList();
int length();
void insertFront(Object o);
void insertRear(Object o);
void remFront();
boolean isEmpty();
IIterator newIterator();
}
interface IReadIterator {
// interface for processing both mutable and immutable lists
void first();
void next();
boolean atEnd();
Object currentItem();
}
interface IIterator extends IReadIterator {
/* Destructive operations */
void insert(Object o);
// inserts before current item; when atEnd(), does insertRear(o)
void remove();
}
// Exception classes for Lists and Iterators

class ListException extends RuntimeException {
ListException(String s) { super(s); }
}
class IteratorException extends RuntimeException {
IteratorException(String s) { super(s); }
}
class List implements IList {
// ** fields **
private Node head = new Node();  // allocate header node
private Node last = head;
private int length = 0;
// ** constructors **
// relying on default constructor
// ** toString() **
public String toString() {
  IIterator i = new Iterator();
  String result = "(";
  for (i.first() ; ! i.atEnd(); i.next())
    result = result + " " + i.currentItem();
  return result + " )";
}
// ** methods of IList **
public IList newList() { return new List(); }
public int length() { return length; }
public void insertFront(Object o) {
  Node oldSucc = head.succ;
  Node newNode = new Node(o,oldSucc);
  head.succ = newNode;
  if (last == head) last = newNode;
  length++;
}
public void insertRear(Object o)  {

Node newNode = new Node(o,null);
last.succ = newNode;
last = newNode;
length++;
}
public void remFront() {
if (isEmpty()) throw new
  ListException("remFront() applied to EmptyList");
else {
  Node newSucc = head.succ.succ;
  head.succ = newSucc;
  if (newSucc == null) last = head;
  length--;
}
}
public boolean isEmpty() { return head == last; }
public IIterator newIterator() {
return new Iterator();
}
// ** hidden classes Node and Iterator **/
private static class Node {
/* fields */
Object item;
Node succ;
/* constructors */
Node(Object i, Node s) {
  item = i;
  succ = s;
}
Node() {  // allocate header
  item = null;
  succ = null;
}
// fields are accessed directly by code in List class

}
private class Iterator implements IIterator {
// NOTE: Iterator points to predecessor of current item.
// Hence, current item is pred.succ
/* fields */
Node pred;
/* Constructors */
Iterator() {
  pred = head;
}
/* methods in IIterator interface */
public void first() {
  // reposition cursor to refer to first item (if one exists)
  pred = head;
}
public void next() {
  // advance cursor
  if (atEnd()) throw new
      IteratorException("No next element in Iteration");
  pred = pred.succ;
}
public Object currentItem() {
  // returns current item
  if (atEnd()) throw new
    IteratorException("No current element in " + List.this);
  return pred.succ.item;
}
public boolean atEnd() { return pred == last; }
// returns true iff cursor points to imaginary element beyond last
public void insert(Object o) {
  // pre: current is either a list element or an imaginary
  //      element just beyond the last element
  // post: Node containing o is inserted before current item,

           //       current is unchanged (pred is changed, last may be)
           Node oldSucc = pred.succ;
           Node newNode = new Node(o, oldSucc); // allocate new node
           pred.succ = newNode;                 // insert it
           pred = newNode;                      // update current
           if (oldSucc == null) last = newNode; // update last if needed
           length++;
         }
         public void remove() {
           // pre: pred != last (current is valid)
           // post: pred.succ becomes pred.succ.succ
           if (atEnd()) // no element available to remove!
             throw new IteratorException(
               "Iterator.remove() applied at end of List");
           Node deadNode = pred.succ;
           pred.succ = deadNode.succ;
           if (last == deadNode) last = pred;
           length--;
         }
       }
     }

BiLists and Their Iterators

The lists of the last few sections can be efficiently scanned in only one direction, starting at the front and proceeding element by element to the end. We would now like to develop a more general form of list that supports both forward and backward traversal. The new list implementation will use doubly-linked lists, and we will call the new lists BiLists and their iterators BiIterators.

As with our previous lists, we define a pair of interfaces, IBiList for lists, and IBiIterator for iterators. Since BiLists and BiIterators will support all the same operations as Lists and Iterators, we will make these interfaces subinterfaces of IBiList and IBiIterator.

IBiList supports an additional operation for removing nodes at the rear of the list, and provides an additional factory method for producing IBiIterators.

     interface IBiList extends IList {
       void remRear();
       BiIIterator newBiIterator();
     }

IBiIterator supports backward traversal of lists, and so requires methods for moving to the end of a list, moving back an element, and a test for whether the iterator is at the front of the list.

     interface IBiIterator extends IIterator {
       void last();
       void prev();
       boolean atBeginning();
     }

Since a IBiIterator is also necessarily an Iterator I, a BiIterator I instance can be substituted for an Iterator I instance. Thus the newIterator and newBiIterator methods can share the same implementation.

An implementation of BiList and BiIterator is given below. In contrast to the List implementation of the last section, all the classes are top-level, and so imperative operations are not as well hidden. The BiIterator must now have a field to record the BiList it operates on, and this must be initialized at construction time. As an exercise, try converting the implementation so it uses nested and inner classes as in the List case.

The underlying list structure is doubly-linked and circular. The dummy node acts as both a marker for the beginning and the end of the list. Since nodes have pointers to both next and previous nodes, the insertion and deletion methods are a little more tricky and require more elaborate pointer juggling. When implementing doubly-linked lists yourself, it helps to draw diagrams that show what points to what at each stage of one of these operations.

In the following code defining the interface IBiList class BiList the in-

terfaces IList, ReadIterator, and Iterator I and classes ListException and

IteratorException are identical to those in the code for the (singly-linked)

class MutList.
     interface IBiList extends IList {
       // Extra operation required for bi-directional traversal
       // Standard implementation uses double linking
       void remRear();
       IBiIterator newBiIterator();
       // duplicate of newIterator() with more precise return type
     }
     interface IBiIterator extends IIterator {
       // extended iterator for IBiList

void last();
void prev();
}
class BiList implements IBiList {
// ** fields **
Node head = new Node();  // allocate circularly linked header node
int length = 0;
// ** constructors **
// relying on default constructor
// ** toString
public String toString() {
  IBiIterator i = new BiIterator(this);
  String result = "(";
  for (i.first() ; ! i.atEnd(); i.next())
    result = result + " " + i.currentItem();
  return result + " )";
}
// ** methods in Interface IBiList
public IList newList() { return new BiList(); }
public IBiList newBiList() { return new BiList(); }
public int length() { return length; }
public void insertFront(Object o) {
  Node oldSucc = head.succ;
  Node newNode = new Node(o,head,oldSucc);  // allocate new Node
  // insert new Node
  head.succ = newNode;
  oldSucc.pred = newNode;
  length++;
}
public void insertRear(Object o)  {
  Node oldPred = head.pred;

Node newNode = new Node(o,oldPred,head);  // allocate new Node
// insert new Node
head.pred = newNode;
oldPred.succ = newNode;
length++;
}
public void remFront() {
if (isEmpty())
  throw new ListException("remFront() applied to EmptyList");
else {
  Node newSucc = head.succ.succ;
  head.succ = newSucc;
  newSucc.pred = head;
  length--;
}
}
public void remRear() {
if (isEmpty())
  throw new ListException("remRear() applied to EmptyList");
else {
  Node newPred = head.pred.pred;
  head.pred = newPred;
  newPred.succ = head;
  length--;
}
}
public boolean isEmpty() { return head == head.succ; }
public IIterator newIterator() {
// weaker typing for BiIterator when viewed as Iterator
return new BiIterator(this);
}
public IBiIterator newBiIterator() {
return new BiIterator(this);
}
}
// Implementation classes (not hidden!)

class Node {
// ** fields **
Object item;
Node pred,succ;
// ** constructors
Node(Object i, Node p, Node s) {
  item = i;
  pred = p;
  succ = s;
}
Node() {  // allocate header
  item = null;
  pred = this;
  succ = this;
}
}
class BiIterator implements IBiIterator {
// ** fields **
BiList listThis;
Node current;
// ** constructors **
BiIterator(BiList l) {
  listThis = l;                  // associated List instance
  current = listThis.head.succ;  // current is first item (if one exists)
}
// ** methods in BiIterator interface **
public void first() {
  current = listThis.head.succ;  // current is first item (if one exists)
}
public void last() {
  current = listThis.head.pred;  // current is last item (if one exists)
}
public void next() {
  current = current.succ;        // wraps around end
}

public void prev() {
current = current.pred;        // wraps around end
}
public Object currentItem() {
if (current == listThis.head) throw
  new IteratorException("No current element in " + listThis);
return current.item;
}
public boolean atEnd() { return current == listThis.head; }
public void insert(Object o) {
// pre: true
// post: Node containing o is inserted before current item,
//       current is unchanged
Node oldPred = current.pred;
Node newNode = new Node(o, oldPred, current);  // allocate new node
current.pred = newNode;                        // insert it
oldPred.succ = newNode;
listThis.length++;
}
public void remove() {
// pre:  current is valid
// post: current becomes current.succ
if (current == listThis.head) throw
  new IteratorException(
    "BiIterator.remove() applied at end of BiList " + listThis);
Node cPred = current.pred;
Node cSucc = current.succ;
cPred.succ = cSucc;
cSucc.pred = cPred;
current = cSucc;
listThis.length--;
}
}

2.1.9 Alternate Representations of Lists

So far we have focused on linked lists, but these are not the only possible implementations. Arrays Here is a sketch of how we might implement the BiList interface using arrays.

1. newList: allocate a new array and initialize the elements in order (O(n))

2. isEmpty: trivial if along with the array we maintain a count of how many elements are actually in use (O(1))

3. insertFront: expensive. If the front is already occupied, we have to shuffle all the contents one place further down the array (O(n);

4. insertRear: cheap (O(1));

5. remRear: cheap (O(1));

6. newIterator: we don’t present an implementation, but the details are not hard. The iterator need only keep the index of the current element. But note that inserting or deleting from the middle now requires shuffling elements (O(n) average case).

If we run out of room we can resize the array used to store the list elements. If we double the size of the array at each resizing, then the average number of times an element is copied due to resizing is approximately 1. To prove this, let the initial size of the array be I, and suppose that the final size of the array is N, and there were k resizes. Then

N = I · 2^{k}

and we observe that 1. the first I elements move k times;

2. the next I elements move k-1 times;

3. the next 2I elements move k-2 times;

4. ...

5. the last N/2 = 2^{k−1} · I elements move 0 = k − k times.

Using some summation facts we can show that the total number of array element copy operations is exactly N - I. Thus the average number of copy operations per element in the final array is (N-I)/N which is always less than 1, and approaches 1 in the limit as N gets much larger than I (i.e. as the number of resizings gets large). We say that the amortized cost of copying array elements is (bounded by a) constant. The strategy of doubling the size of the array on each resize operation appears to be an efficient one.

Finger Exercise 2.1.9.1: Suppose that instead of doubling the array size, we increased it by some constant amount. That is, after k resizings, the size of the array is I + k · J for some constant J. What would the amortized cost of element copying be then?

2.1.10 Hybrid Representations of Sequences

For some application such as a text editor the best representation of sequences may be a hybrid of linked and sequential allocation, sometimes called a rope implementation. Such a hybrid links together sequentially allocated blocks of elements (arrays). If the size of blocks is bounded by a constant, then the asymptotic complexity of sequence operations in the hybrid implementation is identical to the corresponding singly or doubly linked list implementation. But the leading constant in the polynomial approximating the running time may be much lower.

Graphical User Interfaces

Nearly all contemporary software applications have a graphical user interface. A well-designed graphical interface is far more expressive and easier to use than a text based interface. In this section, we will show how to write simple graphical user interfaces in Java.

3.1 GUI Programming

Graphical user interface programming is inherently more complex than ordinary applications programming because the graphical interface computation is driven by a stream of graphical input actions. All of the input actions performed by a program user including moving the mouse, clicking a mouse button, and typing a keystroke are processed by code in the computer operating system. This code determines when an input action of potential interest to the application occurs. Such an input action is called an “event”. Typically mouse movement alone does not constitute an event; the operating system updates the position of the cursor on the screen as the mouse is moved. When a mouse button is clicked or a key is typed, the operating system interrupts the application program and informs it that the specified event has occurred. The Java virtual machine includes an event monitor that processes these interruptions. This event processing code filters input events just as the operating system code filters inputs. For some events such as typing a key (other than return), the Java event monitor simply echoes the character on the screen in the appropriate place. For other events such as a mouse click on a button, the Java event monitor generates a program Event object that it places on a queue of pending Events for processing by the running Java program.

Every Java program that creates graphical components has an extra thread of execution that processes the program Event objects in the event queue. For each program Event object, the thread calls the “listener” method that has been registered by the Java program for handling this kind of program event.

3.1.1 Model-View-controller Pattern

A well-organized graphical application has three components:

• a model consisting of the application with no external interface;

• a view consisting of one or more graphical frames that interact with the user and the application; and

• a controller consisting of the “main” program that constructs the model and the view and links them together.

A model is a “raw” program module with a programming interface consisting a collection of publicly visible methods or procedures. In Java, the application is typically a single object (containing references to many other objects) and the programming interface is the collection of methods supported by that object.

When a program with a graphical interface starts, the controller

1. creates the model (application),

2. creates the view consisting of one or more graphical frames and attaches commands to the graphical input controls (buttons, text boxes, etc.) of the view,

3. activates the graphical components in the view, and

4. terminates.

Of course, program execution continues after the controller terminates because the extra thread of execution that processes program events is still running. After the controller terminates, all program execution is triggered by user input actions.

The commands attached to the graphical input controls are operations on the model implemented using the model’s programming interface. Recall the command pattern from Section 1.11. In Java, each of the graphical input controls in the graphics (AWT/Swing) library has an associated command interface that the installed commands implement. In the Java graphics library, these commands are called “listeners” because the are dormant until a graphical input event occurs (e.g., a button is “pressed”). In the programming literature, these commands are often called “callbacks” because they call methods “back” in the model which is logically disjoint from the code running in the view.

To explain how to write programs using the model-view-controller pattern, we will explore a simple example, namely a click-counter application that maintains and displays a simple integer counter ranging from 0 to 999. The graphical display will show the current value of the counter and include three buttons: an increment button, a decrement button, and reset button.

We will start with the problem of writing the view components of the application.

3.1.2 How to Write a View

Most view components have a small number of distinct states that determine how the view is configured and how it will respond to the next program event. As a result, view component programs typically consist of:

• a constructor than initializes the view,

• a registration method for each program event source (e.g., a button) that takes a callback (command) argument and registers this callback as the listener for this event source, and

• a setter method for each distinct view state that sets the fields of the view to the appropriate values.

The controller uses the registration methods to attach callbacks to program event sources in the view. The callbacks use the setter methods to change the state of the view in response to program events.

For our click counter example, the view will have the following format: which we decompose into three possible states:

1. the Min state where the DEC and 0 buttons are deactivated because the counter has its minimum value of 0.

2. the Counting state where are three buttons are activated, and

3. the Max state where the INC button is deactivated because the counter has reached its maximum value.

The listener must take into account the state of the model to update view.

Warning Most introductory books are written by authors who do not understand the model-view-controller pattern and the proper use of callbacks. Callbacks are not supported in C or C++ because there is no concept of procedures as data objects (simply passing a pointer to a procedure does not work!). As a result, textbook authors with a C/C++ background are accustomed to using ugly alternatives to callbacks which they continue to use in the context of Java. A common and particularly onerous abuse of the Java callback interface is implementing the requisite listener interfaces by methods in the main viewer class, which is typically a frame or an applet. This approach limits each event category to a single callback, e.g, one callback method for all buttons which is coded in the main viewer class.

This approach has four serious disadvantages.

• First, to determine which component produced a given event, the viewer class must uniquely label each event source (or maintain a table of event references).

• Second, when the listener receives an event, it must classify the event source using a sequence of tests or look it up in a table to determine what block of code should be used to process the event.

• Third, the code to process the event embedded in the view relies on the interface provided by the application, corrupting the model-view-controller decomposition.

• Fourth, if another graphical component in the same category is added to the view (e.g, a new button) then the code for the callback method for that event category must be modified.

Since the command pattern (procedures as data objects) completely eliminates this mess, the “view class as listener” approach to event processing is indefensible. Nevertheless, it is widely taught even by some reputed “experts” on Java programming. In fact, I am familiar with only one popular Java book that teaches good programming practice in conjunction with GUI programming, namely Thinking in Java by Bruce Eckel.

Coding the View Class The following code defines a view class that supports the schematic display given above:

     import java.awt.*;
     import java.applet.*;
     import java.awt.event.*;
     import javax.swing.*;
     class ClickCounterView {
       // ** fields **
       private JButton incButton;
       private JButton resetButton;
       private JButton decButton;
       private JLabel valueDisplay;
       // ** constructors **
       public ClickCounterView(JApplet itsApplet) {
         JPanel controlPanel = new JPanel();

CHAPTER 3. GRAPHICAL USER INTERFACES 147

itsApplet.getContentPane().setLayout(new BorderLayout());
valueDisplay = new JLabel("000", JLabel.CENTER);
itsApplet.getContentPane().add(valueDisplay, "Center");
incButton = new JButton("+");
resetButton = new JButton("0");
decButton = new JButton("-");
controlPanel.add(incButton);
controlPanel.add(resetButton);
controlPanel.add(decButton);
itsApplet.getContentPane().add(controlPanel, "South");
}
// ** methods **
public void setValueDisplay(String setTo) {
valueDisplay.setText(setTo);
}
public void addIncListener(ActionListener a) {
incButton.addActionListener(a);
}
public void addDecListener(ActionListener a) {
decButton.addActionListener(a);
}
public void addResetListener(ActionListener a) {
resetButton.addActionListener(a);
}
public void setMinimumState() {
incButton.setEnabled(true);
resetButton.setEnabled(false);
decButton.setEnabled(false);
}
public void setCountingState() {
incButton.setEnabled(true);
resetButton.setEnabled(true);
decButton.setEnabled(true);
}

       public void setMaximumState() {
         incButton.setEnabled(false);
         resetButton.setEnabled(true);
         decButton.setEnabled(true);
       }
     }

The structure of this program is very simple. Most of the length is due to Java’s wordy syntax and long variable names. What does it do?

The Java AWT/Swing library includes a large number of classes for defining graphical components that can be displayed on the screen. The AWT library relies on the window manager of the underlying operating system to implement common graphical components like windows, menus, and buttons. The Swing extension to the AWT library provides “pure Java” equivalents of these graphical elements, eliminating the vagaries in graphical style among window systems. For every graphical component class C in the AWT library, the Swing extension includes an equivalent “pure Java” component class JC. For example, the AWT library includes a component class Button to represent a button in a graphical window. Hence, the Swing extension includes the corresponding class JButton. With a few exceptions, each Swing component class can be used in place of the corresponding AWT class.

All of the component classes in AWT/Swing are all descendants of the AWT abstract class Component (surprise!). The ClickCounterView class mentions three of these component classes, namely JPanel, JButton, and JLabel which are all subclasses of the Swing abstract class JComponent (which is a subclass of Component). A JPanel object is simply a rectangular region that can be incorporated in a graphical container (such as the JFrame class in the Swing library) which is subsequently displayed on the screen. A panel typically contains other graphical elements (e.g. buttons, drawing canvases, text, pictures) which are displayed as part of the panel. (A blank panel is not very interesting!) A JButton is a graphical button and a JLabel is a single line of text that can be used as a graphical component.

In AWT/Swing library, graphical components that can contain other graphical components are called containers and belong to type Container. Within a container, the layout of the graphical components inside it is determined by a layout manager, a Java object of type LayoutManager. One of the important tasks in programming a user interface is determining which layout manager and combination of parameter values to use. A good layout policy will produce an attractive logical layout for a variety of different frame shapes and sizes.

The layout manager BorderLayout used in ClickCounterView uses compass points to constrain the relative position of graphical components. The four

compass points, "North", "East","South", and "West" plus the "Center" po-

sition are supported by the layout manager as directives when graphical components are installed within the panel that it manages.

In the program text above, the ClickCounterView object constructor creates panels within the frame provided by the JApplet object that is passed to it. An applet is a “top level” container; the browser that is executing the applet provides it with a blank area of the screen in which it can paint its various graphical elements. The constructor for ClickCounterView creates a JLabel displayValue text line to hold the current click count and

a JPanel controlPanel containing three buttons, incButton, resetButton,

and decButton with adorning String labels +, 0, and -, respectively. The displayValue is placed in the center of the Applet and the three buttons are placed in the controlPanel using the default layout manager FlowLayout. This layout manager places graphical components in rows from left-to-right just like a text editor places characters when you type text into a buffer.

The ClickCounterView class defines seven public methods to access and update its components:

• the setValueDisplay method update the valueDisplay to the specified

     String;
   • the methods addIncListener, addDecListener, and addResetListener

attach their command arguments to the buttons incButton, resetButton, and decButton, respectively; and

   • the methods setMinimumState, setCountingState, and setMaximumState

which enable and disable the buttons appropriately for each of the three states described above.

3.1.3 How to Write a Simple Model

From the perspective of GUI design, the critical issue in developing a model is defining the interface for manipulating the model. This interface should be as transparent as possible, without making a commitment to a particular user interface.

In our click counter example program, the model class is utterly trivial. In accordance with the model-view-controller pattern, it is does not presume any particular user interface. The only feature of the counter targeted at supporting a user interface is the toString method which pads the output String with leading zeroes to produce the specified display width of 3 digits.

     class ClickCounter {

// ** fields **
private static final int MAXIMUM = 999;
private static final int MINIMUM = 0;
private static final int STRING_WIDTH = 3;
private static int count = MINIMUM;
// ** constructor
public ClickCounter() {}
// ** methods
public boolean isAtMinimum() { return count == MINIMUM; }
public boolean isAtMaximum() { return count == MAXIMUM; }
public int inc() {
  if (! this.isAtMaximum()) count++;
  return count;
}
public int dec() {
  if (! this.isAtMinimum()) count--;
  return count;
}
public void reset() { count = MINIMUM; }
public int getCount() { return count; }
// ** toString() **
public String toString() {
  StringBuffer buffer =
    new StringBuffer(Integer.toString(count));
  while (buffer.length() < STRING_WIDTH) buffer.insert(0,0);
  return buffer.toString();
}
}

3.1.4 How to Write a Controller

From a programming perspective, the controller is the most interesting part of this example. It glues together the model and view using callbacks and then terminates. Of course, whenever the view receives an event, it invokes callbacks, code defined in the controller, to process them. The controller’s callback code performs whatever updates are required to the model an to the view.

public class ClickCounterControl extends JApplet {
  // ** fields **
  private ClickCounter counter;
  private ClickCounterView view;
  // ** constructors **
  // relying on default constructor
  // ** methods **
  // relying on inheritance from JApplet
  public void init() {
    counter = new ClickCounter();
    view = new ClickCounterView(this);
    view.setMinimumState();
    view.setValueDisplay(counter.toString());

    view.addIncListener(new ActionListener(){
      public void actionPerformed(ActionEvent event) {
if (counter.isAtMaximum()) return;
if (counter.isAtMinimum()) view.setCountingState();
counter.inc();
view.setValueDisplay(counter.toString());
if (counter.isAtMaximum()) view.setMaximumState();
      }
    });
    view.addDecListener(new ActionListener(){
      public void actionPerformed(ActionEvent event) {
if (counter.isAtMinimum()) return;
if (counter.isAtMaximum()) view.setCountingState();
counter.dec();
view.setValueDisplay(counter.toString());
if (counter.isAtMinimum()) view.setMinimumState();
      }
    });

         view.addResetListener(new ActionListener(){
           public void actionPerformed(ActionEvent event) {
     counter.reset();
             view.setMinimumState();
             view.setValueDisplay(counter.toString());
           }
         });
       }
     }

3.2 What is Concurrent Programming?

Until now, we have been exclusively concerned with sequential programs that execute a single stream of operations. Even the GUI programming in the previous section avoided concurrent execution by terminating the controller as soon as it finished setting up the model and view. Concurrent computation makes programming much more complex. In this section, we will explore the extra problems posed by concurrency and outline some strategies for managing them.

In a concurrent program, several streams of operations may execute concurrently. Each stream of operations executes as it would in a sequential program except for the fact that streams can communicate and interfere with one another. Each such sequence of instructions is called a thread. For this reason, sequential programs are often called single-threaded programs. When a multi-threaded program executes, the operations in its various threads are interleaved in an unpredictable order subject to the constraints imposed by explicit synchronization operations that may be embedded in the code. The operations for each stream are strictly ordered, but the interleaving of operations from a collection of streams is undetermined and depends on the vagaries of a particular execution of the program. One stream may run very fast while another does not run at all. In the absence of fairness guarantees (discussed below), a given thread can starve unless it is the only “runnable” thread.

A thread is runnable unless it executes a special operation requiring synchronization that waits until a particular condition occurs. If more than one thread is runnable, all but one thread may starve (make no progress because none of its operations are being executed) unless the thread system makes a fairness guarantee. A fairness guarantee states that the next operation in a runnable thread eventually will execute. The Java language specification currently makes no fairness guarantees but most Java Virtual Machines guarantee fairness.

Threads can communicate with each other in a variety of ways that we will discuss in detail later in this section. The Java programming language relies primarily on shared variables to support communication between processes, but it also supports an explicit signaling mechanism.

In general, writing concurrent programs is extremely difficult because the multiplicity of possible interleavings of operations among threads means that program execution is non-deterministic. For this reason, program bugs may be difficult to reproduce. Furthermore, the complexity introduced by multiple threads and their potential interactions makes programs much more difficult to analyze and reason about. Fortunately, many concurrent programs including most GUI applications follow stylized design patterns that control the underlying complexity.

To demonstrate some of the subtle problems that arise with this sort of programming, consider the following example. We have two threads, A and B, that both have access to a variable ct. Suppose that, initially, ct is 0, but there are places in both A and B where ct is incremented.

A  B
...  ...

    ct++;  ct++;

To increment a variable x, (i) the value v of x must be fetched from memory, (ii) a new value v’ based on v, and (iii) v’ must be stored in the memory location allocated to variable x. These are three separate actions, and there is no guarantee that no other thread will access the variable until all three are done. So it’s possible, for instance, that the order of operations from these two threads occurs as follows:

     A fetches ct = 0
     B fetches ct = 0
     A computes the value ct++ = 1
     A stores the value 1 in ct
     B computes new value ct++ = 1
     B stores the value 1 in ct

With this order of the operations, the final value for ct is 1. But in other possible orderings (e.g., if A performs all of its actions first), the final value would be 2.

A simple strategy for preventing this form of interference (often called a race condition) is to make the entire access/modify/store cycle for updating shared variables atomic, despite the fact that the cycle is performed using several machine instructions. Atomic operations appear to execute as a single machine instruction because all other threads are forced to pause executing while the atomic operation executes. As a result, it is impossible for another thread to observe the value of the updated variables while the operation is in progress. A block of code that requires atomic execution is called a critical section. Some programming languages that support concurrency include begin/end brackets for enclosing critical sections.

The critical section mechanism works well in the context of running multithreaded programs on a computer with a single processor (a uniprocessor) since it reduces to ensuring that a critical section is not interruptible (permitting another thread to run). But it is clumsy and inefficient on a multiprocessor because it forces all processors but one to stop execution for the duration of a critical section. (Existing virtual machines treat new operations that force garbage collection as critical sections.

A much better mechanism for preventing interference in concurrent programs that may be executed on multiprocessors is locking data objects. When a data object is locked by a thread, no other thread can access or modify the data object until the locking thread releases it. In essence, locking relaxes the concept of atomic execution so that it is relative to a particular object. Threads can continue executing until they try to access a locked object.

Java relies on object locking to prevent interference. An object can be locked for the duration of a method invocation simply by prefixing the method declaration with the work synchronized. For instance, to define a synchronized increment method, we would write:

     synchronized void inc() { ct++; }

We can also declare static methods to be synchronized, which locks the class object (which contain all of the static variables of the class) rather than an instance object.

An unusual feature of Java’s lock mechanism is the fact that locking an object only inhibits the execution of operations that are declared as synchronized. Methods that are not declared as synchronized will be executed even when an object is locked! There is a strong argument for this capability: it supports the definition of classes that partition operations in two groups: those that require synchronization and those that do not. But it also invites subtle synchronization bugs if the synchronized modifier is inadvertently omitted from one method definition.

Of course, concurrency only arises in Java when a program uses more than one thread. To support the explicit creation of new threads, Java includes a built-in abstract class Thread, that has an abstract method run(). We can make a new thread by (i) defining a class extending Thread that defines the method run(), (ii) constructing a new instance of this class, and (iii) calling the start() method on this new instance. The start() method actually creates a new thread corresponding to the receiver object (a Thread) and invokes the run() method of that thread, much as the main() method is invoked in the root class when you run a Java Virtual Machine. For example,

     class Foo extends Thread {
       // must have
       public void run() {
       ...
       }
     }

When a constructor for Foo is called, all of computation for the object allocation and constructor invocation is performed in the current thread; a new thread is not created until the start() method is invoked for a Thread() object. To create and start a new Foo thread, the current thread can simply execute the code

     Thread t = new Foo();
     t.start();

Alternatively, the current thread can execute the run() method of the Thread object t simply by performing the operation

     t.run()

instead of

     t.start()

Assume that a new Foo thread t has been created and started. At some point in the execution of the original thread (now running concurrently with thread t) can wait for thread t to terminate by executing the method invocation:

     t.join();
     // waits for the thread object to terminate.

So we can view the relationship of the two threads of control as follows:

     main
       |
     t.start
       |\
       | \
       | |
       | /
       |/
     t.join
       |
       |

Synchronizing multiple threads does incur some overhead. For example, consider the following Java code:

class PCount extends Thread {
// ** fields ***
static int sharedCtr = 0;
static final int cntPerIteration = 100000;
static final int noOfIterations = 10;
int id;
// ** constructors **
PCount(int i) { id = i; }
// ** methods **
void inc() {
  sharedCtr++;
}
public void run() {
  for (int i = 0; i < cntPerIteration; i++) inc();
  System.out.println("Iteration #" + id +
    " has completed; sharedCtr = " + sharedCtr);
}
public static void main(String[] args)
    throws InterruptedException {
  Thread[] tPool = new Thread[noOfIterations];
  for (int j = 0; j < noOfIterations; j++) {
    tPool[j] = new PCount(j);
  }
  for (int j = 0; j < noOfIterations; j++) {
    tPool[j].start();
  }
  for (int j = 0; j < noOfIterations; j++) {
    tPool[j].join();
  }
  System.out.println("Computation complete.  sharedCtr = "
    + sharedCtr);
}
}

In each iteration, main creates a new thread. Afterwards, all are synchronized and a final value is determined.

The counter is not locked in this example, and so updates may be lost because of the problems described above. The likelihood with which update losses may occur varies depending on the number of threads. For example, in a test that I ran a few months ago

• for 1 million iterations, the program 65

• for 100,000 iterations, the program lost none.

Apparently, even with 100,000 threads, each iteration occurred within a single time slice.

Synchronizing the threads fixes the problem of lost updates, but it really slows the program down; even for 100,000 iterations.

In modern event-handling models such as those in Java and DrScheme, we have a single event handler that executes events serially. This protocol saves the overhead of synchronization and eliminates potential deadlocks (which we will discuss later).

Synchronized methods and statements

We’ve already discussed synchronized methods above. We can likewise declare synchronized blocks of statements using the following syntax:

     synchronized(expr) {
        ...
     }

where ¡tt¿expr¡/tt¿ must evaluate to a reference type (i.e. for most purposes, an object reference). The code between the braces needn’t have any connection to ¡tt¿expr¡/tt¿, although this would be perverse in most situations.

For an example, consider implementing a bounded buffer protocol. Bounded buffers are used in Unix to allow interprocess communication via ¡em¿pipes¡/em¿. When the output of one process is piped to the input of another, a bounded buffer is set up between the processes. The first process writes into the buffer, and the second process reads from it.

When the second process attempts to read data but there is none for it to read, it must wait until data is available. There are two general schemes that can be used to implement this waiting:

• busy-waiting: the reader process executes a loop in which it tests whether there is any data waiting for it. This approach works but is undesirable because the reader is allocated CPU time just to test if it can do useful computation;

• blocking: the operating system provides the routines for reading from and writing to the buffer. When the reader process attempts to read from an empty buffer, it is ¡em¿blocked¡/em¿. The operating system marks the reader as unable to run until data is available in the buffer, and removes it from the collection of currently runnable processes. When the writer process writes data, the system ”wakes up” the reader process and makes it one of the runnable processes again. In this way, very little CPU time is expended on coordinating the writer and reader processes.

Naturally, Unix uses the second approach. Note that a symmetric situation occurs when the writer process attempts to write to the buffer and it is full: the writer is blocked pending the availability of space in the buffer. The writer may be reawakened when the reader reads some data and the routine for reading from the buffer determines that there is available space for more data to be written.

We can program a Java thread so that it busy waits on a locked object, but this is almost always a bad programming practice. Fortunately, Java provides a facility that enables us to avoid busy-waiting. There are two primitives,

   • wait()
   • notify()

which can be used inside synchronized methods and blocks. wait() puts the calling thread to sleep on a queue of threads that are waiting for some change in the status of a locked object. The thread can only be awakened if some other running thread calls notify() on the same locked object. When notify() is invoked for the locked object, a check is done to see whether any change has been made to it, and then some waiting thread from the associated queue is allowed to run. (Notify arbitrarily picks one thread).

The awakened process will, if prudently written, check whether the condition it is waiting for actually holds. If it does, the thread proceeds, and otherwise it should suspend itself again with another call to wait(). Usually, this is implemented in a loop. This may look a bit like busy-waiting, but because wait() suspends the calling thread until something of interest happens, most of the time the thread is idle.

There is also a notifyAll() method, which works just like notify() except that all waiting threads are allowed to run.

3.2.1 Deadlock

In the preceding subsection, we showed how object locking can be used to prevent interference. Unfortunately, locking is not a panacea. The excessive use of locking can severely degrade system performance or, even worse, lock up the system so that all computational progress halts until the program is terminated. The essence of concurrent programming is organizing computations so that neither interference or deadlock can occur.

To illustrate deadlock, let us consider a classical problem called The Dining Philosophers in the theory of concurrent programming. It is unrealistic and fanciful, but the synchronization behavior that it models can happen in real systems.

A collection of N philosophers sits at a round table, where N > 1. N forks are placed on the table, one between each pair of adjacent philosophers. No philosopher can eat unless he has two forks and he can only use the two forks separating him from his two neighbors. Obviously, adjacent philosophers cannot eat at the same time. Each philosopher alternately eats and sleeps, waiting when necessary for the requisite forks before eating.

Our task is to write code simulating the dining philosophers so that no philosopher starves. An obvious protocol that each philosopher might follow is:

     while (true) {
       grab left fork;
       grab right fork;
       eat;
       release left fork;
       release right fork
       sleep;
     }

Now assume that actions of the philosophers are perfectly interleaved: the first philosopher grabs his left fork, then the second philosopher grabs his left fork, and so on until the Nth philosopher grabs his left fork. Then the first philosopher tries to grab his right fork, the second philosopher tries to grab his right fork, and so on. They all have to wait because no right fork is available and they all starve.

Theoretical computer scientists have proven that there is no deterministic uniform solution to this problem. (By uniform, we mean that every philosopher executes exactly the same code with no access to identifying state information such as the name of the “current” philosopher.) But many non-uniform solutions exist. For example, we could number the philosophers around the table from 0 to N. Even numbered philosophers ask for the left fork first, odd numbered ones ask for the right fork first.

Another common solution to this sort of deadlock is to order the resources (in this case forks) and force the processes (philosophers) to grab forks in ascending order. This solution is very general and is widely used in practice.

Consider the case where we have three philosophers: P1, P2, P3 and three forks F1, F2, F3 where P1 sits between F1 and F2, P2 sits between F2 and F3, and P3 sits between F3 and F1. We can order the forks in the obvious ascending order F1, F2, F3.

Now, no matter what, all the philosophers will be able to eat because the linear ordering prevents cycles in the “is waiting for” relation among philosophers.

Finger Exercise: Sketch an informal for this assertion!

For instance, if P0 gets F0, and P1 gets F1, P2 must wait until F0 is free. So P1 will get F2 (since there will be no contention), and finish eating. This will release F1 and F2, allowing P0 to get F1 and finish eating. Finally, this will release F0, allowing P2 to get F0 and F2 (since there will be no further contention) and finish eating.